### v2

# Linear response for skew-product maps with contracting fibres

Abstract.
We study linear response for families of skew-product dynamical systems with contracting fibres. Our approach is based on a sectional transfer operator acting on families of probability measures along the fibres. The operator allows to describe invariant measures of the skew-product in terms of sample measures over the base dynamics, regardless of invertibility or non-invertibility of the base map. Under general assumptions, we establish existence and uniqueness of invariant sample measures and prove their differentiability, with respect to system parameters, in suitable topologies. As an application we obtain linear response for Bernoulli convolutions, which are of prime importance in the study of number theoretic problems and fractals. Another application of our results yields linear response for physical measures of solenoidal attractors with intermittency, an example of a hyperbolic system which cannot be handled by traditional transfer operator techniques.
Key words and phrases:
Linear Response, Bernoulli Convolution, Solenoid with Intermittency
2020 Mathematics Subject Classification:
37A05, 37C40, 37C75
[…]
1. Introduction
Ergodic theory is a field of mathematics that is concerned with studying dynamical systems from probabilistic point of view. It has deep connections with many areas of mathematics such as analysis, geometry and number theory. Starting with the work of Sinai, Ruelle and Bowen in the 70’s, significant advances have been made in the study of uniformly hyperbolic systems from statistical point of view [42, 20]. However, unlike their uniformly hyperbolic counterparts, nonuniformly and partially hyperbolic systems are less understood despite several landmark results, including [2, 3, 14, 26, 45]. These systems represent a broader and more complicated class of dynamical systems. When analysing probabilistic aspects of hyperbolic dynamical systems, it is important to understand how stable dynamical invariants, such as SRB measures [5, 23, 32], mixing rates [30] and entropy [35] are. In this domain, a far-reaching direction studies differentiability of such quantities with respect to parameters of the system [35].
More precisely, given a family of dynamical systems $T_{\alpha}:M\to M$, with $\alpha$ taking values in an interval $J\subset\mathbb{R}$, it is natural to ask whether dynamical quantities are differentiable, or not, with respect to the parameter $\alpha$ in a suitable topology. In particular, if for each $\alpha$ the system $T_{\alpha}$ admits a unique physical measure $\mu_{\alpha}$, it is natural to ask whether $\alpha\mapsto\mu_{\alpha}$ is differentiable in a suitable topology. When differentiability holds, it is customary to say that the system has linear response. Such a property was first proved by Ruelle in the framework of uniformly hyperbolic systems [41] followed by Dolgopyat in the setting of partially hyperbolic systems [27]. Since then many authors studied such a phenomena and higher order differentiability in a smooth uniformly hyperbolic setting, most notably by Gouëzel and Liverani [33], Butterley and Liverani [21], among others. However, beyond uniform hyperbolicity, the situation is more subtle. For instance, even for simple piecewise smooth expanding maps if perturbations (or parameters) are transversal to the topological class, linear response can fail to hold, see the work of Baladi [11, 12], Baladi and Smania [15] and related work on unimodal maps [13, 16, 17, 24]. On the other hand, there are other recent examples where linear response holds despite the perturbation being transversal to the topological class [9, 22]. The situation in smooth partially hyperbolic setting is not better. Despite the positive results of Dolgopyat [27], Zhang [47] studied a class of partially hyperbolic skew products where it was shown that linear response fails and that the map $\alpha\mapsto\mu_{\alpha}$ is only Hölder for certain partially hyperbolic systems.
Motivated by examples of partially hyperbolic systems that do not fall within the classes considered by [27] and [47], we consider in this work skew product maps with uniformly contacting fibres. Our approach is based on a sectional transfer operator acting on families of probability measures along the fibres. The operator allows to describe invariant measures of the skew-product in terms of sample measures over the base dynamics, regardless of invertibility or non-invertibility of the base map. Hence, this approach turns out to be not only fruitful to study linear response for certain partially hyperbolic systems but also to study linear response for measures, called Bernoulli convolutions, used in fractal geometry to analyse number theoretic questions posed by Erdős [29, 43]. Under general assumptions, we establish existence and uniqueness of invariant sample measures and prove their differentiability, with respect to system parameters, in suitable topologies. In one of our applications we obtain strong linear response for Bernoulli convolutions covering also the biased case addressed in [40]. Another application of our results yields linear response for physical measures of solenoidal attractors with intermittency, an example of a hyperbolic system which cannot be handled by traditional transfer operator techniques [19].
1.1. General setup and abstract results
Let $\Omega,X$ be compact Riemannian manifolds (possibly with boundary), and $d_{\Omega},d_{X}$ the respective Riemannian distances. Let also $m$ denote the Lebesgue (volume) measure defined on the Borel sets of $\Omega$. Throughout the paper, we use $\operatorname{diam}(X)$ to denote the diameter of $X$. Given an interval $J\subset\mathbb{R}$, we consider a family of skew-product maps $T_{\alpha}:\Omega\times X\to\Omega\times X$, with $\alpha\in J$, defined for $(\omega,x)\in\Omega\times X$ by
(1.1) $T_{\alpha}(\omega,x)=\big(f_{\alpha}(\omega),g_{\alpha}(\omega,x)\big).$
For each $\omega\in\Omega$, we write $g_{\alpha,\omega}:X\to X$ for the fibre map
$g_{\alpha,\omega}(x)=g_{\alpha}(\omega,x).$
We introduce below a set of general assumptions that provide an abstract framework for our main results.
  1. (A1)
    There exists a constant $0<\lambda<1$ such that
    $d_{X}(g_{\alpha,\omega}(x),g_{\alpha,\omega}(y))\leq\lambda d_{X}(x,y),$
    for all $x,y\in X$, all $\omega\in\Omega$ and all $\alpha\in J$.
  1. (A2)
    For each base map $f_{\alpha}:\Omega\to\Omega$, the following hold:
    1. (a)
      there exists a countable $m$ modulo zero partition of $\Omega$ into open sets on each of which $f_{\alpha}$ is a $C^{1}$ diffeomorphism onto its image;
    2. (b)
      the map $f_{\alpha}$ has an invariant probability measure $\eta_{\alpha}$ that is absolutely continuous with respect to $m$, with density $\rho_{\alpha}\in L^{1}(m)$.
It is well known (see e.g. [7]) that for every probability measure $\mu$ invariant under the skew-product map $T_{\alpha}$, whose marginal on $\Omega$ is $\eta_{\alpha}$, there exists an essentially unique family of sample measures $\{\nu_{\alpha,\omega}\}_{\omega\in\Omega}$ on $X$ such that
(1.2) $\mu_{\alpha}=\int_{\Omega}\nu_{\alpha,\omega}\,d\eta_{\alpha}(\omega).$
Under assumptions (A1)-(A2) we are able to construct these sample measures explicitly, in a manner that facilitates the analysis of their differentiability in some families of skew-product maps. To this end, let $\mathcal{M}_{1}(X)$ denote the space of Borel probability measures on $X$. Considering in $\mathcal{M}_{1}(X)$ the Wasserstein-1 distance $W_{1}$, it becomes a complete metric space, and since we assume $X$ a compact set, then $\mathcal{M}_{1}(X)$ is bounded with the distance $W_{1}$; see Appendix A for details. Define the space of probability-valued sections on $\Omega$,
$\mathbf{P}=\big\{\bm{\nu}:\Omega\to\mathcal{M}_{1}(X)\;\big|\;\bm{\nu}\text{ is measurable}\big\},$
with the identification of sections that coincide $\eta_{\alpha}$-almost everywhere. We consider $\mathbf{P}$ equipped with the essential supremum metric. Since $(\mathcal{M}_{1}(X),W_{1})$ is bounded, this metric is well defined, and $\mathbf{P}$ is in fact a complete metric space.
We now introduce an operator that provides a natural framework to study the evolution of the sample measures $\nu_{\omega}$ under the skew-product dynamics. For a given $\bm{\nu}\in\mathbf{P}$ and for $\eta_{\alpha}$-almost every $\omega\in\Omega$, the sectional transfer operator $\mathcal{K}_{\alpha}$ is defined by
(1.3) $(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}=\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|\,\rho_{\alpha}(\omega)}\,(g_{\alpha,\theta})_{*}\nu_{\theta}.$
Since the invariant density of the base map is preserved under its transfer operator, the contributions from all preimages of any point $\omega\in\Omega$ sum to one. In this sense, $\mathcal{K}_{\alpha}$ acts by pushing forward measures along the preimages, weighted according to a reverse Markov kernel. By the measurability of all functions involved in its definition, $\mathcal{K}_{\alpha}$ defines a map from $\mathbf{P}$ into itself. It can be interpreted as the transfer operator associated with the skew-product map $T_{\alpha}$ and its definition is motivated by the duality relation
$\langle\!\langle\mathcal{K}_{\alpha}\bm{\nu},\Phi\rangle\!\rangle=\langle\!\langle\bm{\nu},\Phi\circ T_{\alpha}\rangle\!\rangle,$
which holds for any bounded measurable function $\Phi:\Omega\times X\to\mathbb{R}$, where the pairing $\langle\!\langle\cdot,\cdot\rangle\!\rangle$ integrates the fiberwise action of $\nu_{\omega}$ against $\Phi(\omega,\cdot)$ with respect to $\eta_{\alpha}$; see Lemma 2.4.
Under the assumptions above, we obtain the following simple result, which serves primarily as a stepping stone and motivation for the more substantial developments that follow.
Theorem A.
Assume that (A1)(A2) hold. Then $\mathcal{K}_{\alpha}$ defines a contraction on $\mathbf{P}$, and therefore admits a unique fixed point $\bm{\nu}_{\alpha}\in\mathbf{P}$. Moreover,
$\mu_{\alpha}=\int_{\Omega}\nu_{\alpha,\omega}\,d\eta_{\alpha}(\omega)$
defines a $T_{\alpha}$-invariant probability measure on $\Omega\times X$.
We now introduce an additional set of assumptions under which linear response for the invariant measure $\mu_{\alpha}$ will be established. Our strategy is to first prove the differentiability of $\bm{\nu}_{\alpha}$ with respect to the parameter $\alpha$ in some suitable Banach space. This derivative cannot be taken directly in $\mathbf{P}$, since the space of probability measures lacks a vector space structure. We therefore make use of continuous inclusions
(1.4) $\mathcal{M}_{1}(X)\subset\operatorname{Lip}(X)^{*}\subset C^{k}(X)^{*},$
for $k\geq 1$, with respect to the natural norms on these dual spaces; see Appendix 4 for details. We also consider the corresponding subspaces
(1.5) $\operatorname{Lip}_{0}(X)^{*}\subset C^{k}_{0}(X)^{*},$
consisting of functionals on the respective dual that vanish on constant functions. In order to study the linear response of the skew-product measure, we begin by requiring some regularity for the fixed point $\bm{\nu}_{\alpha}$ of the sectional transfer operator. To begin with, we rewuire continuity with respect to the base point, which naturally leads us to consider the space of continuous probability-valued sections
$\mathbf{P}_{0}=\bigl\{\bm{\nu}:\Omega\to\mathcal{M}_{1}(X)\;\big|\;\bm{\nu}\text{ is }C^{0}\bigr\}.$
Under very general conditions, this space is invariant under the action of the sectional transfer operator. Given $\bm{\nu}_{\alpha},\bm{\nu}_{\alpha_{0}}\in\mathbf{P}_{0}$, the difference quotient
(1.6) $\frac{\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}}{\alpha-\alpha_{0}}$
takes values in the vector space
(1.7) $\mathbf{E}_{\mathrm{L}}^{0}=\bigl\{\,\bm{\nu}:\Omega\to\operatorname{Lip}_{0}(X)^{*}\;\big|\;\bm{\nu}\text{ is $C^{0}$}\bigr\}.$
which, when equipped with the supremum norm $\|\cdot\|_{\mathrm{L}}$, is a Banach space. Aiming for differentiability of the map $\alpha\mapsto\bm{\nu}_{\alpha}$ in the Banach space $\mathbf{E}_{\mathrm{L}}^{0}$ would however be too ambitious, given the weak regularity of the associated test functions. For this reason, we introduce for $k\geq 1$ the larger spaces
(1.8) $\mathbf{E}_{k}^{0}=\bigl\{\,\bm{\nu}:\Omega\to C_{0}^{k}(X)^{*}\;\big|\;\bm{\nu}\text{ is }C^{0}\bigr\},$
which are Banach spaces when endowed with the supremum norm $\|\cdot\|_{k}$. Note that, due to the continuous embeddings in (1.4), the sapce $\mathbf{E}_{\mathrm{L}}^{0}$ is continuously embedded in $\mathbf{E}_{k}^{0}$, for every $k\geq 1$.
The introduction of these spaces allows for weaker norms, under which the verification of the next conditions becomes feasible. We remark that the expression used to define the sectional transfer operator in (1.3) is still valid when applied to sections in these spaces; see equation (A.7) in Appendix A for the general definition of the push-forward.
  1. (A3)
    For all $\alpha\in J$, the operator $\mathcal{K}_{\alpha}$ acts as a bounded linear endomorphism of some $\mathbf{E}_{k}^{0}$. Moreover,
    1. (a)
      there exists $0<\lambda<1$ such that $\|\mathcal{K}_{\alpha}\|\leq\lambda$, for all $\alpha\in J$;
    2. (b)
      $\mathcal{K}_{\alpha}$ maps the vector subspace $\mathbf{E}_{\mathrm{L}}^{0}$ of $\mathbf{E}_{k}^{0}$ into itself;
    3. (c)
      the map $\alpha\mapsto\mathcal{K}_{\alpha}\bm{\nu}$ is continuous in $\mathbf{E}_{k}^{0}$ at $\alpha_{0}$, for any $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}^{0}$;
    4. (d)
      the map $\alpha\mapsto\mathcal{K}_{\alpha}\bm{\nu}_{\alpha_{0}}$ is differentiable in $\mathbf{E}_{k}^{0}$ at $\alpha_{0}$.
As a by-product of our approach, we obtain the following linear response result for the sample measures. We emphasise that the specific nature of the operator $\mathcal{K}_{\alpha}$ and of the spaces $\mathbf{E}_{\mathrm{L}}^{0},\mathbf{E}_{k}^{0}$ is immaterial for this theorem.
Theorem B.
Assume that (A1)(A3) hold. Then, the map $\alpha\mapsto\bm{\nu}_{\alpha}$ is differentiable in the space $\mathbf{E}_{k}^{0}$ at $\alpha_{0}$.
For the linear response of the skew-product measure, we need to impose additional regularity conditions both on the density $\rho_{\alpha}$ of the absolutely continuous invariant probability of the base map and on the fixed point $\bm{\nu}_{\alpha}$ of the operator $\mathcal{K}_{\alpha}$.
  1. (A4)
    The map $\alpha\mapsto\rho_{\alpha}$ is differentiable in $L^{1}(m)$ at $\alpha_{0}$.
  1. (A5)
    For every $\alpha\in J$, we have $\bm{\nu}_{\alpha}\in\mathbf{P}_{0}$ and $\|\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}\|_{k}\to 0$ as $\alpha\to\alpha_{0}.$
Under these additional assumptions, we obtain our first general linear response result for the skew-product measure provided by Theorem A. The order of differentiability $C^{k}$ of the observables appearing in the subsequent results is implicitly tied to the index $k$ of the space $\mathbf{E}_{k}^{0}$ for which properties (A3) and (A5) hold.
Theorem C.
Assume that (A1)(A5) hold. Then, for every $C^{k}$ function $\Phi:\Omega\times X\to\mathbb{R}$, the map
$\alpha\longmapsto\int_{\Omega}\int_{X}\Phi(\omega,x)d\nu_{\alpha,\omega}(x)\,d\eta_{\alpha}(\omega)$
is differentiable at $\alpha_{0}$.
In order to allow for applications in which the sample measures cannot be directly derived via Theorem B, we introduce a kind of minimal space in which differentiability of the sample measures implies differentiability of the skew-product measure. This added flexibility will be crucial for the application to systems whose base map has an inducing scheme, particularly, the intermittent solenoid presented in Subsection 1.4.2. To this end, we define for $k\geq 1$
(1.9) $\mathbf{L}_{k}^{0}=\Big\{\bm{\nu}:\Omega\to C^{k}_{0}(X)^{*}\;\big|\;\bm{\nu}\text{ is measurable and }\int_{\Omega}\|\nu_{\omega}\|_{k^{*}}\,d\eta_{\alpha_{0}}<\infty\Big\},$
and set for each $\bm{\nu}\in\mathbf{L}_{k}^{0}$,
(1.10) $\|\bm{\nu}\|_{\mathbf{L}_{k}}=\int_{\Omega}\|\nu_{\omega}\|_{k^{*}}\,d\eta_{\alpha_{0}}(\omega),$
where $\|\cdot\|_{k^{*}}$ denotes the standard norm in $C^{k}(X)^{*}$; see Appendix A. This defines a norm on $\mathbf{L}_{k}^{0}$ which makes it a Banach space. In fact, $\mathbf{L}_{k}^{0}$ coincides with the space of measurable functions from $\Omega$ to the Banach space $C_{0}^{k}(X)^{*}$ which are Bochner integrable with respect to the measure $\eta_{\alpha_{0}}$. Although we do not make this explicit in the notation, it should be noted that the space $\mathbf{L}_{k}$ is defined using the measure $\eta_{\alpha_{0}}$ corresponding to the reference parameter $\alpha_{0}$, the one at which we aim to study differentiability.
Theorem D.
Assume that (A1)(A2) and (A4)(A5) hold, and that the map $\alpha\mapsto\bm{\nu}_{\alpha}$ is differentiable in $\mathbf{L}_{k}^{0}$ at $\alpha_{0}$. Then, for every $C^{k}$ observable $\Phi\colon\Omega\times X\to\mathbb{R}$, the map
$\alpha\longmapsto\int_{\Omega}\int_{X}\Phi(\omega,x)\,d\nu_{\alpha,\omega}(x)\,d\eta_{\alpha}(\omega)$
is differentiable at $\alpha_{0}$.
Since convergence in the space $\mathbf{E}_{k}^{0}$ implies convergence in the space $\mathbf{L}_{k}^{0}$, Theorem C is a direct consequence of Theorem B and Theorem D. The proof of Theorem B will be presented in Subsection 3.1. The proof of Theorem D will be presented in Subsection 3.2.
Remark 1.1.
Theorem D is our result of most general applicability concerning linear response for skew-product maps. The assumption of harder verification being in general the differentiability of the map $\alpha\mapsto\bm{\nu}_{\alpha},$ even if only in the space $\mathbf{L}_{k}^{0}$. Recall that (A3) is used in Theorem B to obtain this conclusion in the space $\bm{\mathrm{E}}_{k}^{0}$, and for the general class of systems considered in Subsection 1.2, this differentiability follows indeed from Theorem B. However, this approach cannot be applied to the intermittent solenoid considered in Subsection 1.4.2, essentially because Theorem B requires the base dynamics to be at least of class $C^{2}$, a regularity assumption that is not satisfied by those systems.
As an alternative, in Subsection 1.3 we present an approach for skew-product systems whose base map admits an inducing scheme. This allows us to deduce the differentiability of the map $\alpha\mapsto\bm{\nu}_{\alpha}$ in the space $\mathbf{L}_{k}^{0}$ from the differentiability of the sample invariant measures of the induced skew-product in $\mathbf{E}_{k}^{0}$. For the induced system, this latter property can be established by applying Theorem B. We develop first in Subsection 1.2 a theory for skew-product maps whose base dynamics are given by piecewise expanding full-branch maps.
In Table 1 below, we summarise the section spaces and norms used in the analysis throughout this work; see Appendix A for the norms in the dual spaces used in the fibres.
Space Definition Metric
$\mathbf{P}$ $\{\bm{\nu}:\Omega\to\mathcal{M}_{1}(X)\mid\bm{\nu}\text{ is measurable}\}$ $d_{\mathrm{P}}(\bm{\mu},\bm{\nu})=\displaystyle\operatorname*{ess\,sup}_{\omega\in\Omega}\|\mu_{\omega}-\nu_{\omega}\|_{\mathrm{Lip}^{*}}$
$\mathbf{P}_{0}$ $\{\bm{\nu}:\Omega\to\mathcal{M}_{1}(X)\mid\bm{\nu}\text{ is $C^{0}$}\}$
$\mathbf{E}_{\mathrm{L}}^{0}$ $\{\bm{\nu}:\Omega\to\mathrm{Lip}_{0}(X)^{*}\mid\bm{\nu}\text{ is $C^{0}$}\}$ $\|\bm{\nu}\|_{\mathrm{L}}=\displaystyle\sup_{\omega\in\Omega}\|\nu_{\omega}\|_{\mathrm{Lip}^{*}}$
$\mathbf{E}_{\mathrm{L}}$ $\{\bm{\nu}:\Omega\to\mathrm{Lip}(X)^{*}\mid\bm{\nu}\text{ is $C^{0}$}\}$
$\mathbf{E}_{k}^{0}$ $\{\bm{\nu}:\Omega\to C^{k}_{0}(X)^{*}\mid\bm{\nu}\text{ is $C^{0}$}\}$ $\|\bm{\nu}\|_{k}=\displaystyle\sup_{\omega\in\Omega}\|\nu_{\omega}\|_{k^{*}}$
$\mathbf{E}_{k}$ $\{\bm{\nu}:\Omega\to C^{k}(X)^{*}\mid\bm{\nu}\text{ is $C^{0}$}\}$
$\mathbf{E}_{k}^{\prime}$ $\{\bm{\nu}:\Omega\to C^{k}(X)^{*}\mid\bm{\nu}\text{ is $C^{1}$}\}$ $\|\bm{\nu}\|_{k}^{\prime}=\|\bm{\nu}\|_{k}+\|\bm{\nu}^{\prime}\|_{k}$
$\mathbf{L}_{k}^{0}$ $\{\bm{\nu}:\Omega\to C^{k}_{0}(X)^{*}\mid\bm{\nu}\in L^{1}(\eta_{\alpha_{0}})\}$ $\|\bm{\nu}\|_{\mathbf{L}_{k}}=\int_{\Omega}\|\nu_{\omega}\|_{k^{*}}\,d\eta_{\alpha_{0}}(\omega)$
Table 1. Spaces of sections and metrics.
In the two applications of our results, we do not consider values of $k$ beyond k=3{k=3} for the spaces $\mathbf{E}_{k}^{0}$ and beyond $k=2$ for the spaces $\mathbf{E}_{k}$. The spaces $\mathbf{E}_{\mathrm{L}}$$\mathbf{E}_{k}$, and $\mathbf{E}_{k}^{\prime}$ have not appeared so far; they will be introduced later as auxiliary spaces in the statements and proofs of some of our results. The distinction between $\mathbf{E}_{\mathrm{L}}$$\mathbf{E}_{k}$ and $\mathbf{E}_{\mathrm{L}}^{0}$$\mathbf{E}_{k}^{0}$ is subtle, though essential: the latter consist of those sections whose values are, fibrewise, functionals vanishing on constant functions. This choice reflects two complementary requirements: on the one hand, the operator $\mathcal{K}_{\alpha}$ must act as a contraction on the spaces $\mathbf{E}_{\mathrm{L}}^{0}$$\mathbf{E}_{k}^{0}$; on the other hand, the spaces $\mathbf{E}_{k}$$\mathbf{E}_{k}^{\prime}$ must be sufficiently large to intersect $\mathbf{P}$ and to contain the fixed point of $\mathcal{K}_{\alpha}$ in $\mathbf{P}$. Finally, we emphasise that the use of the Lipschitz norm in the definition of the spaces $\mathbf{E}_{\mathrm{L}}^{0}$$\mathbf{E}_{\mathrm{L}}$, rather than the $C^{0}$ norm, is essential: the operator $\mathcal{K}_{\alpha}$ acts fibrewise as a contraction in the corresponding $\operatorname{Lip}_{0}(X)^{*}$ norm, a property which generally fails in the $C^{0}_{0}(X)^{*}$ setting.
1.2. Base maps with full branches
This section is devoted to skew-product systems whose base dynamics are given by piecewise smooth full branch maps. This setting will be particularly relevant for the applications presented in Section 1.4. Our goal is to introduce verifiable criteria for families of skew-product maps
$T_{\alpha}:\Omega\times X\to\Omega\times X,\qquad\alpha\in J,$
of the form (1.1), under which several of the assumptions of Section 1.1 can be checked more directly. For simplicity, we assume throughout that $\Omega$ is a compact interval in $\mathbb{R}$ and that $X$ is a compact, convex subset of $\mathbb{R}^{d}$ for some $d\geq 1$. While this framework admits natural generalisations, the present setting is sufficient for all the applications considered in this paper.
We assume there exists an at most countable family $\{\Omega_{\alpha,i}\}_{i\in\mathbb{I}}$ of smoothness domains for each base map $f_{\alpha}\colon\Omega\to\Omega$. We now introduce a collection of additional assumptions on the family $\{T_{\alpha}\}_{\alpha\in J}$.
  1. (B1)
    For each $i\in\mathbb{I}$ and $\alpha\in J$, the restriction $f_{\alpha}|_{\Omega_{\alpha,i}}\colon\Omega_{\alpha,i}\to\Omega$ is onto, and its local inverse branch has a $C^{1}$ extension $\theta_{\alpha,i}\colon\Omega\to\Omega_{\alpha,i}$ satisfying $|\theta_{\alpha,i}^{\prime}|\leq 1$.
  2. (B2)
    For each $i\in\mathbb{I}$, the map $\alpha\mapsto\theta_{\alpha,i}\in C^{1}(\Omega,\Omega)$ is continuous, and there exists $C_{\theta}>0$ such that
    $\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\big\|\theta_{\alpha,i}\big\|_{C^{1}}\leq C_{\theta}.$
  1. (B3)
    For each $\alpha\in J$ and $\omega\in\Omega$, we have $g_{\alpha,\omega}\in C^{3}(X)$. Moreover,
    1. (a)
      for each $i\in\mathbb{I}$, the map $\alpha\mapsto g_{\alpha,\theta_{\alpha,i}}\in C^{1}(\Omega\times X,X)$ is continuous;
    2. (b)
      there exists $\lambda<1$ such that defining for $i\in\mathbb{I}$, $\alpha\in J$ and $k=1,2,3$
      $L_{k}=L_{k}(\alpha,i)=\sup_{\omega\in\Omega}\sup_{x\in X}\|D^{k}g_{\alpha,\theta_{\alpha,i}(\omega)}(x)\|,$
      one has
      $\max\left\{L_{1}+L_{2}+L_{3},\,L_{1}^{2}+3L_{1}L_{2}\right\}\leq\lambda;$
    3. (c)
      there exists $C_{g}>0$ such that
      $\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\|g_{\alpha,\theta_{\alpha,i}}\|_{C^{1}}\leq C_{g}.$
For each $i\in\mathbb{I}$ and $\alpha\in J$, we introduce the weight function $p_{\alpha,i}$, defined for $\eta_{\alpha}$-almost every $\omega\in\Omega$ by
(1.11) $p_{\alpha,i}(\omega)=\frac{\rho_{\alpha}\bigl(\theta_{\alpha,i}(\omega)\bigr)}{|f_{\alpha}^{\prime}\bigl(\theta_{\alpha,i}(\omega)\bigr)|\,\rho_{\alpha}(\omega)}.$
With this notation, the sectional transfer operator $\mathcal{K}_{\alpha}$ defined in (1.3) admits the explicit representation
(1.12) $(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}=\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,\bigl(g_{\alpha,\theta_{\alpha,i}(\omega)}\bigr)_{*}\nu_{\alpha,\theta_{\alpha,i}(\omega)}.$
The reverse Markov kernel property translates in this setting to the simple relation
(1.13) $\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)=1.$
Our final assumption of this part concerns the continuity and summability properties of these weight functions. Throughout the text, we use a dot over a function to denote its partial derivative with respect to the parameter $\alpha\in J$.
  1. (B4)
    For each $i\in\mathbb{I}$, the map $\alpha\mapsto p_{\alpha,i}\in C^{1}(\Omega,\mathbb{R})$ is continuous, and there exists $C_{p}>0$ such that
    $\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\|p_{\alpha,i}\|_{C^{1}}\leq C_{p}\quad\text{and}\quad\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\|\dot{p}_{\alpha,i}\|_{C^{0}}\leq C_{p}.$
We refer the reader to Remark 4.2 for a detailed explanation of the role played by the constant $0<\lambda<1$ in (B3)(B3)(b), and for the motivation behind its choice. Note that if the maps $g_{\alpha,\omega}$ are affine then (B3)(B3)(b) means that $g_{\alpha,\omega}$ are uniform smooth contractions in $X$. Clearly (B3)(B3)(b) implies (A1) in this setting. The constant $\lambda<1$ is crucial in establishing the contraction properties of the sectional transfer operator on suitably chosen Banach spaces. In the present setting, we take $\mathbf{E}_{3}$ as the main reference space where the linear response of the sample measures is obtained.
Theorem E.
Assume that (A2) and (B1)(B4) hold. Then,
  1. (1)
    $\mathcal{K}_{\alpha}$ has a unique fixed point $\bm{\nu}_{\alpha}$ in $\mathbf{P}$;
  2. (2)
    $\bm{\nu}_{\alpha}\in\mathbf{P}_{0}$ and $\|\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}\|_{3}\to 0$ as $\alpha\to\alpha_{0};$
  3. (3)
    $\bm{\nu}_{\alpha}:\Omega\to C^{2}(X)^{*}$ is a $C^{1}$ function;
  4. (4)
    the map $\alpha\mapsto\bm{\nu}_{\alpha}$ is differentiable in $\mathbf{E}_{3}^{0}$ at $\alpha_{0}$.
Since condition (B3)(B3)(b) implies (A1), the existence and uniqueness of a fixed point for $\mathcal{K}_{\alpha}$ in $\mathbf{P}$ follow from Theorem A. The remaining conclusions are proved in Section 4. Specifically, the second conclusion follows from Corollary 4.6 and Corollary 4.12. The differentiability of $\bm{\nu}_{\alpha}$ is established in Proposition 4.16 as a by-product of the argument used to show that the assumptions of (A3) are satisfied, allowing Theorem B to be applied for the final conclusion.
Corollary F.
Assume that (A2), (A4) and (B1)(B4) hold and $\bm{\nu}_{\alpha}$ is the unique fixed point of $\mathcal{K}_{\alpha}$ in $\mathbf{P}$. Then, for every $C^{3}$ function $\Phi:\Omega\times X\to\mathbb{R}$, the map
$\alpha\longmapsto\int_{\Omega}\int_{X}\Phi(\omega,\cdot)d\nu_{\alpha,\omega}\,d\eta_{\alpha}(\omega)$
is differentiable at $\alpha_{0}$.
Since convergence in $\mathbf{E}_{3}^{0}$ implies convergence in $\mathbf{L}_{3}$, Corollary F follows from Theorems D and E once assumption (A5) has been verified for $k=3$. This verification will be carried out in Section 4, specifically in Corollaries 4.6 and 4.12. We conclude this part with a remark concerning the choice of the spaces $\mathbf{E}_{3}^{0}$ and $\mathbf{E}_{\mathrm{L}}^{0}$.
Remark 1.2.
In the present setting, to obtain the differentiability with respect to the parameter required by (A3)(A3)(d), there are two points where the $C^{2}$ regularity of the test functions is essential:
  1. (1)
    the differentiability of the invariant section (Proposition 4.16);
  2. (2)
    the continuity of the push-forward operator (Lemma 4.7).
In both cases, one cannot expect to improve the conclusions if the test functions are only assumed to be $C^{1}$; see Remark 4.10 and Remark 4.14.
However, since the first condition above is trivially satisfied when the invariant section is constant, and the second already provides the continuity required in (A3)(A3)(c) in the space $\mathbf{E}_{2}$, it is natural to ask whether differentiability of the invariant section could be obtained in $\mathbf{E}_{2}$, thereby strengthening the conclusion in the case where the section does not depend on the base points.
Even in this simplified scenario, this is not possible. Indeed, when verifying (A3)(A3)(d), Lemma 4.7 is applied in a crucial way to the gradient of a test function, which requires the test function to be of class $C^{3}$; see the estimates for the terms $C_{\varepsilon}$ and $D_{\varepsilon}$ in the proof of Proposition 4.18. Assuming the invariant section is constant has absolutely no effect on the derivation of these estimates.
Finally, to avoid exceeding $C^{2}$ regularity for the test functions, it is essential that the continuity in (A3)(A3)(c) is only required for elements in the space $\mathbf{E}_{\mathrm{L}}^{0}$. This allows the application of Lemma 4.7 while keeping the fibres at the level of $\operatorname{Lip}(X)^{*}$, ensuring continuity of the push-forward operator when tested against $C^{2}$ observables.
1.3. Base maps with inducing schemes
This section is largely motivated by Remark 1.1, and its purpose is to obtain the conclusion of Theorem B under milder assumptions, by indirectly applying Theorem B to an appropriate inducing scheme. As before, we consider a family of skew-product maps
$T_{\alpha}:\Omega\times X\to\Omega\times X,\qquad\alpha\in J,$
For many base maps $f_{\alpha}$, one can construct an inducing scheme, that is, there exist a subset $\bar{\Omega}\subset\Omega$ and a first return time function $\tau_{\alpha}:\bar{\Omega}\to\mathbb{N}$ such that the induced base map
$\bar{f}_{\alpha}:\bar{\Omega}\to\bar{\Omega}$
is defined for each $\omega\in\bar{\Omega}$ by
$\bar{f}_{\alpha}(\omega)=f_{\alpha}^{\tau_{\alpha}(\omega)}(\omega).$
  1. (C1)
    For each $\alpha\in J$, the maps $\bar{f}_{\alpha}$ satisfies:
    1. (a)
      there exists a countable $m$ modulo zero partition $\{\bar{\Omega}_{\alpha,i}\}_{i\in\mathbb{I}}$ of $\bar{\Omega}$ such that $\bar{f}_{\alpha}|_{\bar{\Omega}_{\alpha,i}}\colon\bar{\Omega}_{\alpha,i}\to\bar{\Omega}$} is a $C^{1}$ diffeomorphism for all $i\in\mathbb{I}$;
    2. (b)
      there exists $\sigma>1$ such that $|\bar{f}_{\alpha}^{\prime}|_{\bar{\Omega}_{\alpha,i}}|\geq\sigma$ for all $i\in\mathbb{I}$;
    3. (c)
      there exist $C,\gamma>0$ such that, for all $i\in\mathbb{I}$ and all $x,y\in\bar{\Omega}_{\alpha,i}$
      $\log\frac{\bar{f}^{\prime}_{\alpha}(x)}{\bar{f}_{\alpha}^{\prime}(y)}\leq C\left|\bar{f}_{\alpha}(x)-\bar{f}_{\alpha}(y)\right|^{\gamma}.$
    4. (d)
      $\tau_{\alpha}$ is integrable with respect to Lebesgue measure on $\bar{\Omega}$.
Under these conditions, the induced base map $\bar{f}_{\alpha}$ has a unique absolutely continuous invariant probability measure $\bar{\eta}_{\alpha}$ with a density $\bar{\rho}_{\alpha}\in L^{1}(\bar{\Omega})$. Moreover, an absolutely continuous invariant probability measure $\eta_{\alpha}$ for the original base map $f_{\alpha}$ can be obtained via the unfolding formula
(1.14) $\eta_{\alpha}=\frac{1}{E}\sum_{k=0}^{\infty}(f_{\alpha}^{k})_{*}\big(\bar{\eta}_{\alpha}|\{\tau_{\alpha}>k\}\big),\qquad E=\int_{\bar{\Omega}}\tau_{\alpha}\,d\bar{\eta}_{\alpha}.$
This classical procedure allows one to transfer many statistical properties from the induced map to the original system. The induced skew-product map is
(1.15) $\bar{T}_{\alpha}:\bar{\Omega}\times X\to\bar{\Omega}\times X,\qquad\bar{T}_{\alpha}(\omega,x)=(\bar{f}_{\alpha}(\omega),\bar{g}_{\alpha}(\omega,x)),$
with the induced fibre map
$\bar{g}_{\alpha}:\bar{\Omega}\times X\to X$
given for any $(\omega,x)\in\bar{\Omega}\times X$ by
$\bar{g}_{\alpha}(\omega,x)=g_{\alpha,f_{\alpha}^{\tau_{\alpha}(\omega)-1}(\omega)}\circ\cdots\circ g_{\alpha,f_{\alpha}(\omega)}\circ g_{\alpha,\omega}(x).$
Equivalently, $\bar{g}_{\alpha}(\omega,\cdot)=g_{\alpha,\omega}^{(\tau_{\alpha}(\omega))}$ with
(1.16) $g_{\alpha,\omega}^{(k)}=g_{\alpha,f_{\alpha}^{k-1}(\omega)}\circ\cdots\circ g_{\alpha,\omega}.$
Our goal is now to provide conditions under which some of the results of the previous sections can be applied, yielding linear response for the induced skew-product maps $\bar{T}_{\alpha}$ and, from this, deriving the linear response of the original skew-product maps $T_{\alpha}$. To implement this inducing strategy, we will work with certain Banach spaces from Table 1, both on the base space $\Omega$ and on the inducing domain $\bar{\Omega}$. The distinction between these two settings will be indicated by overlining the corresponding space for the induced system and keeping it not overlined for the original one.
Whenever needed, we regard a section $\bar{\bm{\nu}}$ on $\bar{\Omega}$ as a section on $\Omega$ by extending it by zero outside $\bar{\Omega}$. We define the sectional unfolding operator $\mathcal{U}_{\alpha}$ which constructs a full section on $\Omega$ from a section $\bar{\bm{\nu}}$ on the induced domain $\bar{\Omega}$ by
(1.17) $\mathcal{U}_{\alpha}(\bar{\bm{\nu}})=\sum_{k=0}^{\infty}\mathcal{K}_{\alpha}^{k}\bigl(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}}\bigr),$
where $\mathcal{K}_{\alpha}$ is the sectional transfer operator associated with $T_{\alpha}$ and $\mathbf{1}_{A}$ stands for the indicator function of any set $A\subset\Omega$. Note that the sum in (1.17) is pointwise finite, and therefore well defined as a section on $\Omega$.
The main result of this section will be obtained under conditions that, in particular, ensure the existence of a unique fixed point $\bar{\bm{\nu}}_{\alpha}$ in $\bar{\mathbf{P}}$ for the sectional transfer operator $\bar{\mathcal{K}}_{\alpha}$ associated with the induced system $\bar{T}_{\alpha}$. We require certain regularity properties of the unfolding operator, which we state next. As has been customary, these include conditions imposed both for generic parameter values and at a distinguished parameter $\alpha_{0}$.
  1. (C2)
    The unfolding operator $\mathcal{U}_{\alpha}$ defines a bounded operator from $\bar{\mathbf{E}}_{3}^{0}$ to $\mathbf{L}_{3}^{0}$. Moreover,
    1. (a)
      there exists $C>0$ such that $\|\mathcal{U}_{\alpha}\|\leq C$, for all $\alpha\in J$;
    2. (b)
      the map $\alpha\mapsto\mathcal{U}_{\alpha}(\bar{\bm{\nu}})$ is continuous in $\mathbf{L}_{3}^{0}$ at $\alpha_{0}$, for every $\bar{\bm{\nu}}\in\bar{\mathbf{E}}_{3}^{0}$;
    3. (c)
      the map $\alpha\mapsto\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha_{0}})$ is differentiable in $\mathbf{L}_{3}^{0}$ at $\alpha_{0}$.
We are now in a position to state our main result for skew-product systems whose base maps admits an inducing scheme as described above.
Theorem G.
Assume that (C1)(C2) hold and that (B1)(B4) hold for $\bar{T}_{\alpha}$. Then,
  1. (1)
    $\bar{\mathcal{K}}_{\alpha}$ has a unique fixed point $\bar{\bm{\nu}}_{\alpha}$ in $\bar{\mathbf{P}}$;
  2. (2)
    $\bm{\nu}_{\alpha}=\mathcal{U}_{\alpha}\bar{\bm{\nu}}_{\alpha}$ is a fixed point of $\mathcal{K}_{\alpha}$ in $\mathbf{P}$;
  3. (3)
    the map $\alpha\mapsto\bm{\nu}_{\alpha}$ is differentiable in $\mathbf{L}_{3}^{0}$.
This result will be proved in Section 5. We remark that the existence of a unique fixed point $\bar{\bm{\nu}}_{\alpha}$ in $\bar{\mathbf{P}}$ (in fact, in $\mathbf{P}_{0}$) for $\bar{\mathcal{K}}_{\alpha}$ is a direct consequence of Theorem E, since assumptions (B1)(B4) are assumed for $\bar{T}_{\alpha}$, and assumption (A2) for $\bar{T}_{\alpha}$ follows directly from (C1).
1.4. Applications
In this section, we apply our abstract framework to two landmark problems that serve as primary motivations for this work: the Bernoulli convolutions and the intermittent solenoid. These applications illustrate the power of our approach in handling two distinct but fundamental challenges in linear response theory.
1.4.1. Bernoulli convolutions
Bernoulli convolutions are a classical object at the interface of probability theory, harmonic analysis, and fractal geometry. They arise as probability measures on the real line generated by random power series of the form
$Y_{\alpha}=\sum_{n=0}^{\infty}a_{n}\alpha^{n},$
where $(a_{n})_{n\geq 0}$ are i.i.d. random variables with $\mathbb{P}(a_{n}=1)=\mathbb{P}(a_{n}=-1)=\beta$, and $0<\alpha,\beta<1$. We denote by $\nu_{\alpha,\beta}$ the distribution of $Y_{\alpha}$.
For the classical unbiased case $\beta=1/2$, a result of Jessen and Wintner [34] shows that $\nu_{\alpha,1/2}$ is always pure, that is, either absolutely continuous or singular with respect to Lebesgue measure. More precisely, for $\alpha<1/2$, the measure $\nu_{\alpha}$ is singular and supported on a Cantor set of zero Lebesgue measure. The regime $\alpha>1/2$ is considerably more subtle: Solomyak [43] proved that $\nu_{\alpha,1/2}$ is absolutely continuous for Lebesgue almost every $\alpha\in(1/2,1)$, whereas Erdős [29] showed that $\nu_{\alpha,1/2}$ is singular whenever $1/\alpha$ is a Pisot number. The connection of this unbiased case to deterministic dynamics is provided by the baker map; see e.g. [6].
Here we address the general biased case with $\beta\in(0,1)$ studied in [40] by Peres and Solomyak. Consider the family of skew-product map $T_{\alpha,\beta}\colon[-1,1]^{2}\to[-1,1]^{2}$ defined by
$T_{\alpha,\beta}(\omega,x)=(f_{\beta}(\omega),g_{\alpha}(\omega,x)),$
where the base map $f_{\beta}\colon[-1,1]\to[-1,1]$ is given by
$f_{\beta}(\omega)=\begin{cases}\frac{1}{\beta}\omega-\frac{\beta-1}{\beta},&\omega\in[-1,2\beta-1),\\[5.69054pt] \frac{1}{1-\beta}\omega-\frac{\beta-1}{\beta},&\omega\in[2\beta-1,1];\end{cases}$
and the fibre maps $g_{\alpha}\colon[-1,1]\times[-1,1]\to[-1,1]$, for $\alpha\in(0,1)$, are given by
$g_{\alpha}(\omega,x)=\begin{cases}\alpha x-(1-\alpha),&\omega\in[-1,2\beta-1),\\[5.69054pt] \alpha x+(1-\alpha),&\omega\in[2\beta-1,1].\end{cases}$
This system admits a unique physical measure $\mu_{\alpha,\beta}=m\times\nu_{\alpha,\beta}$, where $m$ denotes Lebesgue measure on the base. The vertical marginal $\nu_{\alpha,\beta}$ is the unique invariant probability measure of the iterated function system $\{g_{\alpha,1},g_{\alpha,2};\beta,1-\beta\}$ satisfying the relation
$\nu_{\alpha,\beta}=\tfrac{1}{\beta}(g_{\alpha,1})_{*}\nu_{\alpha,\beta}+\tfrac{1}{1-\beta}(g_{\alpha,2})_{*}\nu_{\alpha,\beta},$
where $g_{\alpha,1}=g_{\alpha}(\omega,\cdot)$ for $\omega<2\beta-1$ and $g_{\alpha,2}=g_{\alpha}(\omega,\cdot)$ for $\omega\geq 2\beta-1$. Equivalently, $\nu_{\alpha,\beta}$ is the distribution of the random series
$X_{\alpha}=(1-\alpha)\sum_{n=0}^{\infty}a_{n}\alpha^{n},$
where $(a_{n})_{n\geq 0}$ are i.i.d. random variables with $\mathbb{P}(a_{n}=1)=\mathbb{P}(a_{n}=-1)=\beta$; this corresponds to a simple rescaling of the original Bernoulli convolution.
Kloeckner [37] proved that in the unbiased case the map $\alpha\mapsto\nu_{\alpha,1/2}$ is differentiable for almost every $\alpha\in(1/2,1)$ – more precisely, for a full-measure subset of the parameters considered in [43] – when tested against smooth observables, a result usually referred to as weak linear response. Recently we were made aware that in the upcoming paper [31] by Fu, the author studies weak linear response for unbiased Bernoulli convolutions against Hölder observables. In particular, it is shown that whenever the Hausdorff dimension of $\nu_{\alpha,1/2}$ is smaller than one, weak differentiability against Hölder observables fails to hold, while for $C^{1}$ observables weak linear response always hold. By contrast, the framework developed in this paper allows us to obtain linear response in a strong sense, valid for all parameters $\alpha,\beta\in(0,1)$.
Theorem H.
The probability measure $\nu_{\alpha,\beta}$ depends differentiably on $\alpha,\beta\in(0,1)$ with values in $C^{3}([-1,1])^{*}$.
Theorem H follows as a direct application of the final item of Theorem E. For that, we are going to consider the dependence on the parameters $\alpha,\beta$ separately. Nevertheless, we use the sectional transfer operator in this setting labelled by both parameters
$(\mathcal{K}_{\alpha,\beta}\bm{\nu})_{\omega}=\tfrac{1}{\beta}(g_{\alpha,1})_{*}\nu_{\theta_{\alpha,\beta,1}(\omega)}+\tfrac{1}{1-\beta}(g_{\alpha,2})_{*}\nu_{\theta_{\alpha,\beta,2}(\omega)}.$
Since $\mathcal{K}_{\alpha,\beta}$ preserves the closed subspace of constant sections of $\mathbf{P}$, its fixed point $\bm{\nu}_{\alpha,\beta}$ is necessarily a constant section. Consequently, the space $\mathbf{E}_{3}$ in which differentiability is obtained can naturally be identified with the dual space $C^{3}([-1,1])^{*}$. It therefore suffices to verify the assumptions of Theorem E.
(A2), (B1), (B2): These assumptions are trivially satisfied: the base map is $f_{\beta}$ for all $\alpha,\beta$, with two inverse branches taking values in $1,2\beta-1)$ and $2\beta-1,1]$, and its absolutely continuous invariant probability measure is Lebesgue measure.
(B3): Since both fibre maps are affine contractions, all the required conditions are immediately verified. The only subtlety concerns the contraction constant $\lambda=\alpha$, which cannot be chosen uniformly for all $\alpha\in(0,1)$. However, as differentiability is a local property in parameter space, $\lambda$ can always be chosen locally uniformly.
(B4): This assumption is trivially verified, as the weight functions are constant on each branch with values $\beta$ and $1-\beta$.
1.4.2. Solenoid with intermittency
We consider a class of skew-product systems that are partially hyperbolic but not uniformly hyperbolic, due to the presence of a neutral fixed point in the base dynamics. Such examples arise naturally in the study of intermittent behaviour and provide a benchmark for statistical stability beyond the uniformly hyperbolic setting. This example differs from the classical construction of the hyperbolic solenoidal map as a skew-product in that we replace the doubling map in the base with an intermittent LSV (Liverani-Saussol-Vaienti) map [39]. Linear response for LSV maps was studied in [8, 18, 38] and in a random and sequential setting in [10, 28].
Let $D\subset\mathbb{R}^{2}$ denote the unit disk and fix $\lambda\in(0,1/2)$. For each $\alpha\in(0,1)$, consider the skew-product map
$T_{\alpha}\colon[0,1]\times D\longrightarrow[0,1]\times D,\qquad T_{\alpha}(\omega,x)=(f_{\alpha}(\omega),g(\omega,x),$
where the base map $f_{\alpha}\colon[0,1]\to[0,1]$ is given by
$f_{\alpha}(\omega)=\begin{cases}\omega\bigl(1+2^{\alpha}\omega^{\alpha}\bigr),&\omega\in[0,\tfrac{1}{2}],\\[5.69046pt] 2\omega-1,&\omega\in(\tfrac{1}{2},1],\end{cases}$
and the fibre maps are defined for $(\omega,x)\in[0,1]\times D$ by
$g(\omega,x)=\Bigl(\tfrac{1}{2}\cos(2\pi\omega),\,\tfrac{1}{2}\sin(2\pi\omega)\Bigr)+\lambda x.$
The map $T_{\alpha}$ admits a unique physical measure $\mu_{\alpha}$ whose basin covers Lebesgue-almost every point in $0,1]\times D$; see [1, Theorem 4.28]. The dynamics is partially hyperbolic, with uniform contraction along the fibres and non-uniform expansion in the base direction, where a neutral fixed point gives rise to intermittent behaviour. This lack of uniform hyperbolicity is precisely the mechanism responsible for the polynomial decay of correlations established in [4]. The statistical properties of this system fit naturally into the induced framework developed in this paper.
Theorem I.
For every $C^{3}$ observable $\Phi\colon[0,1]\times D\to\mathbb{R}$, the map
$\alpha\longmapsto\int_{[0,1]\times D}\Phi\,d\mu_{\alpha}$
is differentiable at every parameter $\alpha_{0}\in(0,1)$.
This follows as an application of our previous results in the following way. First, note that the system satisfies the assumptions of Theorem A, which guarantees the existence of a skew-product measure as in the conclusion of that theorem, with marginal $\eta_{\alpha}$ on $\Omega$. Since $\eta_{\alpha}$ is a physical measure for the base map, it follows from [36, Theorem A] that this skew-product measure is a physical measure for $T_{\alpha}$, and thus coincides with $\mu_{\alpha}$ by uniqueness. The conclusion of Thoerem I then follows from Theorem D applied with $k=3$, together with Theorem G. Recall that the applicability of Theorem D relies on the final conclusion of Theorem G as well as on assumptions (A1)(A2) and (A4)(A5), which we now proceed to verify.
(A1): Follows trivially from the definition of the fibre maps.
(A2): Is well known in the literature; see, for instance, [1, Section 3.5.2].
(A4): Is established in [8, Theorem 3.1].
(A5): The first part follows from Corollary 4.6, while the second part is a consequence of Corollary 4.12, both under assumptions (B1)(B2) and a weaker version (B4’) of assumption (B4) that we state in the beginning of Subsection 4.2, which we address next.
(B1)(B2): These properties are immediate, since the base maps $f_{\alpha}$ satisfy $f^{\prime}_{\alpha}\geq 1$ and admit only two inverse branches, which can easily be seen to depend continuously on the parameter in the $C^{1}$ topology.
(B4’): In this case, we have two weight functions
$p_{\alpha,i}(\omega)=\frac{\rho_{\alpha}\bigl(\theta_{\alpha,i}(\omega)\bigr)}{|f_{\alpha}^{\prime}\bigl(\theta_{\alpha,i}(\omega)\bigr)|\,\rho_{\alpha}(\omega)},\quad i=1,2.$
It suffices to show that these quantities depend continuously on $\alpha$ in the $C^{0}$ norm. It is straightforward to verify that both the inverse branches $\theta_{\alpha,i}$ and the derivative $f_{\alpha}^{\prime}$ depend continuously on $\alpha$. Assume that $p_{\alpha,1}$ corresponds to the right-hand branch, which is uniformly expanding. It follows from (1.14) that, up to a normalising constant, $\rho_{\alpha}\circ\theta_{\alpha,1}$ coincides with $\bar{\rho}_{\alpha}\circ\theta_{\alpha,1}$. This density is bounded away from zero and infinity and depends continuously (in fact, smoothly) on the parameter $\alpha$, by [8, Theorem 3.2]. On the other hand, by the results of [46, Section 6.3], there exists a constant $C>0$ such that
(1.18) $C^{-1}\omega^{-\alpha}\leq\rho_{\alpha}(\omega)\leq C\,\omega^{-\alpha}.$
Combining this estimate with [8, Theorem 3.1], we deduce that $1/\rho_{\alpha}$ depends continuously on $\alpha$ in the $C^{0}$ norm. Consequently, $p_{\alpha,1}$ depends continuously on $\alpha$ in the $C^{0}$ norm. Finally, since $p_{\alpha,0}+p_{\alpha,1}=1$ by (1.13), the same continuity property holds for $p_{\alpha,0}$.
The verification of the existence of an inducing scheme for which (B1)(B4) (for the induced map) and (C1)(C2) hold is deferred to Subsection 5.3, thereby ensuring the applicability of Theorem G.
2. Invariant measures
In this section, we establish the conclusions of Theorem A under assumptions (A1)(A2). We proceed in two main steps. In Subsection 2.1, we show that the operator $\mathcal{K}_{\alpha}$ is a contraction on the complete metric space $\mathbf{P}$, which guarantees the existence of a unique fixed point $\bm{\nu}_{\alpha}\in\mathbf{P}$. In Subsection 2.2, we use this fixed point to construct the corresponding skew-product measure and prove its invariance under the dynamics of $T_{\alpha}$.
2.1. Sample measures
We consider the space $\mathcal{M}_{1}(X)$ of Borel probability measures on $X$, equipped with the Wasserstein–$1$ distance, induced by the norm on $\mathrm{Lip}(X)^{*}$, given for $\mu,\nu\in\mathcal{M}_{1}(X)$ by
$W_{1}(\mu,\nu)=\sup_{\mathrm{Lip}(\phi)\leq 1}\left\{\int\phi\,d\mu-\int\phi\,d\nu\right\}.$
see the Appendix A for details.
We start by recording the standard fact that the push-forward of probability measures by a Lipschitz map contracts distances in the Wasserstein–$1$ metric by at most the Lipschitz constant.
Lemma 2.1.
If $h:X\to X$ has $\operatorname{Lip}(h)\leq L$, then for any $\mu,\nu\in\mathcal{M}_{1}(X)$,
$W_{1}(h_{*}\nu,\;h_{*}\mu)\leq L\,W_{1}(\nu,\mu).$
Proof.
Let $h:X\to X$ satisfy $\mathrm{Lip}(h)\leq L$. For any $\mu,\nu\in\mathcal{M}_{1}(X)$, we have
$W_{1}(\nu,\mu)=\sup_{\mathrm{Lip}(\phi)\leq 1}\left|\int_{X}\phi\,\nu-\int_{X}\phi\,d\mu\right|.$
Also,
$W_{1}(h_{*}\nu,\;h_{*}\mu)$ $=\sup_{\mathrm{Lip}(\phi)\leq 1}\left|\int_{X}\phi\,d(h_{*}\nu)-\int_{X}\phi\,d(h_{*}\mu)\right|$
$=\sup_{\mathrm{Lip}(\phi)\leq 1}\left|\int_{X}\phi\circ h\,\nu-\int_{X}\phi\circ h\,d\mu\right|.$
Since $\mathrm{Lip}(\phi\circ h)\leq\mathrm{Lip}(\phi)\,\mathrm{Lip}(h)\leq L$, we have
$W_{1}(h_{*}\nu,\;h_{*}\mu)\leq\sup_{\mathrm{Lip}(\psi)\leq L}\left|\int_{X}\psi\,\nu-\int_{X}\psi\,d\mu\right|.$
Rescaling $\phi=\psi/L$, we have $\mathrm{Lip}(\phi)\leq 1$, which gives
$\sup_{\mathrm{Lip}(\psi)\leq L}\left|\int_{X}\psi\,d\nu-\int_{X}\psi\,d\mu\right|=L\sup_{\mathrm{Lip}(\phi)\leq 1}\left|\int_{X}\phi\,d\nu-\int_{X}\phi\,d\mu\right|.$
Thus,
$W_{1}(h_{*}\nu,\;h_{*}\mu)\leq L\,W_{1}(\nu,\mu),$
which completes the proof. ∎
We equip $\mathbf{P}$ with the supremum metric, defined for any $\bm{\mu},\bm{\nu}\in\mathbf{P}$ by
$d_{\mathrm{P}}(\bm{\mu},\bm{\nu})=\operatorname*{ess\,sup}_{\omega\in\Omega}W_{1}(\mu_{\omega},\nu_{\omega}).$
Since $(\Omega,d_{\Omega})$ is a compact metric space and $(\mathcal{M}_{1}(X),W_{1})$ is complete, it follows that $d_{\mathrm{P}}$ is a well-defined metric on $\mathbf{P}$ and that $(\mathbf{P},d_{\mathrm{P}})$ is complete. Moreover, the measurability of the functions involved in the definition of $\mathcal{K}_{\alpha}$ ensures that $\mathcal{K}_{\alpha}$ maps $\mathbf{P}$ into itself.
Lemma 2.2.
Let $0<\lambda<1$ be the constant given by (A1). For all $\bm{\mu},\bm{\nu}\in\mathbf{P}$, we have
$d_{\mathrm{P}}({\mathcal{K}_{\alpha}}\bm{\nu},{\mathcal{K}_{\alpha}}\bm{\mu})\leq\lambda d_{\mathrm{P}}(\bm{\nu},\bm{\mu}).$
Proof.
Given $\bm{\mu},\bm{\nu}\in\mathbf{P}$, by convexity of $W_{1}$ (recall it is induced by the norm in $\operatorname{Lip}(X)^{*}$),
(2.1) $W_{1}\big(({\mathcal{K}_{\alpha}}\bm{\nu})_{\omega},({\mathcal{K}_{\alpha}}\bm{\mu})_{\omega}\big)\leq\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|\rho_{\alpha}(\omega)}W_{1}((g_{\alpha,\theta})_{*}\nu_{\theta},(g_{\alpha,\theta})_{*}\mu_{\theta}).$
Using assumption (A1) and Lemma 2.1, we obtain for each $\theta$
$W_{1}\left((g_{\alpha,\theta})_{*}\nu_{\theta},(g_{\alpha,\theta})_{*}{\mu}_{\theta}\right)\leq\lambda\,W_{1}(\nu_{\theta},\mu_{\theta}),$
which together with (2.1) and (2.1), yields
$W_{1}\big(({\mathcal{K}}\bm{\nu})_{\omega},\,({\mathcal{K}}\bm{\mu})_{\omega}\big)$ $\leq\lambda\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|\rho_{\alpha}(\omega)}W_{1}(\nu_{\theta},\mu_{\theta})$
(2.2) $\leq\lambda\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|\rho_{\alpha}(\omega)}d_{\mathrm{P}}(\bm{\nu},\bm{\mu})$
Now observe that the density $\rho_{\alpha}$ of the invariant measure $\eta_{\alpha}$ satisfies the Perron–Frobenius equation
$\rho_{\alpha}(\omega)=\sum_{\theta\in f_{\alpha}^{-1}(\omega)}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|}.$
which together with (2.2) yields
$W_{1}\big(({\mathcal{K}}\bm{\nu})_{\omega},\,({\mathcal{K}}\bm{\mu})_{\omega}\big)\leq\lambda\,d_{\mathrm{P}}(\bm{\nu},\bm{\mu}).$
Taking the essential supremum over $\omega\in\Omega$, we get
$d_{\mathrm{P}}\big({\mathcal{K}_{\alpha}}\bm{\nu},\,{\mathcal{K}_{\alpha}}\bm{\mu}\big)\leq d_{\mathrm{P}}(\bm{\nu},\bm{\mu}),$
thus completing the proof. ∎
Since $(\mathbf{P},d_{\mathrm{P}})$ is a complete metric space, Banach fixed point theorem and the previous lemma give the first conclusion of Theorem A.
Corollary 2.3.
$\mathcal{K}_{\alpha}$ has a unique fixed point in $\mathbf{P}$.
2.2. Skew-product measure
Our goal here is to deduce the second conclusion of Theorem A. Given any $\bm{\nu}\in\mathbf{P}$ and any bounded measurable function $\Phi:\Omega\times X\to\mathbb{R}$, we define the pairing
$\langle\!\langle\bm{\nu},\Phi\rangle\!\rangle=\int_{\Omega}\langle\nu_{\omega},\Phi(\omega,\cdot)\rangle\,d\eta_{\alpha}(\omega).$
Lemma 2.4.
For any $\bm{\nu}\in\mathbf{P}$ and any bounded measurable function $\Phi:\Omega\times X\to\mathbb{R}$, one has
$\langle\!\langle\mathcal{K}_{\alpha}\bm{\nu},\Phi\rangle\!\rangle=\langle\!\langle\bm{\nu},\Phi\circ T_{\alpha}\rangle\!\rangle.$
Proof.
By definition of the pairing and of $\mathcal{K}_{\alpha}$, and using that $d\eta_{\alpha}=\rho_{\alpha}\,dm$, we compute
$\langle\!\langle\mathcal{K}_{\alpha}\bm{\nu},\Phi\rangle\!\rangle$ $=\int_{\Omega}\langle(\mathcal{K}_{\alpha}\bm{\nu})_{\omega},\Phi(\omega,\cdot)\rangle\,\rho_{\alpha}(\omega)\,dm(\omega)$
$=\int_{\Omega}\sum_{\theta\in f_{\alpha}^{-1}(\omega)}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|}\,\langle(g_{\alpha,\theta})_{*}\nu_{\theta},\Phi(\omega,\cdot)\rangle\,dm(\omega)$
$=\int_{\Omega}\sum_{\theta\in f_{\alpha}^{-1}(\omega)}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|}\,\langle\nu_{\theta},\Phi(\omega,g_{\alpha,\theta}(\cdot))\rangle\,dm(\omega).$
The absolute convergence of the series allows us to exchange summation and integration expressed in the duality. By (A2), there exists a countable family $\{\Omega_{i}\}_{i\in\mathbb{I}}$ of measurable subsets of $\Omega$, covering $\Omega$ up to an $m$-null set, such that each restriction $f_{\alpha}|_{\Omega_{i}}:\Omega_{i}\to f_{\alpha}(\Omega_{i})$ is a $C^{1}$ diffeomorphism onto its image and $\det Df_{\alpha}(\theta)\neq 0$ for all $\theta\in\Omega_{i}$. We now perform a change of variables on each inverse branch
$dm(\omega)=|\det Df_{\alpha}(\theta)|\,dm(\theta).$
Therefore,
$\int_{\Omega}\sum_{\theta\in f_{\alpha}^{-1}(\omega)}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|}\,$ $\langle\nu_{\theta},\Phi(\omega,g_{\alpha,\theta}(\cdot))\rangle\,dm(\omega)$
$=\sum_{i\in\mathbb{I}}\int_{\Omega_{i}}\rho_{\alpha}(\theta)\,\langle\nu_{\theta},\Phi(f_{\alpha}(\theta),g_{\alpha,\theta}(\cdot))\rangle\,dm(\theta)$
$=\int_{\Omega}\langle\nu_{\theta},\Phi(f_{\alpha}(\theta),g_{\alpha,\theta}(\cdot))\rangle\,d\eta_{\alpha}(\theta).$
By definition of the skew-product map $T_{\alpha}$, this equals
$\langle\!\langle\bm{\nu},\Phi\circ T_{\alpha}\rangle\!\rangle,$
which concludes the proof. ∎
Corollary 2.5.
If $\bm{\nu}_{\alpha}\in\mathbf{P}$ is a fixed point of the operator $\mathcal{K}_{\alpha}$, then
$\mu_{\alpha}=\int_{\Omega}\nu_{\alpha,\omega}\,d\eta_{\alpha}(\omega)$
is a $T_{\alpha}$-invariant probability measure.
Proof.
Let $\Phi:\Omega\times X\to\mathbb{R}$ be an arbitrary bounded measurable function. We have
$\int_{\Omega\times X}\Phi\circ T_{\alpha}\,d\mu_{\alpha}=\langle\!\langle\bm{\nu}_{\alpha},\Phi\circ T_{\alpha}\rangle\!\rangle=\langle\!\langle\mathcal{K}_{\alpha}\bm{\nu}_{\alpha},\Phi\rangle\!\rangle=\langle\!\langle\bm{\nu}_{\alpha},\Phi\rangle\!\rangle=\int_{\Omega\times X}\Phi\,d\mu_{\alpha},$
where the first and fourth equalities follow from the definition of the pairing, the second equality uses the duality given by Lemma 2.4, and the third uses that $\bm{\nu}_{\alpha}$ is a fixed point of $\mathcal{K}_{\alpha}$. Hence, $(T_{\alpha})_{*}\mu_{\alpha}=\mu_{\alpha}$, showing that $\mu_{\alpha}$ is $T_{\alpha}$-invariant. ∎
3. Linear response
Here we consider a family of skew-product maps $T_{\alpha}:\Omega\times X\to\Omega\times X$, with $\alpha\in J$. Our goal in the next two subsections is to prove Theorem B and Theorem D.
3.1. Sample measures
Here we prove Theorem B. Assume that $\mathcal{K}_{\alpha}$ defines a bounded linear operator on a Banach space $\mathbf{E}_{k}^{0}$, for some $k\geq 0$, and (A1)(A3) hold. Let $\bm{\nu}_{\alpha}$ the unique fixed point of $\mathcal{K}_{\alpha}$ in $\mathbf{P}$. We need to show that there exists some $\dot{\bm{\nu}}_{\alpha_{0}}\in\mathbf{E}_{k}^{0}$ such that
(3.1) $\lim_{\alpha\to\alpha_{0}}\left\|\frac{\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}}{\alpha-\alpha_{0}}-\dot{\bm{\nu}}_{\alpha_{0}}\right\|_{k}=0.$
Since ${\mathcal{K}}_{\alpha}\bm{\nu}_{\alpha}=\bm{\nu}_{\alpha}$ and ${\mathcal{K}}_{\alpha_{0}}\bm{\nu}_{\alpha_{0}}=\bm{\nu}_{\alpha_{0}}$, we may write
$(I-{\mathcal{K}}_{\alpha})(\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}})=({\mathcal{K}}_{\alpha}-{\mathcal{K}}_{\alpha_{0}})\bm{\nu}_{\alpha_{0}}.$
By assumption (A3)(A3)(a) and a standard result on Banach spaces, we have that $I-\mathcal{K}_{\alpha}$ is invertible. Moreover, the inverse is given by the von Neumann series
(3.2) $(I-\mathcal{K}_{\alpha})^{-1}=\sum_{n=0}^{\infty}\mathcal{K}_{\alpha}^{n}.$
Hence,
(3.3) $\frac{\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}}{\alpha-\alpha_{0}}={(I-{\mathcal{K}}_{\alpha})^{-1}}{\frac{{\mathcal{K}}_{\alpha}-{\mathcal{K}}_{\alpha_{0}}}{\alpha-\alpha_{0}}}\bm{\nu}_{\alpha_{0}}.$
Define
(3.4) $\bm{\tau}_{\alpha}={\frac{{\mathcal{K}}_{\alpha}-{\mathcal{K}}_{\alpha_{0}}}{\alpha-\alpha_{0}}}\bm{\nu}_{\alpha_{0}}.$
It follows from assumption (A3)(A3)(b) that $\bm{\tau}_{\alpha}\in\mathbf{E}_{\mathrm{L}}^{0}$ and by assumption (A3)(A3)(d) there exists some $\bm{\tau}_{\alpha_{0}}\in\mathbf{E}_{k}^{0}$ such that
(3.5) $\lim_{\alpha\to\alpha_{0}}\left\|\bm{\tau}_{\alpha}-{\bm{\tau}}_{\alpha_{0}}\right\|_{k}=0.$
Combining (3.3) and (3.4), it remains to prove that $(I-\mathcal{K}_{\alpha})^{-1}\bm{\tau}_{\alpha}$ converges in $\mathbf{E}_{k}^{0}$, when $\alpha\to\alpha_{0}$. In fact, we will show that it converges to $(I-{\mathcal{K}}_{\alpha_{0}})^{-1}\bm{\tau}_{\alpha_{0}}$. Writing
$\|(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha}-(I-{\mathcal{K}}_{\alpha_{0}})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k}\,\leq\;$ $\|(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha}-(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k}$
$+\|(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha_{0}}-(I-{\mathcal{K}}_{\alpha_{0}})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k},$
we are left to prove that
  1. (i)
    $\|(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha}-(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k}\to 0$, as $\alpha\to\alpha_{0}$;
  2. (ii)
    $\|(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha_{0}}-(I-{\mathcal{K}}_{\alpha_{0}})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k}\to 0$, as $\alpha\to\alpha_{0}$.
We first prove (i). We have
(3.6) $\|(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha}-(I-{\mathcal{K}}_{\alpha})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k}\leq\|(I-{\mathcal{K}}_{\alpha})^{-1}\|\cdot\|\bm{\tau}_{\alpha}-{\bm{\tau}}_{\alpha_{0}}\|_{k}.$
It follows from (A3)(A3)(a) and (3.2) that
(3.7) $\|(I-\mathcal{K}_{\alpha})^{-1}\|\leq\sum_{n=0}^{\infty}\|\mathcal{K}_{\alpha}\|^{n}\leq\sum_{n=0}^{\infty}\lambda^{n}=\frac{1}{1-\lambda}.$
Using (3.5), (3.6) and (3.7), we obtain (i).
To prove (ii), we use the second resolvent identity
$(I-{\mathcal{K}}_{\alpha})^{-1}-(I-{\mathcal{K}}_{\alpha_{0}})^{-1}=(I-\mathcal{K}_{\alpha})^{-1}(\mathcal{K}_{\alpha}-\mathcal{K}_{\alpha_{0}})(I-\mathcal{K}_{\alpha_{0}})^{-1},$
which combined with (3.7) yields
(3.8) $\|(I-\mathcal{K}_{\alpha})^{-1}\bm{\tau}_{\alpha_{0}}-(I-\mathcal{K}_{\alpha_{0}})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k}\leq\frac{1}{1-\lambda}\|(\mathcal{K}_{\alpha}-\mathcal{K}_{\alpha_{0}})(I-\mathcal{K}_{\alpha_{0}})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k}.$
Set
$\bm{\tau}_{0}=(I-\mathcal{K}_{\alpha_{0}})^{-1}\bm{\tau}_{\alpha_{0}}.$
By (3.5), the vector $\bm{\tau}_{\alpha_{0}}$ is the limit in $\mathbf{E}_{k}^{0}$ of $\bm{\tau}_{\alpha}\in\mathbf{E}_{\mathrm{L}}^{0}$, as $\alpha\to\alpha_{0}$. Thus, $\bm{\tau}_{\alpha_{0}}$ belongs to the closure $\overline{\mathbf{E}}_{\mathrm{L}}^{0}$ of $\mathbf{E}_{\mathrm{L}}^{0}$ in the topology induced by the norm in $\mathbf{E}_{k}^{0}$. Moreover, assumption (A3)(A3)(b) gives that $\mathcal{K}_{\alpha_{0}}$ maps $\mathbf{E}_{\mathrm{L}}^{0}$ into itself. Since we assume $\mathcal{K}_{\alpha_{0}}$ a bounded operator on $\mathbf{E}_{k}^{0}$, then $\mathcal{K}_{\alpha_{0}}$ maps $\overline{\mathbf{E}}_{\mathrm{L}}^{0}$ into itself. Consequently, the series in (3.2) defining $(I-\mathcal{K}_{\alpha_{0}})^{-1}$ also maps $\overline{\mathbf{E}}_{\mathrm{L}}^{0}$ into itself. This ensures that $\bm{\tau}_{0}\in\overline{\mathbf{E}}_{\mathrm{L}}^{0}$.
Now, we fix an arbitrary $\varepsilon>0$. By definition of the closure, there exists an approximating element $\bm{\tau}_{\varepsilon}\in\mathbf{E}_{\mathrm{L}}^{0}$ such that
(3.9) $\|\bm{\tau}_{0}-\bm{\tau}_{\varepsilon}\|_{k}<\frac{(1-\lambda)\varepsilon}{3\lambda}.$
We now bound the right-hand side of (3.8) using the triangle inequality:
(3.10) $\|(\mathcal{K}_{\alpha}-\mathcal{K}_{\alpha_{0}})\bm{\tau}_{0}\|_{k}\leq\|\mathcal{K}_{\alpha}(\bm{\tau}_{0}-\bm{\tau}_{\varepsilon})\|_{k}+\|(\mathcal{K}_{\alpha}-\mathcal{K}_{\alpha_{0}})\bm{\tau}_{\varepsilon}\|_{k}+\|\mathcal{K}_{\alpha_{0}}(\bm{\tau}_{\varepsilon}-\bm{\tau}_{0})\|_{k}.$
By assumption (A3)(A3)(a) and the choice of $\bm{\tau}_{\varepsilon}$ in (3.9), the first and third terms are bounded as
(3.11) $\|\mathcal{K}_{\alpha}(\bm{\tau}_{0}-\bm{\tau}_{\varepsilon})\|_{k}\leq\lambda\|\bm{\tau}_{0}-\bm{\tau}_{\varepsilon}\|_{k}<\frac{(1-\lambda)\varepsilon}{3},$
and
(3.12) $\|\mathcal{K}_{\alpha_{0}}(\bm{\tau}_{\varepsilon}-\bm{\tau}_{0})\|_{k}\leq\lambda\|\bm{\tau}_{\varepsilon}-\bm{\tau}_{0}\|_{k}<\frac{(1-\lambda)\varepsilon}{3}.$
Since $\bm{\tau}_{\varepsilon}\in\mathbf{E}_{\mathrm{L}}^{0}$, we invoke assumption (A3)(A3)(c), which provides a $\delta>0$ such that if $|\alpha-\alpha_{0}|<\delta$, then
(3.13) $\|(\mathcal{K}_{\alpha}-\mathcal{K}_{\alpha_{0}})\bm{\tau}_{\varepsilon}\|_{k}<\frac{(1-\lambda)\varepsilon}{3}.$
Substituting (3.11), (3.12), and (3.13) into (3.10), we obtain
$\|(\mathcal{K}_{\alpha}-\mathcal{K}_{\alpha_{0}})\bm{\tau}_{0}\|_{k}<(1-\lambda)\varepsilon.$
Finally, substituting this back into (3.8), we have for $|\alpha-\alpha_{0}|<\delta$
$\|(I-\mathcal{K}_{\alpha})^{-1}\bm{\tau}_{\alpha_{0}}-(I-\mathcal{K}_{\alpha_{0}})^{-1}\bm{\tau}_{\alpha_{0}}\|_{k}\leq\frac{1}{1-\lambda}\cdot(1-\lambda)\varepsilon=\varepsilon.$
Since $\varepsilon>0$ was arbitrary, we obtain (ii), thereby completing the proof of Theorem B.
3.2. Skew-product measure
Here we prove Theorem D. Let $\bm{\nu}_{\alpha}$ be the unique fixed point of $\mathcal{K}_{\alpha}$ in $\mathbf{P}_{0}$ and
$\mu_{\alpha}=\int_{\Omega}\nu_{\alpha,\omega}\,d\eta_{\alpha}(\omega)$
be the measure given by Theorem A. Assume that (A4)(A5) hold and the map $\alpha\mapsto\bm{\nu}_{\alpha}$ is differentiable in the space $\mathbf{L}_{k}^{0}$ at $\alpha_{0}$, for some $k\geq 1$. Given a $C^{k}$ observable $\Phi:\Omega\times X\to\mathbb{R}$, define for each $\alpha\in J$
$\psi_{\alpha}(\omega)=\int_{X}\Phi(\omega,x)\,d\nu_{\alpha,\omega}(x),$
so that
$\int_{\Omega\times X}\Phi\,d\mu_{\alpha}=\int_{\Omega}\psi_{\alpha}\,d\eta_{\alpha}.$
Lemma 3.1.
Each $\psi_{\alpha}:\Omega\to\mathbb{R}$ is a continuous function; moreover, $\|\psi_{\alpha}-\psi_{\alpha_{0}}\|_{C^{0}}\to 0$, as $\alpha\to\alpha_{0}$.
Proof.
First we show that $\psi_{\alpha}$ is a continuous function. For all $\omega,\omega^{\prime}\in\Omega$, we have
$|\psi_{\alpha}(\omega)-\psi_{\alpha}(\omega^{\prime})|$ $=\left|\int_{X}\Phi(\omega,\cdot)\,d\nu_{\alpha,\omega}-\int_{X}\Phi(\omega^{\prime},\cdot)\,d\nu_{\alpha,\omega}\right|$
$\leq\int_{X}|\Phi(\omega,\cdot)-\Phi(\omega^{\prime},\cdot)|\,d\nu_{\alpha,\omega}+\left|\int_{X}\Phi(\omega^{\prime},\cdot)\,d\nu_{\alpha,\omega}-\int_{X}\Phi(\omega^{\prime},\cdot)\,d\nu_{\alpha,\omega^{\prime}}\right|.$
Let $L=\|\Phi\|_{C^{k}}$. Since $L$ works as a Lipschitz constant for the map $\omega\mapsto\Phi(\omega,\cdot)$ and each $\nu_{\alpha,\omega}$ is a probability measure, we may write
$|\psi_{\alpha}(\omega)-\psi_{\alpha}(\omega^{\prime})|$ $\leq\int_{X}L\,d_{\Omega}(\omega,\omega^{\prime})\,d\nu_{\alpha,\omega}+L\left|\int_{X}\frac{\Phi(\omega^{\prime},\cdot)}{L}\,d\nu_{\alpha,\omega}-\int_{X}\frac{\Phi(\omega^{\prime},\cdot)}{L}\,d\nu_{\alpha,\omega^{\prime}}\right|$
$=Ld_{\Omega}(\omega,\omega^{\prime})+L\Big\langle\nu_{\alpha,\omega}-\nu_{\alpha,\omega^{\prime}},\frac{\Phi(\omega^{\prime},\cdot)}{L}\Big\rangle$
$\leq Ld_{\Omega}(\omega,\omega^{\prime})+L\|\nu_{\alpha,\omega}-\nu_{\alpha,\omega^{\prime}}\|_{k^{*}}.$
Since the first part of (A5) implies that $\bm{\nu}_{\alpha}$ is a continuous map taking values in $C^{k}(X)^{*}$, it follows that $\psi_{\alpha}$ is continuous.
Now we prove that $\psi_{\alpha}$ converges to $\psi_{\alpha_{0}}$ in the supremum norm. Given any $\omega\in\Omega$,
$|\psi_{\alpha}(\omega)-\psi_{\alpha_{0}}(\omega)|$ $=\left|\int_{X}\Phi(\omega,\cdot)\,d\nu_{\alpha,\omega}-\int_{X}\Phi(\omega,\cdot)\,d\nu_{\alpha_{0},\omega}\right|$
$=L\Big|\Big\langle\nu_{\alpha,\omega}-\nu_{\alpha_{0},\omega},\frac{\Phi(\omega,\cdot)}{L}\Big\rangle\Big|$
$\leq L\|\nu_{\alpha,\omega}-\nu_{\alpha_{0},\omega}\|_{k^{*}}.$
Taking supremum in $\omega\in\Omega$ on both sides and using the second part of assumption (A5), the second conclusion follows. ∎
This lemma will be used to complete the proof of Theorem A. By Assumption (A4), there exists $\dot{\rho}_{\alpha_{0}}\in L^{1}(m)$} such that
(3.14) $\lim_{\alpha\to\alpha_{0}}\int_{\Omega}\Big|\frac{\rho_{\alpha}-\rho_{\alpha_{0}}}{\alpha-\alpha_{0}}-\dot{\rho}_{\alpha_{0}}\Big|\,dm=0.$
Define the signed measure $\dot{\eta}_{\alpha_{0}}=\dot{\rho}_{\alpha_{0}}\,m$. We complete the proof of Theorem D by showing that
(3.15) $\frac{d}{d\alpha}\Big|_{\alpha_{0}}\int_{\Omega\times X}\Phi\,d\mu_{\alpha}\;=\;{\int_{\Omega}\langle\dot{\nu}_{\alpha_{0},\omega},\Phi(\omega,\cdot)\rangle\,d\eta_{{\alpha_{0}}}(\omega)}\;+\;{\int_{\Omega}\psi_{\alpha_{0}}\,d\dot{\eta}_{\alpha_{0}}}.$
We may write
$\int_{\Omega\times X}\Phi\,d\mu_{\alpha}-\int_{\Omega\times X}\Phi\,d\mu_{\alpha_{0}}=\int_{\Omega}\psi_{\alpha}\,d\eta_{{\alpha}}-\int_{\Omega}\psi_{\alpha_{0}}\,d\eta_{\alpha_{0}}=A_{\alpha}+B_{\alpha},$
with
$A_{\alpha}=\int_{\Omega}\big(\psi_{\alpha}-\psi_{\alpha_{0}}\big)\,d\eta_{{\alpha_{0}}}\quad\text{and}\quad B_{\alpha}=\int_{\Omega}\psi_{\alpha}\,d\big(\eta_{{\alpha}}-\eta_{\alpha_{0}}\big).$
We first show that
(3.16) $\lim_{\alpha\to\alpha_{0}}\frac{A_{\alpha}}{\alpha-\alpha_{0}}={\int_{\Omega}\langle\dot{\nu}_{\alpha_{0},\omega},\Phi(\omega,\cdot)\rangle\,d\eta_{{\alpha_{0}}}(\omega)}.$
We have
$\frac{A_{\alpha}}{{\alpha-\alpha_{0}}}=\int_{\Omega}\frac{\psi_{\alpha}(\omega)-\psi_{\alpha_{0}}(\omega)}{{\alpha-\alpha_{0}}}\,d\eta_{{\alpha_{0}}}(\omega)=\int_{\Omega}\int_{X}\Phi(\omega,\cdot)\,d\left(\frac{\nu_{\alpha,\omega}-\nu_{\alpha_{0},\omega}}{{\alpha-\alpha_{0}}}\right)d\eta_{{\alpha_{0}}}(\omega).$
Letting $C=\|\Phi\|_{C^{k}}$, we may write for all $\omega\in\Omega$,
$\left|\frac{A_{\alpha}}{{\alpha-\alpha_{0}}}-\int_{\Omega}\langle\dot{\nu}_{\alpha_{0},\omega},\Phi(\omega,\cdot)\rangle d\eta_{{\alpha_{0}}}(\omega)\right|$ $=C\left|\int_{\Omega}\Big\langle\frac{\nu_{\alpha,\omega}-\nu_{\alpha_{0},\omega}}{{\alpha-\alpha_{0}}}-\dot{\nu}_{\alpha_{0},\omega}\,,\frac{\Phi(\omega,\cdot)}{C}\Big\rangle d\eta_{{\alpha_{0}}}(\omega)\right|$
$\leq C\int_{\Omega}\Big\|\frac{\nu_{\alpha,\omega}-\nu_{\alpha_{0},\omega}}{{\alpha-\alpha_{0}}}-\dot{\nu}_{\alpha_{0},\omega}\Big\|_{k^{*}}\,d\eta_{{\alpha_{0}}}(\omega)$
$\leq C\Big\|\frac{\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}}{{\alpha-\alpha_{0}}}-\bm{\dot{\nu}}_{\alpha_{0}}\Big\|_{\mathbf{L}_{k}}.$
Since we assume that the map $\alpha\mapsto\bm{\nu}_{\alpha}$ is differentiable in $\mathbf{L}_{k}^{0}$ at $\alpha_{0}$, the latter expression converges to zero as $\alpha\to\alpha_{0}$, and hence (3.16) holds. It remains to show that
(3.17) $\lim_{\alpha\to\alpha_{0}}\frac{B_{\alpha}}{\alpha-\alpha_{0}}=\int_{\Omega}\psi_{\alpha_{0}}\,d\dot{\eta}_{\alpha_{0}}.$
We have
$\frac{B_{\alpha}}{{\alpha-\alpha_{0}}}=\frac{1}{{\alpha-\alpha_{0}}}\int_{\Omega}\psi_{\alpha}\,d\big(\eta_{{\alpha}}-\eta_{\alpha_{0}}\big)=\frac{1}{{\alpha-\alpha_{0}}}\int_{\Omega}\psi_{\alpha}\big(\rho_{{\alpha}}-\rho_{\alpha_{0}}\big)dm,$
and so
$\frac{B_{\alpha}}{{\alpha-\alpha_{0}}}-{\int_{\Omega}\psi_{\alpha_{0}}\,d\dot{\eta}_{\alpha_{0}}}$ $=\int_{\Omega}\psi_{\alpha}\frac{\rho_{{\alpha}}-\rho_{\alpha_{0}}}{\alpha-\alpha_{0}}dm-{\int_{\Omega}\psi_{\alpha_{0}}\dot{\rho}_{\alpha_{0}}}dm$
$=\int_{\Omega}\psi_{\alpha}\frac{\rho_{{\alpha}}-\rho_{\alpha_{0}}}{\alpha-\alpha_{0}}dm-{\int_{\Omega}\psi_{\alpha}\dot{\rho}_{\alpha_{0}}}dm+{\int_{\Omega}\psi_{\alpha}\dot{\rho}_{\alpha_{0}}}dm-{\int_{\Omega}\psi_{\alpha_{0}}\dot{\rho}_{\alpha_{0}}}dm$
$=\int_{\Omega}\psi_{\alpha}\left(\frac{\rho_{{\alpha}}-\rho_{\alpha_{0}}}{\alpha-\alpha_{0}}-\dot{\rho}_{\alpha_{0}}\right)dm+\int_{\Omega}\left(\psi_{\alpha}-\psi_{\alpha_{0}}\right)d\dot{\eta}_{\alpha_{0}}$
$\leq\|\psi_{\alpha}\|_{C^{0}}\int_{\Omega}\Big|\frac{\rho_{\alpha}-\rho_{\alpha_{0}}}{\alpha-\alpha_{0}}-\dot{\rho}_{\alpha_{0}}\Big|dm+\|\psi_{\alpha}-\psi_{\alpha_{0}}\|_{C^{0}}|\dot{\eta}_{\alpha_{0}}|(\Omega),$
where $|\dot{\eta}_{\alpha_{0}}|$ denotes the variation of the signed measure $\dot{\eta}_{\alpha_{0}}$. Using (3.14) and Lemma 3.1, we obtain (3.17).
4. Base maps with full branches
In this section we prove Theorem E and Corollary F. As observed above, to prove Theorem E it suffices to establish (A3), and to deduce Corollary F it suffices to establish (A5), both for the case $k=3$. These conditions are verified in the next four subsections. More precisely:
(A3)(A3)(a) $\longrightarrow$ Lemma 4.3;
(A3)(A3)(b) $\longrightarrow$ Corollary 4.5;
(A3)(A3)(c) $\longrightarrow$ Corollary 4.11;
(A3)(A3)(d) $\longrightarrow$ Corollary 4.19;
(A5) $\longrightarrow$ Corollary 4.6 and Corollary 4.12.
4.1. Contraction
Here we establish the contraction property (A3)(A3)(a) for the operator $\mathcal{K}_{\alpha}$ acting on the space $\mathbf{E}_{3}^{0}$. Incidentally, this is the only place where using the space $\mathbf{E}_{3}^{0}$ yields a better conclusion than using $\mathbf{E}_{3}$. We begin with a standard lemma, whose second part will only be used in the next subsection.
Lemma 4.1.
Assume that $h:X\to X$ is a $C^{k}$ function, for some $k\geq 1$, and define $L_{j}=\|D^{j}h\|_{C^{0}}$ for $j=1,\dots,k$. Let
$\mathcal{C}_{k}(h)=\max_{1\leq s\leq k}\left\{\sum_{m=s}^{k}{B}_{m,s}(L_{1},\dots,L_{m-s+1})\right\}$
where ${B}_{m,s}$ are the partial Bell polynomials. Then
  1. (1)
    for any $\nu\in C^{k}_{0}(X)^{*}$, we have $\|h_{*}\nu\|_{k^{*}}\leq\mathcal{C}_{k}(h)\|\nu\|_{k^{*}}$;
  2. (2)
    for any $\nu\in C^{k}(X)^{*}$, we have $\|h_{*}\nu\|_{k^{*}}\leq\max\{1,\operatorname{diam}(X)+\mathcal{C}_{k}(h)\}\|\nu\|_{k^{*}}$.
Proof.
Recall that for $\nu\in C^{k}_{0}(X)^{*}$, the dual norm satisfies
$\|\nu\|_{k^{*}}=\sup_{|\varphi|_{C^{k}}\leq 1}\langle\nu,\varphi\rangle.$
By the definition of the push-forward, $\langle h_{*}\nu,\varphi\rangle=\langle\nu,\varphi\circ h\rangle$. To prove the result, it is sufficient to show that the composition $\Phi=\varphi\circ h$ satisfies the seminorm bound $|\Phi|_{C^{k}}\leq\mathcal{C}_{k}(h)|\varphi|_{C^{k}}$}. The $m$-th derivative of the composition $\Phi$ is given by Faà di Bruno’s formula:
$D^{m}\Phi(x)=\sum_{s=1}^{m}(D^{s}\varphi\circ h)(x)\cdot{B}_{m,s}\big(Dh(x),D^{2}h(x),\dots,D^{m-s+1}h(x)\big),$
where ${B}_{m,s}$ are the partial Bell polynomials. These polynomials are sums of terms of the form
$\frac{m!}{j_{1}!j_{2}!\cdots j_{m-s+1}!}\prod_{r=1}^{m-s+1}\left(\frac{D^{r}h}{r!}\right)^{j_{r}}$
subject to $\sum_{r=1}^{m-s+1}j_{r}=s$ and $\sum_{r=1}^{m-s+1}rj_{r}=m$. Taking the supremum norm of $D^{m}\Phi$, we obtain the bound
(4.1) $\|D^{m}\Phi\|_{C^{0}}\leq\sum_{s=1}^{m}\|D^{s}\varphi\|_{C^{0}}\cdot{B}_{m,s}(L_{1},L_{2},\dots,L_{m-s+1}).$
Summing (4.1) over $m=1,\dots,k$ to reconstruct the $C^{k}$ seminorm, we find
$|\Phi|_{C^{k}}=\sum_{m=1}^{k}\|D^{m}\Phi\|_{C^{0}}$ $\leq\sum_{m=1}^{k}\sum_{s=1}^{m}\|D^{s}\varphi\|_{C^{0}}\cdot{B}_{m,s}(L_{1},\dots,L_{m-s+1})$
$=\sum_{s=1}^{k}\|D^{s}\varphi\|_{C^{0}}\left(\sum_{m=s}^{k}{B}_{m,s}(L_{1},\dots,L_{m-s+1})\right).$
Therefore, taking
$\mathcal{C}_{k}(h)=\max_{1\leq s\leq k}\left\{\sum_{m=s}^{k}{B}_{m,s}(L_{1},\dots,L_{m-s+1})\right\},$
we have
$|\Phi|_{C^{k}}\leq\mathcal{C}_{k}(h)\sum_{s=1}^{k}\|D^{s}\varphi\|_{C^{0}}=\mathcal{C}_{k}(h)|\varphi|_{C^{k}}.$
It follows that
$\|h_{*}\nu\|_{k^{*}}=\sup_{|\varphi|_{C^{k}}\leq 1}\langle\nu,\varphi\circ h\rangle\leq\sup_{|\varphi|_{C^{k}}\leq 1}\|\nu\|_{k^{*}}|\varphi\circ h|_{C^{k}}\leq\mathcal{C}_{k}(h)\|\nu\|_{k^{*}},$
thus concluding the proof of the first item.
For the second item, we need the full norm $\|\Phi\|_{C^{k}}=|\Phi(x_{0})|+|\Phi|_{C^{k}}$. By the mean value theorem
$|\varphi(h(x_{0}))|\leq|\varphi(x_{0})|+|h(x_{0})-x_{0}|\cdot\|D\varphi\|_{C^{0}}\leq|\varphi(x_{0})|+\text{diam}(X)|\varphi|_{C^{k}}.$
Combining this with the seminorm bound, we get
$\|\Phi\|_{C^{k}}$ $\leq|\varphi(x_{0})|+\text{diam}(X)|\varphi|_{C^{k}}+\mathcal{C}_{k}(h)|\varphi|_{C^{k}}$
$=|\varphi(x_{0})|+(\text{diam}(X)+\mathcal{C}_{k}(h))|\varphi|_{C^{k}}$
$\leq\max\{1,\text{diam}(X)+\mathcal{C}_{k}(h)\}(|\varphi(x_{0})|+|\varphi|_{C^{k}}).$
Consequently, for any $\nu\in C^{k}(X)^{*}$:
$\|h_{*}\nu\|_{k^{*}}=\sup_{\|\varphi\|_{C^{k}}\leq 1}\langle\nu,\varphi\circ h\rangle\leq\max\{1,\text{diam}(X)+\mathcal{C}_{k}(h)\}\|\nu\|_{k^{*}},$
which concludes the proof. ∎
Remark 4.2.
The constant $\mathcal{C}_{k}(h)$ can be explicitly computed for low orders by evaluating the partial Bell polynomials and summing the coefficients corresponding to each derivative order $\|D^{s}\varphi\|_{C^{0}}$. Note that the arguments of ${B}_{m,s}$ are $(L_{1},\dots,L_{m-s+1})$.
  • Case $k=1$: Only $m=1,s=1$ exists and ${B}_{1,1}(L_{1})=L_{1}$.
    Hence, $\mathcal{C}_{1}(h)=L_{1}$.
  • Case $k=2$: We collect coefficients for $s=1$ and $s=2$:
    • For $s=1$: ${B}_{1,1}(L_{1})+{B}_{2,1}(L_{1},L_{2})=L_{1}+L_{2}$.
    • For $s=2$: ${B}_{2,2}(L_{1})=L_{1}^{2}$.
    Hence, $\mathcal{C}_{2}(h)=\max\{L_{1}+L_{2},\,L_{1}^{2}\}$.
  • Case $k=3$: We evaluate the sums for $s=1,2,3$:
    • For $s=1$: ${B}_{1,1}(L_{1})+{B}_{2,1}(L_{1},L_{2})+{B}_{3,1}(L_{1},L_{2},L_{3})=L_{1}+L_{2}+L_{3}$.
    • For $s=2$: ${B}_{2,2}(L_{1})+{B}_{3,2}(L_{1},L_{2})=L_{1}^{2}+3L_{1}L_{2}$.
    • For $s=3$: ${B}_{3,3}(L_{1})=L_{1}^{3}$.
    Hence, $\mathcal{C}_{3}(h)=\max\{L_{1}+L_{2}+L_{3},\,L_{1}^{2}+3L_{1}L_{2},\,L_{1}^{3}\}$.
The value of $\mathcal{C}_{3}(h)$ motivates our choice of the constant $\lambda$ in (B3)(B3)(b). Indeed, since we assume in particular that $L_{1}+L_{2}+L_{3}<1$, it follows that $L_{1}<1$, and therefore the third term in the expression for $\mathcal{C}_{3}(h)$ can be omitted.
For the next result, $0<\lambda<1$ is the constant specified in (B3)(B3)(b).
Lemma 4.3.
For every $\bm{\nu}\in\mathbf{E}_{3}^{0}$ we have
$\|\mathcal{K}_{\alpha}\bm{\nu}\|_{3}\leq\lambda\|\bm{\nu}\|_{3}.$
Proof.
Let $\bm{\nu}\in\mathbf{E}_{3}^{0}$. By definition, for every $\omega\in\Omega$, the functional $\nu_{\omega}$ belongs to $C^{3}_{0}(X)^{*}$, which implies that $\langle\nu_{\omega},1\rangle=0$. The norm on the Banach space $\mathbf{E}_{3}^{0}$ is given by
$\|\bm{\nu}\|_{3}=\sup_{\omega\in\Omega}\|\nu_{\omega}\|_{3^{*}}.$
For a fixed $\omega\in\Omega$, the transfer operator $\mathcal{K}_{\alpha}$ acts on the section $\bm{\nu}$ as
$(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}=\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}.$
Applying the dual norm $\|\cdot\|_{3^{*}}$ to both sides and using the triangle inequality, we obtain:
(4.2) $\|(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}\|_{3^{*}}\leq\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}\|_{3^{*}}.$
For each $i\in\mathbb{I}$, let $h_{i,\omega}=g_{\alpha,\theta_{\alpha,i}(\omega)}$. Since $\nu_{\theta_{\alpha,i}(\omega)}\in C^{3}_{0}(X)^{*}$, we apply the push-forward estimate in Lemma 4.1 for $k=3$. This yields
$\|(h_{i,\omega})_{*}\nu_{\theta_{\alpha,i}(\omega)}\|_{3^{*}}\leq\mathcal{C}_{3}(h_{i,\omega})\|\nu_{\theta_{\alpha,i}(\omega)}\|_{3^{*}},$
where the constant $\mathcal{C}_{3}(h_{i,\omega})$ is determined by the derivative norms $L_{j}=\|D^{j}g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{0}}$ as
$\mathcal{C}_{3}(h_{i,\omega})=\max\{L_{1}+L_{2}+L_{3},\,L_{1}^{2}+3L_{1}L_{2},\,L_{1}^{3}\};$
recall Remark 4.2. By in (B3)(B3)(b), $\lambda$ is a uniform upper bound for $\mathcal{C}_{3}(h_{i,\omega})$ over all $i\in\mathbb{I}$ and $\omega\in\Omega$. Specifically, we assume $\sup_{\alpha,i,\omega}\mathcal{C}_{3}(g_{\alpha,\theta_{\alpha,i}(\omega)})\leq\lambda<1$. Substituting this bound into (4.2), we get
$\|(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}\|_{3^{*}}$ $\leq\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\lambda\|\nu_{\theta_{\alpha,i}(\omega)}\|_{3^{*}}$
$\leq\lambda\left(\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\right)\sup_{\omega^{\prime}\in\Omega}\|\nu_{\omega^{\prime}}\|_{3^{*}}$
$=\lambda\|\bm{\nu}\|_{3},$
where we used the normalization condition $\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)=1$ for all $\omega\in\Omega$. Taking the supremum over $\omega\in\Omega$ on the left-hand side, we conclude the proof. ∎
4.2. Invariance
With the application to the intermittent solenoid in Subsection 1.4.2 in mind, we emphasise that the arguments used to establish (A3)(A3)(b) and (A5) rely solely on assumptions (B1)(B2), together with the following weakened form of assumption (B4).
  1. (B4’)
    For each $i\in\mathbb{I}$, the map $\alpha\mapsto p_{\alpha,i}\in C^{0}(\Omega,\mathbb{R})$ is continuous, and there exists $C^{\prime}_{p}>0$ such that
    $\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\|p_{\alpha,i}\|_{C^{0}}\leq C^{\prime}_{p}.$
Define for each section $\bm{\nu}$ and $\omega\in\Omega$
(4.3) $\mathcal{(}\mathcal{K}_{\alpha,i}\bm{\nu})_{\omega}=p_{\alpha,i}(\omega)(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)},$
so that by (1.12) we have
(4.4) $\mathcal{K}_{\alpha}=\sum_{i\in\mathbb{I}}\mathcal{K}_{\alpha,i}.$
Lemma 4.4.
Assume that  (B1)(B2) and (B4’) hold. Then, the operator $\mathcal{K}_{\alpha}$ maps $\mathbf{E}_{\mathrm{L}}$ into itself.
Proof.
Considering the sum in (4.4), we first show that each $\mathcal{K}_{\alpha,i}$ maps $\mathbf{E}_{\mathrm{L}}$ into itself. Take an arbitrary $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}$. We first show that $(\mathcal{K}_{\alpha,i}\bm{\nu})_{\omega}\in\mathrm{Lip}(X)^{*}$, for all $\omega\in\Omega$. By assumption (B3), each $g_{\alpha,\theta_{\alpha,i}(\omega)}$ is a $C^{3}$ map of $X$, so in particular it is a Lipschitz map. By definition, we have for each $\phi\in\operatorname{Lip}(X)$
$\bigl\langle(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)},\phi\bigr\rangle=\langle\nu_{\theta_{\alpha,i}(\omega)},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle,$
which clearly shows that $(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}$ is an element of $\mathrm{Lip}(X)^{*}$. Since multiplication by the scalar $p_{\alpha,i}(\omega)$ does not affect this fact, we have $(\mathcal{K}_{\alpha,i}\bm{\nu})_{\omega}\in\mathrm{Lip}(X)^{*}$ for all $\omega\in\Omega$.
Now we show the continuity of the map $\omega\mapsto(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}$ in $\mathrm{Lip}(X)^{*}$. Given $\omega,\omega^{\prime}\in\Omega$, we have
$\|(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}-(\mathcal{K}_{\alpha}\bm{\nu})_{\omega^{\prime}}\|_{\mathrm{Lip}^{*}}\leq\sum_{i\in\mathbb{I}}\bigl\|p_{\alpha,i}(\omega)\,(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}-p_{\alpha,i}(\omega^{\prime})\,(g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})})_{*}\nu_{\theta_{\alpha,i}(\omega^{\prime})}\bigr\|_{\mathrm{Lip}^{*}}.$
We can split each term as
$\bigl\|p_{\alpha,i}(\omega)(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}-p_{\alpha,i}(\omega^{\prime})(g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})})_{*}\nu_{\theta_{\alpha,i}(\omega^{\prime})}\bigr\|_{\mathrm{Lip}^{*}}$
$\quad\leq|p_{\alpha,i}(\omega)-p_{\alpha,i}(\omega^{\prime})|\cdot\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}\|_{\mathrm{Lip}^{*}}$
$\quad\quad+|p_{\alpha,i}(\omega^{\prime})|\cdot\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}-(g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})})_{*}\nu_{\theta_{\alpha,i}(\omega^{\prime})}\|_{\mathrm{Lip}^{*}}.$
For the first term, we have by (B3)(B3)(b) a uniform constant $L>0$ such that for any $\phi\in\operatorname{Lip}(X)$ with $\|\phi\|_{\mathrm{L}}\leq 1$ we have
$\|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{\mathrm{L}}$ $=|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}(x_{0})|+\mathrm{Lip}(\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)})$
$\leq 1+\sup_{x\in X}\|Dg_{\alpha,\theta_{\alpha,i}(\omega)}(x)\|$
(4.5) $\leq 1+L.$
Hence
$\bigl\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}\bigr\|_{\mathrm{Lip}^{*}}$ $=\sup_{\|\phi\|_{\mathrm{L}}\leq 1}\bigl|\langle\nu_{\theta_{\alpha,i}(\omega)},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle\bigr|$
$\leq\sup_{\|\phi\|_{\mathrm{L}}\leq 1}\|\nu_{\theta_{\alpha,i}(\omega)}\|_{\mathrm{Lip}^{*}}\|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{\mathrm{L}}$
(4.6) $\leq\|\bm{\nu}\|_{\mathrm{L}}(1+L).$
Hence
$|p_{\alpha,i}(\omega)-p_{\alpha,i}(\omega^{\prime})|\cdot\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}\|_{\mathrm{Lip}^{*}}\leq L\|\bm{\nu}\|_{\mathrm{L}}|p_{\alpha,i}(\omega)-p_{\alpha,i}(\omega^{\prime})|.$
On the other hand, using (4.2)
$\bigl\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}$ $\nu_{\theta_{\alpha,i}(\omega)}-(g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})})_{*}\nu_{\theta_{\alpha,i}(\omega^{\prime})}\bigr\|_{\mathrm{Lip}^{*}}$
$\quad=\sup_{\|\phi\|_{\mathrm{L}}\leq 1}\Bigl|\langle\nu_{\theta_{\alpha,i}(\omega)},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle-\langle\nu_{\theta_{\alpha,i}(\omega^{\prime})},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})}\rangle\Bigr|$
$\quad\leq\sup_{\|\phi\|_{\mathrm{L}}\leq 1}\Bigl|\langle\nu_{\theta_{\alpha,i}(\omega)}-\nu_{\theta_{\alpha,i}(\omega^{\prime})},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle\Bigr|$
$\quad\quad\quad\quad+\sup_{\|\phi\|_{\mathrm{L}}\leq 1}\Bigl|\langle\nu_{\theta_{\alpha,i}(\omega^{\prime})},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}-\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})}\rangle\Bigr|$
$\quad\leq\|\nu_{\theta_{\alpha,i}(\omega)}-\nu_{\theta_{\alpha,i}(\omega^{\prime})}\|_{\mathrm{Lip}^{*}}\,\sup_{\|\phi\|_{\mathrm{L}}\leq 1}\|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{\mathrm{L}}$
$\quad\quad\quad\quad+\|\nu_{\theta_{\alpha,i}(\omega^{\prime})}\|_{\mathrm{Lip}^{*}}\,{\sup_{\|\phi\|_{\mathrm{L}}\leq 1}\|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}-\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})}\|_{C^{0}}}$
$\quad\leq(1+L)\|\nu_{\theta_{\alpha,i}(\omega)}-\nu_{\theta_{\alpha,i}(\omega^{\prime})}\|_{\mathrm{Lip}^{*}}+\|\bm{\nu}\|_{\mathrm{L}}\,\|g_{\alpha,\theta_{\alpha,i}(\omega)}-g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})}\|_{C^{0}}.$
Therefore
$\|(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}-(\mathcal{K}_{\alpha}\bm{\nu})_{\omega^{\prime}}\|_{\mathrm{Lip}^{*}}\leq$ $(1+L)\|\bm{\nu}\|_{\mathrm{L}}|p_{\alpha,i}(\omega)-p_{\alpha,i}(\omega^{\prime})|$
$+(1+L)\|\nu_{\theta_{\alpha,i}(\omega)}-\nu_{\theta_{\alpha,i}(\omega^{\prime})}\|_{\mathrm{Lip}^{*}}$
$+\|\bm{\nu}\|_{\mathrm{L}}\,\|g_{\alpha,\theta_{\alpha,i}(\omega)}-g_{\alpha,\theta_{\alpha,i}(\omega^{\prime})}\|_{C^{0}}.$
Using assumptions (B3)(B2), and (B4’), together with the uniform continuity of $\bm{\nu}$, we deduce that the three differences above can be made arbitrarily small whenever $\omega$ and $\omega^{\prime}$ are sufficiently close. Hence, the map $\omega\mapsto(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}$ is continuous.
Finally, using assumption (B4’) together with the uniform bound (4.6), we deduce that the series (4.4) defining $\mathcal{K}_{\alpha}\bm{\nu}$ converges uniformly in $\mathbf{E}_{\mathrm{L}}$. By standard results on series in Banach spaces, it then follows that $\mathcal{K}_{\alpha}\bm{\nu}\in\mathbf{E}_{\mathrm{L}}$. ∎
In the next result we establish (A3)(A3)(b).
Corollary 4.5.
Assume that  (B1)(B2) and (B4’) hold. Then, $\mathcal{K}_{\alpha}$ maps $\mathbf{E}_{\mathrm{L}}^{0}$ into itself.
Proof.
This is a straightforward consequence of Lemma 4.4, together with the observation that push-forwards preserve functionals in $\mathrm{Lip}_{0}(X)^{*}$, that is, those vanishing on constant functions. ∎
In the net result we obtain the first part of (A5).
Corollary 4.6.
Assume that  (B1)(B2) and (B4’) hold. Then, the fixed point of $\mathcal{K}_{\alpha}$ in $\mathbf{P}$ belongs in $\mathbf{P}_{0}$.
Proof.
Since $\mathbf{P}_{0}$ is a closed subset of $\mathbf{P}$, it suffices to show that it is invariant under $\mathcal{K}_{\alpha}$. Let $\bm{\nu}\in\mathbf{P}_{0}$. By definition, $\mathbf{P}_{0}\subset\mathbf{E}_{\mathrm{L}}$, so the continuity of $\mathcal{K}_{\alpha}\bm{\nu}$ follows from Lemma 4.4. It remains to check that each $(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}$ is a probability measure. This follows immediately from the properties of push-forwards and the weights: since $(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}$ preserves total mass, and $p_{\alpha,i}(\omega)\geq 0$ with $\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)=1$, we conclude that $(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}\in\mathcal{M}_{1}(X)$. ∎
4.3. Continuity
In this subsection, we establish the continuity property (A3)(A3)(c) and the second part of (A5), both on the space $\mathbf{E}_{3}^{0}$. These results will follow as consequences of Proposition 4.9 below. We emphasise that, also in this part, our conclusions rely solely on assumptions (B1)(B2), together with the weaker version (B4’) of assumption (B4) introduced in Subsection 4.2.
Lemma 4.7.
Given $\nu\in\mathrm{Lip}(X)^{*}$ and $C^{1}$ maps $h_{0},h_{1}:X\to X$, there exists $\mathcal{C}(h_{1},h_{2})>0$ depending only on the $C^{1}$ norms of $h_{0}$ and $h_{1}$ such that, for any $\phi\in C^{2}(X)$,
$\big|\langle(h_{1})_{*}\nu-(h_{0})_{*}\nu,\phi\rangle\big|\leq\mathcal{C}(h_{1},h_{2})\,\|\nu\|_{\mathrm{Lip}^{*}}\,\|\phi\|_{C^{2}}\,\|h_{1}-h_{0}\|_{C^{1}}.$
Proof.
By definition of the push-forward,
$\langle(h_{1})_{*}\nu-(h_{0})_{*}\nu,\phi\rangle=\langle\nu,\phi\circ h_{1}-\phi\circ h_{0}\rangle.$
Since $X$ is convex, define $h_{t}=h_{0}+t(h_{1}-h_{0})$ for $t\in[0,1]$. Then
$\phi\circ h_{1}-\phi\circ h_{0}=\int_{0}^{1}\frac{d}{dt}\phi\circ h_{t}\,dt=\int_{0}^{1}(\nabla\phi\circ h_{t})\bm{\cdot}(h_{1}-h_{0})\,dt.$
Set
$\psi_{t}=(\nabla\phi\circ h_{t})\bm{\cdot}(h_{1}-h_{0}).$
Since $\phi\in C^{2}(X)$, each $\psi_{t}$ is Lipschitz and
$\mathrm{Lip}(\psi_{t})$ $\leq\|\nabla\phi\circ h_{t}\|_{C^{0}}\mathrm{Lip}(h_{1}-h_{0})+\mathrm{Lip}(\nabla\phi\circ h_{t})\|h_{1}-h_{0}\|_{C^{0}}$
$\leq\|\phi\|_{C^{1}}\|h_{1}-h_{0}\|_{C^{1}}+\|\phi\|_{C^{2}}\|h_{t}\|_{C^{1}}\|h_{1}-h_{0}\|_{C^{0}}$
(4.7) $\leq\|\phi\|_{C^{2}}\|h_{1}-h_{0}\|_{C^{1}}(1+\|h_{t}\|_{C^{1}}).$
This gives in particular that $t\mapsto\psi_{t}\in\mathrm{Lip}(X)$ is Bochner integrable in the Banach space $\operatorname{Lip}(X)$. By the properties of this integral,
$\left|\langle\nu,\int_{0}^{1}\psi_{t}\,dt\rangle\right|=\left|\int_{0}^{1}\langle\nu,\psi_{t}\rangle\,dt\right|\leq\int_{0}^{1}|\langle\nu,\psi_{t}\rangle|\,dt.$
On the other hand, by the definition of the dual norm,
$|\langle\nu,\psi_{t}\rangle|\leq\|\nu\|_{\mathrm{Lip}^{*}}\|\psi_{t}\|_{\operatorname{Lip}}\quad\text{for all }t\in[0,1].$
Hence
$|\langle(h_{1})_{*}\nu-(h_{0})_{*}\nu,\phi\rangle|\leq\int_{0}^{1}|\langle\nu,\psi_{t}\rangle|\,dt\leq\|\nu\|_{\mathrm{Lip}^{*}}\int_{0}^{1}\|\psi_{t}\|_{\operatorname{Lip}}\,dt.$
Finally, using the uniform bound for $\|\psi_{t}\|_{\operatorname{Lip}}$ obtained in (4.7) and integrating over $t$ we obtain the desired conclusion. ∎
Example 4.8.
With this example we show that Lemma 4.7 cannot be established in the space $C^{1}(X)$. Indeed, let $X=[0,1]$ and define for $\varepsilon>0$ the dipoles
$\nu_{\varepsilon}=\frac{1}{\varepsilon}(\delta_{1}-\delta_{1-\varepsilon}),$
which satisfy $\|\nu_{\varepsilon}\|_{\mathrm{Lip}^{*}}=1$. Consider the family of contractions $h_{t}(z)=(1-t)z$ for $t\in(0,1)$. Note that $\|h_{t}-h_{s}\|_{C^{1}}=|t-s|$. For any $\phi\in C^{1}(X)$, we have
$\langle(h_{t})_{*}\nu_{\varepsilon}-(h_{s})_{*}\nu_{\varepsilon},\phi\rangle=\frac{\phi(1-t)-\phi((1-t)(1-\varepsilon))}{\varepsilon}-\frac{\phi(1-s)-\phi((1-s)(1-\varepsilon))}{\varepsilon}.$
As $\varepsilon\to 0$, this converges to
$\Delta(t,s):=(1-t)\phi^{\prime}(1-t)-(1-s)\phi^{\prime}(1-s).$
Now, choose $\phi\in C^{1}(X)$ such that $\phi^{\prime}(y)=|1-y|^{\beta}$, with $\beta\in(0,1)$. Then
$\Delta(t,s)=(1-t)t^{\beta}-(1-s)s^{\beta}.$
Evaluating the ratio for $s=2t$ with small $t$, we obtain
$\frac{|\Delta(t,2t)|}{\|h_{t}-h_{2t}\|_{C^{0}}}=\frac{|(1-t)t^{\beta}-(1-2t)(2t)^{\beta}|}{t}=t^{\beta-1}|(1-t)-2^{\beta}(1-2t)|.$
As $t\to 0$, the term in the absolute value converges to $|1-2^{\beta}|\neq 0$. Thus, the ratio behaves as $ct^{\beta-1}$, which diverges because $\beta<1$. This confirms that $C^{1}$ regularity is insufficient to provide a uniform Lipschitz constant for the push-forward operator.
The next result is formulated for elements in the space $\mathbf{E}_{\mathrm{L}}$, which encompasses both spaces $\mathbf{E}_{\mathrm{L}}^{0}$ and $\mathbf{P}_{0}$ on which we have to verify properties (A3)(A3)(c) and (A5), respectively.
Proposition 4.9.
Assume that (B1)(B2) and (B4’) hold and let $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}$. Then,
  1. (1)
    the map $\alpha\mapsto\mathcal{K}_{\alpha,i}\bm{\nu}$ is continuous in $\mathbf{E}_{2}$, for each $i\in\mathbb{I}$;
  2. (2)
    the series $\sum_{i\in\mathbb{I}}{\mathcal{K}}_{\alpha,i}\bm{\nu}$ converges in $\mathbf{E}_{2}$ uniformly in $\alpha$.
Proof.
We first show that, for fixed $i\in\mathbb{I}$ and $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}$, the map $\alpha\mapsto\mathcal{K}_{\alpha,i}\bm{\nu}$ is continuous in $\mathbf{E}_{2}$. The action of the operator $\mathcal{K}_{\alpha,i}$ on the section $\bm{\nu}$ is given pointwise by
$(\mathcal{K}_{\alpha,i}\bm{\nu})]_{\omega}=p_{\alpha,i}(\omega)(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}.$
Hence,
$\|[(\mathcal{K}_{\alpha,i}\bm{\nu})]_{\omega}-[(\mathcal{K}_{\alpha_{0},i}\bm{\nu})]_{\omega}\|_{2^{*}}$ $=\|p_{\alpha,i}(\omega)(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}-p_{\alpha_{0},i}(\omega)(g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)})_{*}\nu_{\theta_{\alpha_{0},i}(\omega)}\|_{2^{*}}$
(4.8) $\leq A(\omega)+B(\omega)+C(\omega),$
with
$A(\omega)$ $=|p_{\alpha,i}(\omega)-p_{\alpha_{0},i}(\omega)|\cdot\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}\|_{2^{*}}$
$B(\omega)$ $=|p_{\alpha_{0},i}(\omega)|\cdot\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}-(g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}\|_{2^{*}}$
$C(\omega)$ $=|p_{\alpha_{0},i}(\omega)|\cdot\|(g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)})_{*}(\nu_{\theta_{\alpha,i}(\omega)}-\nu_{\theta_{\alpha_{0},i}(\omega)})\|_{2^{*}}.$
We estimate these terms separately. Using the second item of Lemma 4.1 and Remark 4.2 and since $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}\subset\mathbf{E}_{2}$, we have a constant $L>0$ depending only on the constants $L_{1},L_{2}$ in (B3)(B3)(b) such that
$\big\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}\big\|_{2^{*}}\leq L\,\|\nu_{\theta_{\alpha,i}(\omega)}\|_{2^{*}}\leq L\,\,\|\bm{\nu}\|_{2}.$
Thus,
(4.9) $A(\omega)\leq L\,\|\bm{\nu}\|_{2}\cdot\|p_{\alpha,i}-p_{\alpha_{0},i}\|_{C^{0}}.$
For the second term, we apply Lemma 4.7 to the maps $h_{1}=g_{\alpha,\theta_{\alpha,i}(\omega)}$, $h_{2}=g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}$ and $\nu=\nu_{\theta_{\alpha,i}(\omega)}\in\operatorname{Lip}(X)^{*}$, which together with (B3) yields a uniform constant $C>0$ such that for any $\phi\in C^{2}(X)$
$\big|\langle(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}-(g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)},\phi\rangle\big|\leq C\|\nu_{\theta_{\alpha,i}(\omega)}\|_{\operatorname{Lip}^{*}}\,\|\phi\|_{C^{2}}\,\|g_{\alpha,\theta_{\alpha,i}(\omega)}-g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}\|_{C^{0}}.$
Taking the supremum over $\|\phi\|_{C^{2}}\leq 1$ and since $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}$, we get
$\|(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}-(g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)})_{*}\nu_{\theta_{\alpha,i}(\omega)}\|_{2^{*}}\leq\|\bm{\nu}\|_{\mathrm{L}}\|g_{\alpha,\theta_{\alpha,i}(\omega)}-g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}\|_{C^{0}}.$
Hence,
(4.10) $B(\omega)\leq\|p_{\alpha_{0},i}\|_{C^{0}}\|\bm{\nu}\|_{\mathrm{L}}\|g_{\alpha,\theta_{\alpha,i}(\omega)}-g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}\|_{C^{0}}.$
Using again the second part of Lemma 4.1 and Remark 4.2 together with (B3) for the third term, we obtain
(4.11) $C(\omega)\leq L\|p_{\alpha_{0},i}\|_{C^{0}}\|\nu_{\theta_{\alpha,i}(\omega)}-\nu_{\theta_{\alpha_{0},i}(\omega)}\|_{2^{*}}.$
Putting (4.8) and (4.9)–(4.11) together, we get
$\|[(\mathcal{K}_{\alpha,i}\bm{\nu})]_{\omega}-[(\mathcal{K}_{\alpha_{0},i}\bm{\nu})]_{\omega}\|_{2^{*}}$ $\leq L\|\bm{\nu}\|_{3}\cdot\|p_{\alpha,i}-p_{\alpha_{0},i}\|_{C^{0}}$
$\quad+C\|p_{\alpha_{0},i}\|_{C^{0}}\|\bm{\nu}\|_{2}\|g_{\alpha,\theta_{\alpha,i}(\omega)}-g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}\|_{C^{1}}$
(4.12) $\quad+L\|p_{\alpha_{0},i}\|_{C^{0}}\|\nu_{\theta_{\alpha,i}(\omega)}-\nu_{\theta_{\alpha_{0},i}(\omega)}\|_{2^{*}}.$
Taking the supremum over $\omega\in\Omega$, and using continuity assumptions (B1)(B4) and the uniform continuity of the section $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}\subset\mathbf{E}_{2}$, we conclude the proof of the first part. The second part follows from the estimate obtained in (4.12) and the summability conditions in (B1)(B4). ∎
Remark 4.10.
Even when the weights and the fibre maps do not depend on the base point, the regularity of the space $\mathbf{E}_{2}$ in which Proposition 4.9 is proved cannot be improved. Indeed, $C^{2}$ regularity is essential in the proof of Lemma 4.7, as Example 4.8 illustrates. Moreover, even in the base-independent case, the continuity of the term $B(\omega)$ in the proof of Proposition 4.9 requires comparing push-forwards associated with two nearby but distinct maps, due to the dependence on the parameter.
The next result provides (A3)(A3)(c) in the space $\mathbf{E}_{3}^{0}$.
Corollary 4.11.
Assume that (B1)(B2) and (B4’) hold and let $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}$. If $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}^{0}$, then the map $\alpha\mapsto\mathcal{K}_{\alpha}\bm{\nu}$ is continuous in $\mathbf{E}_{3}^{0}$.
Proof.
First, observe that by basic calculus in Banach spaces, Proposition 4.9 implies that for any $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}^{0}\subset\mathbf{E}_{\mathrm{L}}$, the map $\alpha\mapsto\mathcal{K}_{\alpha}\bm{\nu}$ is continuous as a map into $\mathbf{E}_{2}$. Since we have continuous inclusions $\mathbf{E}_{3}^{0}\subset\mathbf{E}_{3}\subset\mathbf{E}_{2}$, convergence in $\mathbf{E}_{2}$ therefore implies convergence in $\mathbf{E}_{3}^{0}$. ∎
The next result is a consequence of Lemma 4.3 and Proposition 4.9, and provides the second part of (A5).
Corollary 4.12.
Assume that (B1)(B2) and (B4’) hold and let $\bm{\nu}\in\mathbf{E}_{\mathrm{L}}$. If $\bm{\nu}_{\alpha},\bm{\nu}_{\alpha_{0}}$ are respectively the unique fixed points of $\mathcal{K}_{\alpha},\mathcal{K}_{\alpha_{0}}$ in $\mathbf{P}_{0}$, then
$\lim_{\alpha\to\alpha_{0}}\|\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}\|_{3}=0.$
Proof.
Recall that the difference $\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}$ belongs to the space $\mathbf{E}_{\mathrm{L}}^{0}\subset\mathbf{E}_{3}^{0}$. Using the fixed point property,
$\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}=\mathcal{K}_{\alpha}\bm{\nu}_{\alpha}-\mathcal{K}_{\alpha_{0}}\bm{\nu}_{\alpha_{0}}.$
We decompose the right-hand side as
$\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}$ $=(\mathcal{K}_{\alpha}\bm{\nu}_{\alpha}-\mathcal{K}_{\alpha}\bm{\nu}_{\alpha_{0}})+(\mathcal{K}_{\alpha}\bm{\nu}_{\alpha_{0}}-\mathcal{K}_{\alpha_{0}}\bm{\nu}_{\alpha_{0}})$
$=\mathcal{K}_{\alpha}(\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}})+(\mathcal{K}_{\alpha}\bm{\nu}_{\alpha_{0}}-\mathcal{K}_{\alpha_{0}}\bm{\nu}_{\alpha_{0}}).$
Using the contraction given by Lemma 4.3, we get
$\|\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}\|_{3}\leq\lambda\|\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}\|_{3}+\|(\mathcal{K}_{\alpha}-\mathcal{K}_{\alpha_{0}})\bm{\nu}_{\alpha_{0}}\|_{3}.$
Rearranging the terms, we obtain
$(1-\lambda)\|\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}\|_{3}\leq\|(\mathcal{K}_{\alpha}-\mathcal{K}_{\alpha_{0}})\bm{\nu}_{\alpha_{0}}\|_{3}.$
Recalling that $\mathbf{P}_{0}\subset\mathbf{E}_{\mathrm{L}}$, taking $\bm{\nu}=\bm{\nu}_{\alpha_{0}}$ in Corollary 4.11, we obtain
$\lim_{\alpha\to\alpha_{0}}\|\mathcal{K}_{\alpha}\bm{\nu}_{\alpha_{0}}-\mathcal{K}_{\alpha_{0}}\bm{\nu}_{\alpha_{0}}\|_{3}=0.$
Since $1-\lambda>0$, it follows that $\|\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}\|_{3}\to 0$ as $\alpha\to\alpha_{0}$. ∎
4.4. Differentiability
Here we establish the validity of (A3)(A3)(d) for any fixed $\alpha_{0}\in J$, with the argument divided into two parts. We first establish the $C^{1}$ regularity of the invariant section $\bm{\nu}_{\alpha}$, viewed as a map from the base space $\Omega$ into the space $C^{2}(X)^{*}$. We then use this spatial regularity to deduce the parameter differentiability required by (A3)(A3)(d).
Regarding the $C^{1}$ regularity of the invariant section $\omega\mapsto\nu_{\alpha,\omega}$, we remark that one could derive a Lasota–Yorke inequality for the operator $\mathcal{K}_{\alpha}$, taking $\mathbf{E}_{2}^{\prime}$ as the strong space and $\mathbf{E}_{2}$ as the weak one. However, such an estimate would be of no use in the present context for establishing the $C^{1}$ differentiability of the fixed point $\bm{\nu}_{\alpha}$. Indeed, due to the infinite dimensionality of $C^{2}(X)^{*}$, the space $\mathbf{E}_{2}^{\prime}$ is not compactly embedded in $\mathbf{E}_{2}$. Instead, we adopt a graph transform–type approach.
4.4.1. Differentiability of the invariant section
Our objective is to establish that the invariant section $\bm{\nu}_{\alpha}$, defined as the unique fixed point of the sectional transfer operator
$(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}=\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}{\nu}_{\theta_{\alpha,i}(\omega)}$
is of class $C^{1}$ as a mapping from the base space $\Omega$ into $C^{2}(X)^{*}$. By formally differentiating the fixed-point identity $\mathcal{K}_{\alpha}\bm{\nu}_{\alpha}=\bm{\nu}_{\alpha}$ with respect to $\omega$, we observe that the derivative $\bm{\nu}_{\alpha}^{\prime}$ must satisfy a linear functional equation. This equation naturally leads to the definition of a tangent operator, which describes the first-order variation of the invariant section by accounting for the simultaneous infinitesimal changes in the weights, the fiber dynamics, and the base map shifts.
We define the tangent operator $\mathcal{T}:\mathbf{E}_{2}^{0}\to\mathbf{E}_{2}^{0}$ as the affine mapping given for any $\bm{\xi}\in\mathbf{E}_{2}^{0}$ and $\omega\in\Omega$ by
$(\mathcal{T}\bm{\xi})_{\omega}={\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,\theta_{\alpha,i}^{\prime}(\omega)\,(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\xi_{\theta_{\alpha,i}(\omega)}}+{(\mathcal{R}\bm{\nu}_{\alpha})_{\omega}}.$
The term $(\mathcal{R}\bm{\nu}_{\alpha})$ is the image of the invariant section $\bm{\nu}_{\alpha}\in\mathbf{P}$ under the remainder operator $\mathcal{R}$. It will be convenient to have this operator defined more generally for an arbitrary section $\bm{\nu}\in\mathbf{P}$, by prescribing, for each $\omega\in\Omega$ and each test function $\phi\in C^{2}(X)$,
$\langle(\mathcal{R}\bm{\nu})_{\omega},\phi\rangle=\sum_{i\in\mathbb{I}}p^{\prime}_{\alpha,i}(\omega)\,\langle\nu_{\theta_{\alpha,i}(\omega)},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle+\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,\langle\nu_{\theta_{\alpha,i}(\omega)},\partial_{\omega}(\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)})\rangle.$
Note that since $\sum p_{\alpha,i}(\omega)=1$ and $\langle\nu_{\omega},1\rangle=1$ for all $\omega$, it follows that $\mathcal{R}\bm{\nu}\in\mathbf{E}_{2}^{0}$. The linear part of the operator $\mathcal{T}$ describes how derivatives propagate through the base dynamics, whereas the remainder term $\mathcal{R}\bm{\nu}_{\alpha}$ accounts for the variations arising from the $\omega$–dependence of the weights and of the fiber maps. The key observation is that, if the invariant section $\bm{\nu}_{\alpha}$ is differentiable with respect to $\omega$, then its derivative $\bm{\nu}_{\alpha}^{\prime}$ necessarily satisfies the fixed-point identity
$\mathcal{T}\bm{\nu}_{\alpha}^{\prime}=\bm{\nu}_{\alpha}^{\prime}.$
Thus, the existence and uniqueness of a fixed point of $\mathcal{T}$ in $\mathbf{E}_{2}^{0}$ provides the unique candidate for the derivative of the invariant section. The $C^{1}$ regularity of $\bm{\nu}_{\alpha}$ then follows from the uniform convergence of $\partial_{\omega}(\mathcal{K}_{\alpha}^{n}\bm{\nu})$ to this fixed point, starting from any smooth $\bm{\nu}\in\mathbf{P}$, together with the fundamental theorem of calculus in Banach spaces.
Lemma 4.13.
For any $\bm{\nu}\in\mathbf{P}$, we have
$\lim_{n\to\infty}\|\mathcal{R}(\mathcal{K}_{\alpha}^{n}\bm{\nu}-\bm{\nu}_{\alpha})\|_{2}=0.$
Proof.
Fix $\omega\in\Omega$ and $\phi\in C^{2}(X)$ with $\|\phi\|_{C^{2}}\leq 1$. Set $\bm{\nu}^{(n)}=\mathcal{K}_{\alpha}^{n}\bm{\nu}$ and split
$\mathcal{R}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha})=\mathcal{R}_{1}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha})+\mathcal{R}_{2}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha}),$
with
$\langle(\mathcal{R}_{1}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha}))_{\omega},\phi\rangle=\sum_{i\in\mathbb{I}}p_{\alpha,i}^{\prime}(\omega)\langle{\nu}^{(n)}_{\theta_{\alpha,i}(\omega)}-{\nu}_{\alpha,\theta_{\alpha,i}(\omega)},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle$
and
$\langle(\mathcal{R}_{2}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha}))_{\omega},\phi\rangle=\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\langle{\nu}^{(n)}_{\theta_{\alpha,i}(\omega)}-{\nu}_{\alpha,\theta_{\alpha,i}(\omega)},\partial_{\omega}(\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)})\rangle.$
By the chain rule we have $\|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}\leq\|\phi\|_{C^{1}}\|g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}$, and the definition of the dual norm gives
$|\langle(\mathcal{R}_{1}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha}))_{\omega},\phi\rangle|$ $\leq\sum_{i\in\mathbb{I}}|p_{\alpha,i}^{\prime}(\omega)|\,\|{\nu}^{(n)}_{\theta_{\alpha,i}(\omega)}-{\nu}_{\alpha,\theta_{\alpha,i}(\omega)}\|_{1^{*}}\,\|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}$
$\leq\sum_{i\in\mathbb{I}}|p_{\alpha,i}^{\prime}(\omega)|\,\|{\nu}^{(n)}_{\theta_{\alpha,i}(\omega)}-{\nu}_{\alpha,\theta_{\alpha,i}(\omega)}\|_{1^{*}}\,\|\phi\|_{C^{1}}\|g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}.$
Taking supremum over $\phi\in C^{2}(X)$ with $\|\phi\|_{C^{2}}\leq 1$ and noting that $\|\phi\|_{C^{1}}\leq\|\phi\|_{C^{2}}\leq 1$, we get
$\|(\mathcal{R}_{1}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha}))_{\omega}\|_{2^{*}}\leq\sum_{i\in\mathbb{I}}|p_{\alpha,i}^{\prime}(\omega)|\,\|{\nu}^{(n)}_{\theta_{\alpha,i}(\omega)}-{\nu}_{\alpha,\theta_{\alpha,i}(\omega)}\|_{1^{*}}\,\|g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}.$
Similarly, using the chain rule to get the bound $\|\partial_{\omega}(\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)})\|_{C^{1}}\leq\|\phi\|_{C^{1}}\|\partial_{\omega}g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}$, it follows that
$\|(\mathcal{R}_{2}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha}))_{\omega}\|_{2^{*}}\leq\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,\|{\nu}^{(n)}_{\theta_{\alpha,i}(\omega)}-{\nu}_{\alpha,\theta_{\alpha,i}(\omega)}\|_{1^{*}}\,\|\partial_{\omega}g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}.$
Since $\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)=1$, for all $\omega\in\Omega$, using (B3) we deduce
$\sup_{\omega\in\Omega}\sum_{i\in\mathbb{I}}|p_{\alpha,i}^{\prime}(\omega)|\,\|g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{2}}<\infty\quad\text{and}\quad\sup_{\omega\in\Omega}\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,\|\partial_{\omega}g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}<\infty.$
Hence, there exists a uniform constant $C>0$ such that
(4.13) $\|(\mathcal{R}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha}))_{\omega}\|_{2^{*}}\leq C\sup_{\omega\in\Omega}\|{\nu}^{(n)}_{\omega}-{\nu}_{\alpha,\omega}\|_{1^{*}}.$
By Lemma 2.2, the iterates $\bm{\nu}^{(n)}$ converge in Wasserstein (Lipschitz dual norm) to $\bm{\nu}_{\alpha}$, meaning that
$\sup_{\omega\in\Omega}\|{\nu}^{(n)}_{\omega}-{\nu}_{\alpha,\omega}\|_{\mathrm{Lip}^{*}}\longrightarrow 0.$
Since the inclusion $\operatorname{Lip}(X)^{*}\hookrightarrow C^{1}(X)^{*}$ is continuous, it follows that
$\sup_{\omega\in\Omega}\|{\nu}^{(n)}_{\omega}-{\nu}_{\alpha,\omega}\|_{1^{*}}\longrightarrow 0.$
Combining this with (4.13), the proof is complete. ∎
Remark 4.14.
The regularity of the observable $\phi$ is only relevant for the estimate of $\mathcal{R}_{2}$, and one might hope to take $\phi\in C^{1}(X)$, with convergence of measures in $C^{0}(X)^{*}$, which would allow a stronger final result in the space of sections valued in $C^{1}(X)^{*}$. However, the crucial step in the proof uses the convergence of $\mathcal{K}_{\alpha}^{n}\bm{\nu}$ to $\bm{\nu}_{\alpha}$ in $\mathrm{Lip}(X)^{*}$ (Wasserstein distance) to deduce convergence in $C^{1}(X)^{*}$. If we attempted to go one unit lower, we would need to deduce convergence in $C^{0}(X)^{*}$ from convergence in $\mathrm{Lip}(X)^{*}$, which is obviously false.
In the following result, we make use of the constant $0<\lambda<1$ provided by (B3). Observe that, by Remark 4.2, we have
(4.14) $\mathcal{C}_{2}(g_{\alpha,\theta_{\alpha,i}})\leq\max\{L_{1}+L_{2},\,L_{1}^{2}\}\leq\max\{L_{1}+L_{2}+L_{3},\,L_{1}^{2}+3L_{1}L_{2},\,L_{1}^{3}\}\leq\lambda.$
Lemma 4.15.
For every $\bm{\xi},\bm{\zeta}\in\mathbf{E}_{2}^{0}$ we have
$\|\mathcal{T}\bm{\xi}-\mathcal{T}\bm{\zeta}\|_{2}\leq\lambda\|\bm{\xi}-\bm{\zeta}\|_{2}.$
Proof.
For any $\bm{\xi},\bm{\zeta}\in\mathbf{E}_{2}^{0}$, we have for all $\omega\in\Omega$
$\|(\mathcal{T}\bm{\xi})_{\omega}-(\mathcal{T}\bm{\zeta})_{\omega}\|_{2^{*}}\leq\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,|\theta_{\alpha,i}^{\prime}(\omega)|\,\|(g_{\alpha,\theta_{\alpha,i}})_{*}(\xi_{\theta_{\alpha,i}(\omega)}-\zeta_{\theta_{\alpha,i}(\omega)})\|_{2^{*}}.$
Since $\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)=1$, it follows from (B1)(B2), Lemma 4.1 and (4.14) that
$\|(\mathcal{T}\bm{\xi})_{\omega}-(\mathcal{T}\bm{\zeta})_{\omega}\|_{2^{*}}\leq\lambda\|\xi_{\theta_{\alpha,i}(\omega)}-\zeta_{\theta_{\alpha,i}(\omega)}\|_{2^{*}}.$
Taking supremum in $\omega\in\Omega$ on both sides of the last inequality we get the conclusion. ∎
Proposition 4.16.
$\bm{\nu}_{\alpha}:\Omega\to C^{2}(X)^{*}$ is a $C^{1}$ map.
Proof.
By Lemma 4.15 and the Banach fixed point theorem, the operator $\mathcal{T}$ admits a unique fixed point $\bm{\xi}^{*}\in\mathbf{E}_{2}^{0}$, that is,
$\mathcal{T}\bm{\xi}^{*}=\bm{\xi}^{*}.$
Fix a smooth section $\bm{\nu}\in\mathbf{P}$ and consider $\bm{\nu}^{(n)}=\mathcal{K}_{\alpha}^{n}\bm{\nu}$. Define $\bm{\xi}^{(n)}\in\mathbf{E}_{2}^{0}$ by
(4.15) $\xi^{(n)}_{\omega}=\partial_{\omega}{\nu}^{(n)}_{\omega}$.
We define the error at step $n\geq 0$ by
$E_{n}=\|\bm{\xi}^{(n)}-\bm{\xi}^{*}\|_{2}.$
From the iterative relation
$\xi^{(n+1)}_{\omega}=\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,\theta_{\alpha,i}^{\prime}(\omega)\,(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\xi_{\theta_{\alpha,i}(\omega)}+\big(\mathcal{R}\bm{\nu}^{(n)}\big)_{\omega}.$
and the fixed point equation
$\xi^{*}_{\omega}=(\mathcal{T}\bm{\xi}^{*})_{\omega}=\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,\theta_{\alpha,i}^{\prime}(\omega)\,(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\xi_{\theta_{\alpha,i}(\omega)}+\big(\mathcal{R}\bm{\nu}_{\alpha}\big)_{\omega},$
we obtain
$\xi^{(n+1)}_{\omega}-\xi^{*}_{\omega}$ $=\sum_{i\in\mathbb{I}}p_{\alpha,i}(\omega)\,\theta_{\alpha,i}^{\prime}(\omega)\,(g_{\alpha,\theta_{\alpha,i}(\omega)})_{*}\big(\xi_{\theta_{\alpha,i}(\omega)}^{(n)}-\xi_{\theta_{\alpha,i}(\omega)}\big)+\big(\big(\mathcal{R}\bm{\nu}^{(n)}\big)_{\omega}-\big(\mathcal{R}\bm{\nu}_{\alpha}\big)_{\omega}\big)$
$=\big(\mathcal{T}(\bm{\xi}^{(n)}-\bm{\xi}^{*})\big)_{\omega}+\big(\mathcal{R}(\bm{\nu}^{(n)}-\bm{\nu}_{\alpha})\big)_{\omega}$
Taking the norm in $C^{2}(X)^{*}$, supremum over $\omega\in\Omega$ and using Lemma 4.15, we get
$E_{n+1}\leq\lambda E_{n}+\varepsilon_{n}.$
Observing that by Lemma 4.13 we have $\varepsilon_{n}\to 0$ as $n\to\infty$, standard estimates for linear recursions with vanishing forcing terms give
(4.16) $E_{n}\leq\lambda^{n}E_{0}+\sum_{k=0}^{n-1}\lambda^{k}\varepsilon_{n-1-k}\to 0,\quad\text{as }n\to\infty.$
Now, for any $\omega_{0},\omega\in\Omega$, the fundamental theorem of calculus in the space $C^{2}(X)^{*}$ yields
${\nu}^{(n)}_{\omega}={\nu}^{(n)}_{\omega_{0}}+\int_{\omega_{0}}^{\omega}\xi^{(n)}_{\theta}\,d\theta.$
Taking the limit $n\to\infty$, we have
${\nu}^{(n)}_{\omega}\to{\nu}_{\alpha,\omega}\quad\text{and}\quad{\nu}^{(n)}_{\omega_{0}}\to{\nu}_{\alpha,\omega_{0}},$
with convergence in $\operatorname{Lip}(X)^{*}$, by Lemma 2.2, hence in $C^{2}(X)^{*}$, by the continuous inclusion of $\operatorname{Lip}(X)^{*}$ in $C^{2}(X)^{*}$. On the other hand, it follows from (4.16) that
$\left\|\int_{\omega_{0}}^{\omega}\xi^{(n)}_{\theta}d\theta-\int_{\omega_{0}}^{\omega}\xi^{*}_{\theta}d\theta\right\|_{2^{*}}\leq|\omega-\omega_{0}|\sup_{\theta\in\Omega}\|\xi^{(n)}_{\theta}-\xi^{*}_{\theta}\|_{2^{*}}=|\omega-\omega_{0}|E_{n}\to 0.$
Thus,
${\nu}_{\alpha,\omega}={\nu}_{\alpha,\omega_{0}}+\int_{\omega_{0}}^{\omega}\xi^{*}_{\theta}\,d\theta.$
Since $\bm{\xi}^{*}\in\mathbf{E}_{2}^{0}$, the fundamental theorem of calculus gives that $\bm{\nu}_{\alpha}$ is a $C^{1}$ map from $\Omega$ to $C^{2}(X)^{*}$, with derivative $\bm{\nu}_{\alpha}^{\prime}=\bm{\xi}^{*}.$
4.4.2. Parameter differentiability
Here we prove that the map $\alpha\mapsto\mathcal{K}_{\alpha}\bm{\nu}_{\alpha_{0}}$ is differentiable in $\mathbf{E}_{3}^{0}$ at any $\alpha_{0}\in J$, thus obtaining (A3)(A3)(d). We denote differentiation with respect to the parameter $\alpha$ by a dot over the corresponding function. For notational convenience, we extend this convention to operators to indicate the derivative of the associated maps with respect to $\alpha$. We emphasize that this notation is purely formal; in particular, no claim is made regarding the existence of derivatives in the uniform operator norm. We will make use of the following lemma, whose proof follows from the chain rule for second-order derivatives by standard arguments.
Lemma 4.17.
For any $\phi\in C^{3}(X)$ and any two maps $h_{1},h_{2}\in C^{2}(X,X)$, we have
  1. (1)
    $\|\phi\circ h_{1}-\phi\circ h_{2}\|_{C^{2}}\leq C\sum_{k=0}^{2}\|D^{k}h_{1}-D^{k}h_{2}\|_{C^{0}}$, with
    $C=\max\Big\{\|\phi\|_{C^{1}},\ \|\phi\|_{C^{2}}(1+L_{1}+L_{2}),\ \|\phi\|_{C^{3}}(1+L_{1}^{2})\Big\},$
    where $L_{j}=\max\{\|D^{j}h_{1}\|_{C^{0}},\|D^{j}h_{2}\|_{C^{0}}\}$;
  2. (2)
    $\|\phi\circ h\|_{C^{2}}\leq C^{\prime}\|\phi\|_{C^{2}}$, with
    $C^{\prime}=\max\big\{1,\ \operatorname{diam}(X)+\|Dh\|_{C^{0}}+\|D^{2}h\|_{C^{0}},\ \|Dh\|_{C^{0}}^{2}\big\}.$
Proposition 4.18.
Assume that $\bm{\nu}\in\mathbf{P}\cap\mathbf{E}_{2}^{\prime}$. Then,
  1. (1)
    the map $\alpha\mapsto\mathcal{K}_{\alpha,i}\bm{\nu}$ is differentiable in $\mathbf{E}_{3}^{0}$, for each $i\in\mathbb{I}$;
  2. (2)
    the series $\sum_{i\in\mathbb{I}}\dot{\mathcal{K}}_{\alpha,i}\bm{\nu}$ converges in $\mathbf{E}_{3}^{0}$ uniformly in $\alpha$.
Proof.
For the first item, we are going to prove that the derivative $\dot{\mathcal{K}}_{\alpha,i}\bm{\nu}$ of $\mathcal{K}_{\alpha,i}\bm{\nu}$ with respect to the parameter $\alpha\in J$ is given for any $\omega\in\Omega$ and $\phi\in C^{3}(X)$ by
$\langle(\dot{\mathcal{K}}_{\alpha,i}\bm{\nu})_{\omega},\phi\rangle=$ $\;\dot{p}_{\alpha,i}(\omega)\,\langle\nu_{\theta_{\alpha,i}(\omega)},\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle$
$+p_{\alpha,i}(\omega)\,\langle\nu^{\prime}_{\theta_{\alpha,i}(\omega)}\dot{\theta}_{\alpha,i}(\omega),\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle$
$+p_{\alpha,i}(\omega)\,\langle\nu_{\theta_{\alpha,i}(\omega)},D\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\bm{\cdot}\partial_{\alpha}g_{\alpha,\theta_{\alpha,i}(\omega)}\rangle$
(4.17) $+p_{\alpha,i}(\omega)\,\langle\nu_{\theta_{\alpha,i}(\omega)},D\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\bm{\cdot}\partial_{\omega^{\prime}}g_{\alpha,\theta_{\alpha,i}(\omega)}\,\dot{\theta}_{\alpha,i}(\omega)\rangle.$
Given any $\alpha_{0}\in J$ and ${\varepsilon}\neq 0$, set $\alpha=\alpha_{0}+{\varepsilon}$, $\omega_{\varepsilon}=\theta_{\alpha,i}(\omega)$, and $\omega_{0}=\theta_{\alpha_{0},i}(\omega)$. For $\phi\in C^{3}(X)$, consider the difference quotient
$\langle\Delta_{\varepsilon}(\omega),\phi\rangle=\frac{1}{{\varepsilon}}\Big[\langle(\mathcal{K}_{\alpha,i}\bm{\nu})_{\omega},\phi\rangle-\langle(\mathcal{K}_{\alpha_{0},i}\bm{\nu})_{\omega},\phi\rangle\Big].$
We decompose
$\langle\Delta_{\varepsilon}(\omega),\phi\rangle-\langle(\dot{\mathcal{K}}_{\alpha_{0},i}\bm{\nu})_{\omega},\phi\rangle=\langle A_{\varepsilon}(\omega),\phi\rangle+\langle B_{\varepsilon}(\omega),\phi\rangle+\langle C_{\varepsilon}(\omega),\phi\rangle+\langle D_{\varepsilon}(\omega),\phi\rangle$
with
$\langle A_{\varepsilon}(\omega),\phi\rangle$ $=\left(\frac{p_{\alpha,i}(\omega)-p_{\alpha_{0},i}(\omega)}{{\varepsilon}}\right)\langle\nu_{\omega_{\varepsilon}},\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle-\dot{p}_{\alpha_{0},i}(\omega)\langle\nu_{\omega_{0}},\phi\circ g_{\alpha_{0},\omega_{0}}\rangle,$
$\langle B_{\varepsilon}(\omega),\phi\rangle$ $=p_{\alpha_{0},i}(\omega)\left[\left\langle\frac{\nu_{\omega_{\varepsilon}}-\nu_{\omega_{0}}}{{\varepsilon}},\phi\circ g_{\alpha,\omega_{\varepsilon}}\right\rangle-\langle\nu^{\prime}_{\omega_{0}}\dot{\theta}_{\alpha_{0},i}(\omega),\phi\circ g_{\alpha_{0},\omega_{0}}\rangle\right],$
$\langle C_{\varepsilon}(\omega),\phi\rangle$ $=p_{\alpha_{0},i}(\omega)\left\langle\nu_{\omega_{0}},\frac{\phi\circ g_{\alpha,\omega_{\varepsilon}}-\phi\circ g_{\alpha_{0},\omega_{\varepsilon}}}{{\varepsilon}}-D\phi(g_{\alpha_{0},\omega_{0}})\bm{\cdot}\partial_{\alpha}g_{\alpha_{0},\omega_{0}}\right\rangle,$
$\langle D_{\varepsilon}(\omega),\phi\rangle$ $=p_{\alpha_{0},i}(\omega)\left\langle\nu_{\omega_{0}},\frac{\phi\circ g_{\alpha_{0},\omega_{\varepsilon}}-\phi\circ g_{\alpha_{0},\omega_{0}}}{{\varepsilon}}-D\phi(g_{\alpha_{0},\omega_{0}})\bm{\cdot}\partial_{\omega}g_{\alpha_{0},\omega_{0}}\dot{\theta}_{\alpha_{0},i}(\omega)\right\rangle.$
We need to show that each of these four terms converges to 0 as $\varepsilon\to 0$, uniformly in $\omega\in\Omega$ and $\phi\in C^{3}(X)$ with $\|\phi\|_{C^{3}}\leq 1$.
1. Term $A_{\varepsilon}$: By the mean value theorem applied to $\alpha\mapsto p_{\alpha,i}(\omega)$, there exists $\alpha^{\prime}=\alpha_{0}+{\varepsilon}^{\prime}$ with $|{\varepsilon}^{\prime}|<|{\varepsilon}|$ such that
$\frac{1}{\varepsilon}(p_{\alpha,i}(\omega)-p_{\alpha_{0},i}(\omega))=\dot{p}_{\alpha^{\prime},i}(\omega).$
Substituting this into the expression for $\langle A_{\varepsilon}(\omega),\phi\rangle$, we decompose
$|\langle$ $A_{\varepsilon}(\omega),\phi\rangle|=|\dot{p}_{\alpha^{\prime},i}(\omega)\langle\nu_{\omega_{\varepsilon}},\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle-\dot{p}_{\alpha_{0},i}(\omega)\langle\nu_{\omega_{0}},\phi\circ g_{\alpha_{0},\omega_{0}}\rangle|$
$\leq|(\dot{p}_{\alpha^{\prime},i}(\omega)-\dot{p}_{\alpha_{0},i}(\omega))\langle\nu_{\omega_{\varepsilon}},\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle|+|\dot{p}_{\alpha_{0},i}(\omega)(\langle\nu_{\omega_{\varepsilon}},\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle-\langle\nu_{\omega_{0}},\phi\circ g_{\alpha_{0},\omega_{0}}\rangle)|$
$\leq\underbrace{|\dot{p}_{\alpha^{\prime},i}(\omega)-\dot{p}_{\alpha_{0},i}(\omega)|}_{\text{(I)}}\|\phi\|_{C^{0}}+\|\dot{p}_{\alpha_{0},i}\|_{C^{0}}\underbrace{|\langle\nu_{\omega_{\varepsilon}},\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle-\langle\nu_{\omega_{0}},\phi\circ g_{\alpha_{0},\omega_{0}}\rangle|}_{\text{(II)}}.$
Regarding term (II), we apply a telescoping
$|\langle\nu_{\omega_{\varepsilon}},\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle-\langle\nu_{\omega_{0}},\phi\circ g_{\alpha_{0},\omega_{0}}\rangle|$ $\leq|\langle\nu_{\omega_{\varepsilon}}-\nu_{\omega_{0}},\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle|+|\langle\nu_{\omega_{0}},\phi\circ g_{\alpha,\omega_{\varepsilon}}-\phi\circ g_{\alpha_{0},\omega_{0}}\rangle|$
$\leq\underbrace{\|\nu_{\omega_{\varepsilon}}-\nu_{\omega_{0}}\|_{2^{*}}\|\phi\circ g_{\alpha,\omega_{\varepsilon}}\|_{C^{2}}}_{\text{(III)}}+\underbrace{\|\phi\circ g_{\alpha,\omega_{\varepsilon}}-\phi\circ g_{\alpha_{0},\omega_{0}}\|_{C^{0}}}_{\text{(IV)}}.$
It follows from Lemma 4.17 that
(4.18) $\|\phi\circ g_{\alpha,\omega_{\varepsilon}}\|_{C^{2}}\leq\max\big\{1,\operatorname{diam}(X)+\|Dg_{\alpha,\omega_{\varepsilon}}\|_{C^{0}}+\|D^{2}g_{\alpha,\omega_{\varepsilon}}\|_{C^{0}}+\|Dg_{\alpha,\omega_{\varepsilon}}\|_{C^{0}}^{2}\big\}\|\phi\|_{C^{2}},$
which together with the mean value theorem applied to $\alpha\mapsto\nu_{\theta_{\alpha,i}(\omega)}$, yields
$\text{(III)}\leq\varepsilon\|\bm{\nu}^{\prime}\|_{2}\|{\dot{\theta}}_{\alpha,i}\|_{C^{0}}\max\big\{1,\operatorname{diam}(X)+\|Dg_{\alpha,\omega_{\varepsilon}}\|_{C^{0}}+\|D^{2}g_{\alpha,\omega_{\varepsilon}}\|_{C^{0}}+\|Dg_{\alpha,\omega_{\varepsilon}}\|_{C^{0}}^{2}\big\}\|\phi\|_{C^{2}}$
By the mean value theorem, we also have
$\text{(IV)}\leq\|\phi\|_{C^{1}}\|g_{\alpha,\omega_{\varepsilon}}-g_{\alpha_{0},\omega_{0}}\|_{C_{0}}.$
By the expressions obtained for (I)–(IV), taking the supremum over all $\phi$ with $\|\phi\|_{C^{3}}\leq 1$ and using (B1)-(B4), we conclude that $\|A_{\varepsilon}(\omega)\|_{3^{*}}\to 0$ uniformly in $\omega$ as $\varepsilon\to 0$.
2. Term $B_{\varepsilon}$: By the mean value theorem applied to the map $\alpha\mapsto\nu_{\theta_{\alpha,i}(\omega)}$, there exists some $\alpha^{\prime}=\alpha_{0}+\varepsilon^{\prime}$ with $|\varepsilon^{\prime}|<|\varepsilon|$ such that
$\frac{1}{\varepsilon}(\nu_{\omega_{\varepsilon}}-\nu_{\omega_{0}})=\nu^{\prime}_{\omega_{\varepsilon^{\prime}}}\dot{\theta}_{\alpha^{\prime},i}(\omega).$
Substituting this in $\langle B_{\varepsilon}(\omega),\phi\rangle$, we get
$\langle$ $B_{\varepsilon}(\omega),\phi\rangle=|\langle\nu^{\prime}_{\omega_{\varepsilon^{\prime}}}\dot{\theta}_{\alpha^{\prime},i}(\omega),\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle-\langle\nu^{\prime}_{\omega_{0}}\dot{\theta}_{\alpha_{0},i}(\omega),\phi\circ g_{\alpha_{0},\omega_{0}}\rangle|$
$\leq|\langle\nu^{\prime}_{\omega_{\varepsilon^{\prime}}}\dot{\theta}_{\alpha^{\prime},i}(\omega)-\nu^{\prime}_{\omega_{0}}\dot{\theta}_{\alpha_{0},i}(\omega),\phi\circ g_{\alpha,\omega_{\varepsilon}}\rangle|+|\langle\nu^{\prime}_{\omega_{0}}\dot{\theta}_{\alpha_{0},i}(\omega),\phi\circ g_{\alpha,\omega_{\varepsilon}}-\phi\circ g_{\alpha_{0},\omega_{0}}\rangle|$
$\leq\underbrace{\|\nu^{\prime}_{\omega_{\varepsilon^{\prime}}}\dot{\theta}_{\alpha^{\prime},i}(\omega)-\nu^{\prime}_{\omega_{0}}\dot{\theta}_{\alpha_{0},i}(\omega)\|_{2^{*}}}_{\text{(I)}}\|\phi\circ g_{\alpha,\omega_{\varepsilon}}\|_{C^{2}}+\|\bm{\nu}^{\prime}\|_{2}\|{\dot{\theta}}_{\alpha_{0},i}\|_{C^{0}}\underbrace{\|\phi\circ g_{\alpha,\omega_{\varepsilon}}-\phi\circ g_{\alpha_{0},\omega_{0}}\|_{C^{2}}}_{\text{(II)}}.$
To properly bound term (I), note that
$\|\nu^{\prime}_{\omega_{\varepsilon^{\prime}}}\dot{\theta}_{\alpha^{\prime},i}(\omega)-\nu^{\prime}_{\omega_{0}}\dot{\theta}_{\alpha_{0},i}(\omega)\|_{2^{*}}$ $\leq\|(\nu^{\prime}_{\omega_{\varepsilon^{\prime}}}-\nu^{\prime}_{\omega_{0}})\dot{\theta}_{\alpha^{\prime},i}(\omega)\|_{2^{*}}+\|\nu^{\prime}_{\omega_{0}}(\dot{\theta}_{\alpha^{\prime},i}(\omega)-\dot{\theta}_{\alpha_{0},i}(\omega))\|_{2^{*}}$
$\leq\|\nu^{\prime}_{\omega_{\varepsilon^{\prime}}}-\nu^{\prime}_{\omega_{0}}\|_{2^{*}}|\dot{\theta}_{\alpha^{\prime},i}(\omega)|+\|\nu^{\prime}_{\omega_{0}}\|_{2^{*}}|\dot{\theta}_{\alpha^{\prime},i}(\omega)-\dot{\theta}_{\alpha_{0},i}(\omega)|$
$\leq\|\nu^{\prime}_{\omega_{\varepsilon^{\prime}}}-\nu^{\prime}_{\omega_{0}}\|_{2^{*}}\|{\dot{\theta}}_{\alpha^{\prime},i}\|_{C^{0}}+\|\bm{\nu}^{\prime}\|_{2}|\dot{\theta}_{\alpha^{\prime},i}(\omega)-\dot{\theta}_{\alpha_{0},i}(\omega)|.$
Regarding term (II), Lemma 4.17 yields the estimate
$\bigl\|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}-\phi\circ g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}\bigr\|_{C^{2}}\leq C\sum_{k=0}^{2}\|D^{k}g_{\alpha,\theta_{\alpha,i}(\omega)}-D^{k}g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}\|_{C^{0}},$
where
$C=\max\Big\{\|\phi\|_{C^{1}},\;$ $\|\phi\|_{C^{2}}\bigl(1+\|g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{2}}+\|g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}\|_{C^{2}}\bigr),$
$\|\phi\|_{C^{3}}\bigl(1+\|g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{1}}^{2}+\|g_{\alpha_{0},\theta_{\alpha_{0},i}(\omega)}\|_{C^{1}}^{2}\bigr)\Big\}.$
By the expressions obtained for (I)–(II), taking the supremum over all $\phi$ with $\|\phi\|_{C^{3}}\leq 1$, using (B1)(B4) together with (4.18) and the fact that $\bm{\nu}\in\mathbf{E}_{2}^{\prime}$ (and hence $\bm{\nu}$ is uniformly continuous), we conclude that $\|B_{\varepsilon}(\omega)\|_{3^{*}}\to 0$ uniformly in $\omega$ as $\varepsilon\to 0$.
3. Term $C_{\varepsilon}$: By the fundamental theorem of calculus applied to the map $\alpha\mapsto\phi\circ g_{\alpha,\omega_{\varepsilon}}$, for each $x\in X$ we have
$\frac{1}{\varepsilon}(\phi(g_{\alpha_{0}+\varepsilon,\omega_{\varepsilon}}(x))-\phi(g_{\alpha_{0},\omega_{\varepsilon}}(x)))=\int_{0}^{1}\nabla\phi(g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}(x))\cdot\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}(x)\,dt.$
Substituting this identity into the expression for $\langle C_{\varepsilon}(\omega),\phi\rangle$ and applying Fubini’s theorem to exchange the action of the measure $\nu_{\omega_{0}}$ with the integral over $t\in[0,1]$, we write
$\langle C_{\varepsilon}(\omega),\phi\rangle$ $=\int_{0}^{1}\Big(\langle\nu_{\omega_{0}},(\nabla\phi\circ g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}})\cdot\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}\rangle-\langle\nu_{\omega_{0}},(\nabla\phi\circ g_{\alpha_{0},\omega_{0}})\cdot\partial_{\alpha}g_{\alpha_{0},\omega_{0}}\rangle\Big)\,dt.$
We telescope the integrand by adding and subtracting the term $\langle\nu_{\omega_{0}},(\nabla\phi\circ g_{\alpha_{0},\omega_{0}})\cdot\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}\rangle$. This yields:
$\langle C_{\varepsilon}(\omega),\phi\rangle$ $=\int_{0}^{1}\langle\nu_{\omega_{0}},(\nabla\phi\circ g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}-\nabla\phi\circ g_{\alpha_{0},\omega_{0}})\cdot\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}\rangle\,dt$
(4.19) $\quad+\int_{0}^{1}\langle\nu_{\omega_{0}},(\nabla\phi\circ g_{\alpha_{0},\omega_{0}})\cdot(\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}-\partial_{\alpha}g_{\alpha_{0},\omega_{0}})\rangle\,dt.$
For the first integrand, we define the vector-valued measure ${\mu}_{t,\varepsilon}=(\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}})\nu_{\omega_{0}}$, so that it represents the difference of the push-forwards of ${\mu}_{t,\varepsilon}$ by the maps $h_{1}=g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}$ and $h_{0}=g_{\alpha_{0},\omega_{0}}$ acting on the gradient field $\nabla\phi$:
$\langle\nu_{\omega_{0}},(\nabla\phi\circ g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}-\nabla\phi\circ g_{\alpha_{0},\omega_{0}})\cdot\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}\rangle=\langle(g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}})_{*}{\mu}_{t,\varepsilon}-(g_{\alpha_{0},\omega_{0}})_{*}{\mu}_{t,\varepsilon},\nabla\phi\rangle.$
Applying Lemma 4.7 component-wise to the vector measure ${\mu}_{t,\varepsilon}$ with the test function $\nabla\phi\in C^{2}(X)$, and using the fact that $\|{\mu}_{t,\varepsilon}\|_{0^{*}}\leq\|\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\theta_{\alpha,i}}\|_{C^{0}}$, we obtain
$|\langle\nu_{\omega_{0}},(\nabla\phi$ $\circ g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}-\nabla\phi\circ g_{\alpha_{0},\omega_{0}})\cdot\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}\rangle|$
$\leq\|{\mu}_{t,\varepsilon}\|_{0^{*}}\|\nabla\phi\|_{C^{1}}\|g_{\alpha_{0}+t\varepsilon,\theta_{\alpha,i}}-g_{\alpha_{0},\theta_{\alpha_{0},i}}\|_{C^{0}}$
$\leq\|\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\theta_{\alpha,i}}\|_{C^{0}}\|\nabla\phi\|_{C^{1}}\|g_{\alpha_{0}+t\varepsilon,\theta_{\alpha,i}}-g_{\alpha_{0},\theta_{\alpha_{0},i}}\|_{C^{0}}$
(4.20) $\leq\|\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\theta_{\alpha,i}}\|_{C^{0}}\|\phi\|_{C^{2}}\|g_{\alpha_{0}+t\varepsilon,\theta_{\alpha,i}}-g_{\alpha_{0},\theta_{\alpha_{0},i}}\|_{C^{0}}.$
For the second integral, recalling that $\bm{\nu}\in\mathbf{P}$ and using Cauchy-Schwarz inequality,
$|\langle\nu_{\omega_{0}},(\nabla\phi\circ g_{\alpha_{0},\omega_{0}})$ $\cdot(\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}-\partial_{\alpha}g_{\alpha_{0},\omega_{0}})\rangle|$
$\leq\|\nabla\phi\|_{C^{0}}\|\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\theta_{\alpha,i}}-\partial_{\alpha}g_{\alpha_{0},\theta_{\alpha_{0},i}}\|_{C^{0}}$
(4.21) $\leq\|\phi\|_{C^{1}}\|\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\theta_{\alpha,i}}-\partial_{\alpha}g_{\alpha_{0},\theta_{\alpha_{0},i}}\|_{C^{0}}.$
Taking the supremum over $\|\phi\|_{C^{3}}\leq 1$ in (4.19)–(4.21), we conclude:
$\|C_{\varepsilon}(\omega)\|_{3^{*}}\leq\int_{0}^{1}\Big($ $\|\dot{g}_{\alpha_{0},\theta_{\alpha_{0},i}}\|_{C^{0}}\|g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}-g_{\alpha_{0},\omega_{0}}\|_{C^{0}}+\|\partial_{\alpha}g_{\alpha_{0}+t\varepsilon,\omega_{\varepsilon}}-\partial_{\alpha}g_{\alpha_{0},\omega_{0}}\|_{C^{0}}\Big)\,dt.$
Since (B3) holds, the integrand above vanishes uniformly in $\omega\in\Omega$, as $\varepsilon\to 0$.
4. Term $D_{\varepsilon}$: By the mean value theorem applied to $\alpha\mapsto\theta_{\alpha}(\omega)$, for each $\varepsilon\neq 0$ there exists $\alpha^{\prime}=\alpha_{0}+\varepsilon^{\prime}$ with $|\varepsilon^{\prime}|<|\varepsilon|$ such that
$\frac{\theta_{\alpha_{0}+\varepsilon}(\omega)-\theta_{\alpha_{0}}(\omega)}{\varepsilon}=\dot{\theta}_{\alpha^{\prime}}(\omega).$
Let $\omega_{\varepsilon}=\theta_{\alpha_{0}+\varepsilon}(\omega)$. We express the difference quotient in $\langle D_{\varepsilon}(\omega),\phi\rangle$ using the fundamental theorem of calculus on the map $\omega^{\prime}\mapsto\phi\circ g_{\alpha_{0},\omega^{\prime}}$ along the path from $\omega_{0}$ to $\omega_{\varepsilon}$:
$\langle D_{\varepsilon}(\omega),\phi\rangle=\dot{\theta}_{\alpha^{\prime}}(\omega)\int_{0}^{1}$ $\langle\nu_{\omega_{0}},\nabla\phi\circ g_{\alpha_{0},\omega_{t}}\cdot\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{t}}\rangle\,dt$
(4.22) $\quad-\dot{\theta}_{\alpha_{0}}(\omega)\langle\nu_{\omega_{0}},\nabla\phi\circ g_{\alpha_{0},\omega_{0}}\cdot\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{0}}\rangle,$
where $\omega_{t}=\omega_{0}+t(\omega_{\varepsilon}-\omega_{0})$. We telescope the expression by adding and subtracting the term $\dot{\theta}_{\alpha^{\prime}}(\omega)\langle\nu_{\omega_{0}},\nabla\phi\circ g_{\alpha_{0},\omega_{0}}\cdot\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{t}}\rangle$:
$\langle D_{\varepsilon}(\omega),\phi\rangle$ $=\dot{\theta}_{\alpha^{\prime}}(\omega)\int_{0}^{1}\langle\nu_{\omega_{0}},(\nabla\phi\circ g_{\alpha_{0},\omega_{t}}-\nabla\phi\circ g_{\alpha_{0},\omega_{0}})\cdot\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{t}}\rangle\,dt$
(4.23) $\quad+\langle\nu_{\omega_{0}},\nabla\phi\circ g_{\alpha_{0},\omega_{0}}\cdot(\dot{\theta}_{\alpha^{\prime}}(\omega)\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{t}}-\dot{\theta}_{\alpha_{0}}(\omega)\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{0}})\rangle.$
Defining the vector-valued measure $\mu_{t,\varepsilon}=(\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{t}})\nu_{\omega_{0}}$, the integrand is the difference of push-forwards of $\mu_{t,\varepsilon}$ by $h_{1}=g_{\alpha_{0},\omega_{t}}$ and $h_{0}=g_{\alpha_{0},\omega_{0}}$. Applying Lemma 4.7 component-wise to the vector measure $\mu_{t,\varepsilon}$ with the test function $\nabla\phi\in C^{2}(X)$, and using that $\|\mu_{t,\varepsilon}\|_{0^{*}}\leq\|\partial_{\omega^{\prime}}g_{\alpha_{0},\theta_{\alpha,i}}\|_{C^{0}}$, we obtain:
$|\dot{\theta}_{\alpha^{\prime}}(\omega)\langle(g_{\alpha_{0},\omega_{t}})_{*}\mu_{t,\varepsilon}$ $-(g_{\alpha_{0},\omega_{0}})_{*}\mu_{t,\varepsilon},\nabla\phi\rangle|$
$\leq|\dot{\theta}_{\alpha^{\prime}}(\omega)|\|\mu_{t,\varepsilon}\|_{0^{*}}\|\nabla\phi\|_{C^{1}}\|g_{\alpha_{0},\omega_{t}}-g_{\alpha_{0},\omega_{0}}\|_{C^{0}}$
(4.24) $\leq|\dot{\theta}_{\alpha^{\prime}}(\omega)|\|\partial_{\omega^{\prime}}g\|_{C^{0}}\|\phi\|_{C^{2}}\|g_{\alpha_{0},\omega_{t}}-g_{\alpha_{0},\omega_{0}}\|_{C^{0}}.$
For the second part of (4.23), we have:
$|\langle\nu_{\omega_{0}},\nabla\phi\circ g_{\alpha_{0},\omega_{0}}\cdot$ $(\dot{\theta}_{\alpha^{\prime}}(\omega)\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{t}}-\dot{\theta}_{\alpha_{0}}(\omega)\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{0}})\rangle|$
(4.25) $\leq\|\phi\|_{C^{1}}\|\dot{\theta}_{\alpha^{\prime}}(\omega)\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{t}}-\dot{\theta}_{\alpha_{0}}(\omega)\partial_{\omega^{\prime}}g_{\alpha_{0},\omega_{0}}\|_{C^{0}}.$
Taking the supremum over $\|\phi\|_{C^{3}}\leq 1$ and using (B1)(B2), we conclude that $\|D_{\varepsilon}(\omega)\|_{3^{*}}$ vanishes uniformly in $\omega$ as $\varepsilon\to 0$. This finishes the proof of the first item.
To prove the second item, using that $\bm{\nu}\in\mathbf{P}\cap\mathbf{E}_{2}^{\prime}$ and the expression obtained in (4.17), we deduce for any $\phi\in C^{3}(X)$
$|\langle(\dot{\mathcal{K}}_{\alpha,i}\bm{\nu})_{\omega},\phi\rangle|\leq$ $\;|\dot{p}_{\alpha,i}(\omega)|\cdot\|\nu_{\theta_{\alpha,i}(\omega)}\|_{0^{*}}\|\phi\|_{C^{0}}$
$+p_{\alpha,i}(\omega)\cdot\|\nu^{\prime}_{\theta_{\alpha,i}(\omega)}\dot{\theta}_{\alpha,i}(\omega)\|_{2^{*}}\|\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{2}}$
$+p_{\alpha,i}(\omega)\cdot\|\nu_{\theta_{\alpha,i}(\omega)}\|_{0^{*}}\|D\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\bm{\cdot}\partial_{\alpha}g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{0}}$
$+p_{\alpha,i}(\omega)\cdot\|\nu_{\theta_{\alpha,i}(\omega)}\|_{0^{*}}\|D\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\bm{\cdot}\partial_{\omega^{\prime}}g_{\alpha,\theta_{\alpha,i}(\omega)}\dot{\theta}_{\alpha,i}(\omega)\|_{C^{0}}.$
Using Lemma 4.17, we get
$\|\phi\circ g_{\alpha,\theta_{\alpha,i}}\|_{C^{2}}\leq\underbrace{\max\big\{1,\operatorname{diam}(X)+\|Dg_{\alpha,\theta_{\alpha,i}}\|_{C^{0}}+\|D^{2}g_{\alpha,\theta_{\alpha,i}}\|_{C^{0}}+\|Dg_{\alpha,\theta_{\alpha,i}}\|_{C^{0}}^{2}\big\}}_{C(g_{\alpha,\theta_{\alpha,i}})}\|\phi\|_{C^{2}},$
By Cauchy-Schwarz inequality, we have
$\|D\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\bm{\cdot}\partial_{\alpha}g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{0}}\leq\|\phi\|_{C^{1}}\|\partial_{\alpha}g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{0}}$
and
$\|D\phi\circ g_{\alpha,\theta_{\alpha,i}(\omega)}\bm{\cdot}\partial_{\omega^{\prime}}g_{\alpha,\theta_{\alpha,i}(\omega)}\dot{\theta}_{\alpha,i}(\omega)\|_{C^{0}}\leq\|\phi\|_{C^{1}}\|\partial_{\omega^{\prime}}g_{\alpha,\theta_{\alpha,i}(\omega)}\|_{C^{0}}|\dot{\theta}_{\alpha,i}(\omega)|.$
Taking the supremum over $\|\phi\|_{C^{3}}\leq 1$ and over $\omega\in\Omega$, we get
$\|\dot{\mathcal{K}}_{\alpha,i}\bm{\nu}\|_{3^{*}}\leq$ $\;\|\dot{p}_{\alpha,i}\|_{C^{0}}\|\bm{\nu}\|_{\mathrm{L}}$
$+\|p_{\alpha,i}\|_{C^{0}}\|\bm{\nu}^{\prime}\|_{2}C(g_{\alpha,\theta_{\alpha,i}})$
$+\|p_{\alpha,i}\|_{C^{0}}\|\bm{\nu}\|_{\mathrm{L}}\|\partial_{\alpha}g_{\alpha,\theta_{\alpha,i}}\|_{C^{0}}$
$+\|p_{\alpha,i}\|_{C^{0}}\|\bm{\nu}\|_{\mathrm{L}}\|\partial_{\omega^{\prime}}g_{\alpha,\theta_{\alpha,i}}\|_{C^{0}}\|\dot{\theta}_{\alpha,i}\|_{C^{0}}.$
Using (B1)(B4), we get the convergence of the series in $\mathbf{E}_{3}^{0}$ uniformly in $\alpha$. ∎
Corollary 4.19.
The map $\alpha\mapsto\mathcal{K}_{\alpha}\bm{\nu}_{\alpha_{0}}$ is differentiable in $\mathbf{E}_{3}^{0}$ at $\alpha_{0}$.
Proof.
Since Proposition 4.16 gives $\bm{\nu}_{\alpha_{0}}\in\mathbf{E}_{2}^{\prime}$, the conclusion follows from Proposition 4.18 below, by standard calculus in the Banach space $\mathbf{E}_{3}^{0}$. ∎
5. Base maps with inducing schemes
In this section, we prove Theorem G. Consider the skew-product map
$T_{\alpha}:\Omega\times X\to\Omega\times X,$
as in (1.1), and the induced map
$\bar{T}_{\alpha}:\bar{\Omega}\times X\to\bar{\Omega}\times X,$
as in (1.15). Assume that (B1)(B4) hold for the induced system $\bar{T}_{\alpha}$, and that (C1)(C2) also hold for the unfolding operator.
5.1. Sectional transfer operators
In this subsection we establish a fundamental relation between the sectional transfer operator of the original skew-product map and that of the induced skew-product map. Recall that the sectional transfer operator for the original system acts on a section $\bm{\nu}$ defined for $\eta_{\alpha}$-almost every $\omega\in\Omega$ as
(5.1) $(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}=\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}\frac{\rho_{\alpha}(\theta)}{\rho_{\alpha}(\omega)|f^{\prime}_{\alpha}(\theta)|}\,(g_{\alpha,\theta})_{*}\nu_{\theta},$
whilst the sectional transfer operator for the induced system acts on a section $\bar{\bm{\nu}}$ for $\bar{\eta}_{\alpha}$-almost every $\omega\in\bar{\Omega}$ as
(5.2) $(\bar{\mathcal{K}}_{\alpha}\bar{\bm{\nu}})_{\omega}=\sum_{\theta\in\bar{f}_{\alpha}^{-1}(\{\omega\})}\frac{\bar{\rho}_{\alpha}(\theta)}{\bar{\rho}_{\alpha}(\omega)|\bar{f}_{\alpha}^{\prime}(\theta)|}\,(\bar{g}_{\alpha,\theta})_{*}\bar{\nu}_{\theta}.$
For any $k\geq 1$ and $\theta\in f_{\alpha}^{-k}(\{\omega\})$, define the accumulated weight along the $k$-step path as
(5.3) $W_{\alpha}(\theta,k)=\frac{\rho_{\alpha}(\theta)}{\rho_{\alpha}(f_{\alpha}^{k}(\theta))\,|(f_{\alpha}^{k})^{\prime}(\theta)|}.$
Then, the action of $\mathcal{K}_{\alpha}^{k}$ on the section $\bm{\nu}$ is given for $\eta_{\alpha}$-almost every $\omega\in\Omega$ by
(5.4) $(\mathcal{K}_{\alpha}^{k}\bm{\nu})_{\omega}=\sum_{\theta\in f_{\alpha}^{-k}(\{\omega\})}W_{\alpha}(\theta,k)\,(g_{\alpha,\theta}^{(k)})_{*}\nu_{\theta},$
with the map $g_{\alpha,\theta}^{(k)}$ as in (1.16). The following lemma provides an explicit representation of the sectional transfer operator of the induced system in terms of the original operator.
Lemma 5.1.
For any section $\bar{\bm{\nu}}$ defined on $\bar{\Omega}$, we have
$\bar{\mathcal{K}}_{\alpha}\bar{\bm{\nu}}=\sum_{k=1}^{\infty}\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}=k\}}).$
Proof.
Consider an $\bar{\eta}_{\alpha}$-generic $\omega\in\bar{\Omega}$. Recalling that $\bar{f}_{\alpha}(\theta)=f_{\alpha}^{\tau_{\alpha}(\theta)}(\theta)$, we partition the set of preimages in (5.2) according to the first return time $n=\tau_{\alpha}(\theta)$,
$\bar{f}_{\alpha}^{-1}(\{\omega\})=\bigcup_{n=1}^{\infty}A_{n}(\omega),$
with
$A_{n}(\omega)=\{\theta\in\bar{\Omega}\mid f_{\alpha}^{n}(\theta)=\omega\text{ and }\tau_{\alpha}(\theta)=n\}.$
Since the density $\rho_{\alpha}$ in $\bar{\Omega}$ coincides with $\bar{\rho}_{\alpha}$, up to a normalising factor, the induced weight satisfies
$\frac{\bar{\rho}_{\alpha}(\theta)}{\bar{\rho}_{\alpha}(\omega)|\bar{f}_{\alpha}^{\prime}(\theta)|}=W_{\alpha}(\theta,n),$
for each $\theta\in A_{n}(\omega)$, with $W_{\alpha}(\theta,n)$ as in (5.3). Moreover, $\bar{g}_{\alpha,\theta}=g_{\alpha,\theta}^{(n)}$. Thus, we may write the action of the induced operator as
(5.5) $(\bar{\mathcal{K}}_{\alpha}\bar{\bm{\nu}})_{\omega}=\sum_{n=1}^{\infty}\sum_{\theta\in A_{n}(\omega)}W_{\alpha}(\theta,n)\big(g_{\alpha,\theta}^{(n)}\big)_{*}\bar{\nu}_{\theta}.$
Next, consider the right-hand side of the desired equality,
$\mathcal{S}_{\omega}=\sum_{k=1}^{\infty}\bigl(\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}=k\}})\bigr)_{\omega}.$
By the formula for the iterated operator in (5.4), we have for each $k\geq 1$
$\bigl(\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}=k\}})\bigr)_{\omega}=\sum_{\theta\in f_{\alpha}^{-k}(\{\omega\})}W_{\alpha}(\theta,k)\,\mathbf{1}_{\{\tau_{\alpha}=k\}}(\theta)\,\big(g_{\alpha,\theta}^{(k)}\big)_{*}\bar{\nu}_{\theta}.$
Because $\bar{\bm{\nu}}$ is supported on $\bar{\Omega}$, the term for a fixed $k$ is non-zero only for $\theta\in\bar{\Omega}$ with $f_{\alpha}^{k}(\theta)=\omega$ and $\tau_{\alpha}(\theta)=k$, that is, precisely the points in $A_{k}(\omega)$. Since the sets $A_{k}(\omega)$ are disjoint and cover all preimages under $\bar{f}_{\alpha}$, summing over $k$ counts each preimage exactly once, specifically in the term where $k=\tau_{\alpha}(\theta)$. Thus,
$\mathcal{S}_{\omega}=\sum_{k=1}^{\infty}\sum_{\theta\in A_{k}(\omega)}W_{\alpha}(\theta,k)\big(g_{\alpha,\theta}^{(k)}\big)_{*}\bar{\nu}_{\theta}.$
Comparing this to (5.5), we see that $\mathcal{S}_{\omega}=(\bar{\mathcal{K}}_{\alpha}\bar{\bm{\nu}})_{\omega}$, which concludes the proof. ∎
In the next result, we establish a pointwise property showing that the sectional transfer operator $\mathcal{K}_{\alpha}$ “commutes” with the infinite series defining an unfolded section.
Lemma 5.2.
Let $\bar{\bm{\nu}}$ be a section on $\bar{\Omega}$. For every $\eta_{\alpha}$-almost every $\omega\in\Omega$, we have
$\Big(\mathcal{K}_{\alpha}\sum_{k=0}^{\infty}\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})\Big)_{\omega}=\sum_{k=0}^{\infty}\left(\mathcal{K}_{\alpha}^{k+1}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})\right)_{\omega}.$
Proof.
Define the section
$\bm{\xi}=\sum_{k=0}^{\infty}\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}}),$
where the sum is understood pointwise. According to (5.4), for any $\theta\in\Omega$, the term
$\bigl(\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})\bigr)_{\theta}$
is non-zero only if there exists $\sigma\in\bar{\Omega}$ such that $f_{\alpha}^{k}(\sigma)=\theta$ and $\tau_{\alpha}(\sigma)>k$. For any $\sigma\in\bar{\Omega}$, the inequality $\tau_{\alpha}(\sigma)>k$ holds for only finitely many $k$. Consequently, for each $\theta$, the above series contains only finitely many non-zero terms and therefore defines a finite sum of measures. Now, by definition of the sectional transfer operator,
$\bigl(\mathcal{K}_{\alpha}\bm{\xi}\bigr)_{\omega}=\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}W_{\alpha}(\theta,1)\,(g_{\alpha,\theta})_{*}\bm{\xi}_{\theta}.$
Substituting the definition of $\bm{\xi}_{\theta}$, we obtain
$\bigl(\mathcal{K}_{\alpha}\bm{\xi}\bigr)_{\omega}=\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}W_{\alpha}(\theta,1)\,(g_{\alpha,\theta})_{*}\left(\sum_{k=0}^{\infty}\bigl(\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})\bigr)_{\theta}\right).$
For each fixed $\theta$, the inner sum over $k$ is finite, as seen above. Therefore, using only the linearity of the push-forward $(g_{\alpha,\theta})_{*}$, we may rewrite the above expression as
$\bigl(\mathcal{K}_{\alpha}\bm{\xi}\bigr)_{\omega}=\sum_{k=0}^{\infty}\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}W_{\alpha}(\theta,1)\,(g_{\alpha,\theta})_{*}\bigl(\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})\bigr)_{\theta}.$
Finally, by the definition of $\mathcal{K}_{\alpha}$, the inner sum is exactly
$\left(\mathcal{K}_{\alpha}\bigl[\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})\bigr]\right)_{\omega}=\left(\mathcal{K}_{\alpha}^{k+1}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})\right)_{\omega}.$
Summing over $k$ yields the desired identity. ∎
5.2. Unfolding operator
Recall that the unfolding operator $\mathcal{U}_{\alpha}$, introduced in (1.17), allows one to reconstruct a full section $\bm{\nu}$ on $\Omega$ starting from a section $\bar{\bm{\nu}}$ defined on the inducing domain $\bar{\Omega}$, by
$\bm{\nu}=\mathcal{U}_{\alpha}\bar{\bm{\nu}}=\sum_{k=0}^{\infty}\mathcal{K}_{\alpha}^{k}\left(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}}\right).$
Lemma 5.3.
If $\bar{\bm{\nu}}$ is a fixed point of the $\bar{\mathcal{K}}_{\alpha}$, then $\mathcal{U}_{\alpha}\bar{\bm{\nu}}$ is a fixed point of $\mathcal{K}_{\alpha}$.
Proof.
First we show that $\mathcal{U}_{\alpha}$ maps a fixed point of $\bar{\mathcal{K}}_{\alpha}$ to a fixed point of ${\mathcal{K}}_{\alpha}$. Let $\bar{\bm{\nu}}$ is a fixed point of the $\bar{\mathcal{K}}_{\alpha}$ and set $\bm{\nu}=\mathcal{U}_{\alpha}(\bar{\bm{\nu}})$. Applying the sectional transfer operator $\mathcal{K}_{\alpha}$ to both sides of the identity defining $\mathcal{U}_{\alpha}(\bar{\bm{\nu}})$ and using Lemma 5.2, we get
$\mathcal{K}_{\alpha}\bm{\nu}$ $=\mathcal{K}_{\alpha}\left(\sum_{k=0}^{\infty}\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})\right)$
$=\sum_{k=0}^{\infty}\mathcal{K}_{\alpha}^{k+1}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})$
(5.6) $=\sum_{k=1}^{\infty}\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k-1\}}).$
Now, observe that
$\mathbf{1}_{\{\tau_{\alpha}>k-1\}}=\mathbf{1}_{\{\tau_{\alpha}=k\}}+\mathbf{1}_{\{\tau_{\alpha}>k\}}.$
Substituting this into (5.6), we obtain
$\mathcal{K}_{\alpha}\bm{\nu}$ $=\sum_{k=1}^{\infty}\mathcal{K}_{\alpha}^{k}\left(\bar{\bm{\nu}}\cdot(\mathbf{1}_{\{\tau_{\alpha}=k\}}+\mathbf{1}_{\{\tau_{\alpha}>k\}})\right)$
$=\underbrace{\sum_{k=1}^{\infty}\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}=k\}})}_{\text{(A)}}+\underbrace{\sum_{k=1}^{\infty}\mathcal{K}_{\alpha}^{k}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>k\}})}_{\text{(B)}}.$
Since $\bar{\bm{\nu}}$ is the fixed point of $\bar{\mathcal{K}}_{\alpha}$, it follows from Lemma 5.1 that
$\text{(A)}=\bar{\mathcal{K}}_{\alpha}\bar{\bm{\nu}}=\bar{\bm{\nu}}.$
On the other hand,
$\text{(B)}=\mathcal{U}_{\alpha}\bar{\bm{\nu}}-\mathcal{K}_{\alpha}^{0}(\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>0\}})={\bm{\nu}}-\bar{\bm{\nu}}\cdot\mathbf{1}_{\{\tau_{\alpha}>0\}}={\bm{\nu}}-\bar{\bm{\nu}}$
Combining the results for (A) and (B), we obtain
$\mathcal{K}_{\alpha}\bm{\nu}=\bar{\bm{\nu}}+(\bm{\nu}-\bar{\bm{\nu}})=\bm{\nu}.$
This proves that $\bm{\nu}$ is a fixed point of $\mathcal{K}_{\alpha}$. ∎
Lemma 5.4.
If $\bar{\bm{\nu}}\in\bar{\mathbf{P}}$, then $\mathcal{U}_{\alpha}\bar{\bm{\nu}}\in\mathbf{P}$.
Proof.
We prove that for $\eta_{\alpha}$-almost every $\omega\in\Omega$, the fiber measure ${\nu}_{\omega}$ has total mass $1$. By the formula for the iterated sectional transfer operator in (5.4),
${\nu}_{\omega}=\sum_{k=0}^{\infty}\sum_{\theta\in f_{\alpha}^{-k}(\{\omega\})\cap\bar{\Omega}}W_{\alpha}(\theta,k)\,(g_{\alpha,\theta}^{(k)})_{*}\bigl(\bar{\nu}_{\theta}\mathbf{1}_{\{\tau_{\alpha}>k\}}(\theta)\bigr),$
where the weights are given by (5.3). Since $\bar{\nu}_{\theta}$ is a probability measure and the push-forward $(g_{\alpha,\theta}^{(k)})_{*}$ preserves mass, we get
(5.7) ${\nu}_{\omega}(X)=\frac{1}{\rho_{\alpha}(\omega)}\sum_{k=0}^{\infty}\sum_{\theta\in f_{\alpha}^{-k}(\omega)\cap\bar{\Omega}}\frac{\rho_{\alpha}(\theta)}{|(f_{\alpha}^{k})^{\prime}(\theta)|}\mathbf{1}_{\{\tau_{\alpha}>k\}}(\theta).$
Now, recall that
(5.8) $\eta_{\alpha}=\frac{1}{{E}}\sum_{k=0}^{\infty}(f_{\alpha}^{k})_{*}\bigl(\bar{\eta}_{\alpha}|_{\{\tau_{\alpha}>k\}}\bigr),\qquad{E}=\int_{\bar{\Omega}}\tau_{\alpha}\,d\bar{\eta}_{\alpha}.$
Writing $d\eta_{\alpha}=\rho_{\alpha}\,dm$ and $d\bar{\eta}_{\alpha}=\bar{\rho}_{\alpha}\,dm$, the density on the induced base is related to the global density by $\bar{\rho}_{\alpha}(\theta)={E}\rho_{\alpha}(\theta)$, for $\theta\in\bar{\Omega}$. For any Borel set $A\subset\Omega$, the push-forward integral reads
$(f_{\alpha}^{k})_{*}\bigl(\bar{\eta}_{\alpha}|_{\{\tau_{\alpha}>k\}}\bigr)(A)=\int_{(f_{\alpha}^{k})^{-1}(A)\cap\{\tau_{\alpha}>k\}}\bar{\rho}_{\alpha}(\theta)\,dm(\theta).$
Applying the change of variables formula under the local diffeomorphism $f_{\alpha}^{k}$,
$\int_{(f_{\alpha}^{k})^{-1}(A)\cap\{\tau_{\alpha}>k\}}\bar{\rho}_{\alpha}(\theta)\,dm(\theta)=\int_{A}\sum_{\theta\in f_{\alpha}^{-k}(\omega)\cap\bar{\Omega}}\frac{\bar{\rho}_{\alpha}(\theta)}{|(f_{\alpha}^{k})^{\prime}(\theta)|}\mathbf{1}_{\{\tau_{\alpha}>k\}}(\theta)\,dm(\omega).$
Substituting into (5.8) and using $\bar{\rho}_{\alpha}(\theta)={E}\rho_{\alpha}(\theta)$, we obtain
$\eta_{\alpha}(A)=\frac{1}{{E}}\sum_{k=0}^{\infty}\int_{A}\sum_{\theta\in f_{\alpha}^{-k}(\omega)\cap\bar{\Omega}}\frac{{E}\rho_{\alpha}(\theta)}{|(f_{\alpha}^{k})^{\prime}(\theta)|}\mathbf{1}_{\{\tau_{\alpha}>k\}}(\theta)\,dm(\omega).$
Hence,
$\int_{A}\rho_{\alpha}(\omega)\,dm(\omega)=\int_{A}\sum_{k=0}^{\infty}\sum_{\theta\in f_{\alpha}^{-k}(\omega)\cap\bar{\Omega}}\frac{\rho_{\alpha}(\theta)}{|(f_{\alpha}^{k})^{\prime}(\theta)|}\mathbf{1}_{\{\tau_{\alpha}>k\}}(\theta)\,dm(\omega).$
By the Radon–Nikodym theorem, the integrands must coincide $\nu_{\alpha}$-almost everywhere:
$\rho_{\alpha}(\omega)=\sum_{k=0}^{\infty}\sum_{\theta\in f_{\alpha}^{-k}(\omega)\cap\bar{\Omega}}\frac{\rho_{\alpha}(\theta)}{|(f_{\alpha}^{k})^{\prime}(\theta)|}\mathbf{1}_{\{\tau_{\alpha}>k\}}(\theta).$
Finally, substituting this into (5.7) gives ${\nu}_{\omega}(X)=1$, for $\omega$ almost every $\omega\in\Omega$. ∎
We are now in conditions to conclude the proof of Theorem G. Since conditions (B1)(B4) for $\bar{T}_{\alpha}$, together with (C1), imply that (A2) holds for $\bar{T}_{\alpha}$, we apply Theorem E to deduce:
  • $\bar{\mathcal{K}}_{\alpha}$ has a unique fixed point $\bar{\bm{\nu}}_{\alpha}$ in $\bar{\mathbf{P}}_{0}$;
  • the map $\alpha\mapsto\bar{\bm{\nu}}_{\alpha}$ is differentiable in $\bar{\mathbf{E}}_{3}^{0}$.
It follows from Lemma 5.3 and Lemma 5.4 that
(5.9) $\bm{\nu}_{\alpha}=\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha}).$
is a fixed point for $\mathcal{K}_{\alpha}$ in $\mathbf{P}$. In the next result we establish the last conclusion of Theorem G.
Proposition 5.5.
The map $\alpha\mapsto\bm{\nu}_{\alpha}$ is differentiable in $\mathbf{L}_{3}^{0}$ at $\alpha_{0}$.
Proof.
By (C2)(C2)(c) the map $\alpha\mapsto\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha_{0}})$ is differentiable in $\mathbf{L}_{3}^{0}$ at $\alpha_{0}$. Let $\dot{\mathcal{U}}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})\in\mathbf{L}_{3}^{0}$ denote this derivative. We are going to show that
(5.10) $\dot{\bm{\nu}}_{\alpha_{0}}=\dot{\mathcal{U}}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})+\mathcal{U}_{\alpha_{0}}(\dot{\bar{\bm{\nu}}}_{\alpha_{0}})$
is the derivative of $\alpha\mapsto\bm{\nu}_{\alpha}$ at $\alpha_{0}$. Since (5.9) holds for every $\alpha$, we may write
$\frac{\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}}{\alpha-\alpha_{0}}$ $=\frac{\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha})-\mathcal{U}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})}{\alpha-\alpha_{0}}$
$=\frac{\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha})-\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha_{0}})+\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha_{0}})-\mathcal{U}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})}{\alpha-\alpha_{0}}$
$=\mathcal{U}_{\alpha}\left(\frac{\bar{\bm{\nu}}_{\alpha}-\bar{\bm{\nu}}_{\alpha_{0}}}{\alpha-\alpha_{0}}\right)+\frac{\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha_{0}})-\mathcal{U}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})}{\alpha-\alpha_{0}}.$
To show differentiability, we estimate the distance to the candidate derivative in (5.10). By the triangle inequality,
$\left\|\frac{\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}}{\alpha-\alpha_{0}}-\dot{\bm{\nu}}_{\alpha_{0}}\right\|_{\mathbf{L}_{3}}$ $\leq\left\|\mathcal{U}_{\alpha}\left(\frac{\bar{\bm{\nu}}_{\alpha}-\bar{\bm{\nu}}_{\alpha_{0}}}{\alpha-\alpha_{0}}\right)-\mathcal{U}_{\alpha_{0}}(\dot{\bar{\bm{\nu}}}_{\alpha_{0}})\right\|_{\mathbf{L}_{3}}$
$\quad+\left\|\frac{\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha_{0}})-\mathcal{U}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})}{\alpha-\alpha_{0}}-\dot{\mathcal{U}}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})\right\|_{\mathbf{L}_{3}}.$
We split the first term on the right-hand side
$\left\|\mathcal{U}_{\alpha}\left(\frac{\bar{\bm{\nu}}_{\alpha}-\bar{\bm{\nu}}_{\alpha_{0}}}{\alpha-\alpha_{0}}\right)-\mathcal{U}_{\alpha_{0}}(\dot{\bar{\bm{\nu}}}_{\alpha_{0}})\right\|_{\mathbf{L}_{3}}$ $\leq\left\|\mathcal{U}_{\alpha}\left(\frac{\bar{\bm{\nu}}_{\alpha}-\bar{\bm{\nu}}_{\alpha_{0}}}{\alpha-\alpha_{0}}-\dot{\bar{\bm{\nu}}}_{\alpha_{0}}\right)\right\|_{\mathbf{L}_{3}}$
$\quad+\left\|\mathcal{U}_{\alpha}(\dot{\bar{\bm{\nu}}}_{\alpha_{0}})-\mathcal{U}_{\alpha_{0}}(\dot{\bar{\bm{\nu}}}_{\alpha_{0}})\right\|_{\mathbf{L}_{3}}.$
Combining these and using (C2)(C2)(a), we get
$\left\|\frac{\bm{\nu}_{\alpha}-\bm{\nu}_{\alpha_{0}}}{\alpha-\alpha_{0}}-\dot{\bm{\nu}}_{\alpha_{0}}\right\|_{\mathbf{L}_{3}}$ $\leq C\left\|\frac{\bar{\bm{\nu}}_{\alpha}-\bar{\bm{\nu}}_{\alpha_{0}}}{\alpha-\alpha_{0}}-\dot{\bar{\bm{\nu}}}_{\alpha_{0}}\right\|_{3}$
$\quad+\left\|\mathcal{U}_{\alpha}(\dot{\bar{\bm{\nu}}}_{\alpha_{0}})-\mathcal{U}_{\alpha_{0}}(\dot{\bar{\bm{\nu}}}_{\alpha_{0}})\right\|_{\mathbf{L}_{3}}$
$\quad+\left\|\frac{\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha_{0}})-\mathcal{U}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})}{\alpha-\alpha_{0}}-\dot{\mathcal{U}}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})\right\|_{\mathbf{L}_{3}}.$
Taking the limit as $\alpha\to\alpha_{0}$, the first term vanishes because $\dot{\bar{\bm{\nu}}}_{\alpha_{0}}$ is the derivative of the map $\alpha\mapsto\bar{\bm{\nu}}_{\alpha}$ in $\bar{\mathbf{E}}_{3}^{0}$ at $\alpha_{0}$; the second term vanishes due to the continuity of the map $\alpha\mapsto\mathcal{U}_{\alpha}(\dot{\bar{\bm{\nu}}}_{\alpha_{0}})$ in $\mathbf{L}_{3}^{0}$, as stated in (C2)(C2)(b); the third term vanishes because $\dot{\mathcal{U}}_{\alpha_{0}}(\bar{\bm{\nu}}_{\alpha_{0}})$ is the derivative of the map $\alpha\mapsto\mathcal{U}_{\alpha}(\bar{\bm{\nu}}_{\alpha_{0}})$ at $\alpha_{0}$. Hence, the difference quotient converges to $\dot{\bm{\nu}}_{\alpha_{0}}$ in $\mathbf{L}_{3}^{0}$, completing the proof. ∎
5.3. Application: Solenoid with intermittency
Here we show that the solenoidal maps with intermittency $\{T_{\alpha}\}_{\alpha\in J}$, introduced in Subsection 1.4.2, admit inducing schemes for which the conditions of Theorem G are satisfied.
5.3.1. Induced map
Here we introduce the induced maps associated with and verify assumptions (B1)(B4) (for the induced maps) and (C1). Let $\Omega=\Omega_{0}\cup\Omega_{1}$ be the $m$ mod 0 partition into the smoothness domains of the base map $f_{\alpha}$. For any $\alpha\in J$, consider the two branches of $f_{\alpha}$ given by
$f_{\alpha,0}(\omega)=\omega(1+2^{\alpha}\omega^{\alpha}),\quad\omega\in\Omega_{0}$
and
$f_{\alpha,1}(\omega)=2\omega-1,\quad\omega\in\Omega_{1}.$
Given any $n\in\mathbb{N}_{0}$, let $i$ be a word $10^{n}$ (a single 1 followed by $n$ 0’s). Let also $\ell(i)=n+1$ be the length of such a word $i$. Set $\mathbb{I}$ as the (countable) set of those words $i$. Define for each $\omega\in[0,1]$
(5.11) $\Theta_{\alpha,i}(\omega)=f_{\alpha,1}^{-1}\circ\underbrace{f_{\alpha,0}^{-1}\circ\cdots\circ f_{\alpha,0}^{-1}}_{n\text{-times}}(\omega).$
The cylinder sets $i]_{\alpha}=\Theta_{\alpha,i}(\bar{\Omega})$ give the $m$ mod 0 partition of $\bar{\Omega}=[1/2,1]$ into smoothness domains of the full branch induced map
${\bar{f}_{\alpha}}:\bar{\Omega}\to\bar{\Omega},$
defined as
(5.12) $\bar{f}_{\alpha}|_{[i]_{\alpha}}={f^{\ell(i)}_{\alpha}|}_{[i]_{\alpha}}.$
Note that $\bar{\theta}_{\alpha,i}=\Theta_{\alpha,i}|_{\bar{\Omega}}$ is the inverse branch of $\bar{f}_{\alpha}|_{[i]_{\alpha}}$. Now we check (B1)(B4) for the family of induced maps.
(B1): this follows from the fact that $\bar{f}_{\alpha}$, as defined above, is a full-branch, uniformly expanding map.
(B2): this corresponds to $C^{1}$ norm of the transfer operator associated with $\bar{f}_{\alpha}$, evaluated at the constant function $1$. It is is well known that this $C^{1}$-norm is uniformly bounded independent of $\alpha$, in a small neighbourhood of $\alpha_{0}$. This is because in a small neighbourhood of $\alpha_{0}$, the expansion is also uniform in $\alpha$, $\bar{f}_{\alpha}$ is $C^{2}$ with $C^{2}$ extensions, and bounded distortion with a bound independent of $\alpha$.
(B3): this follows from the contraction that we impose on fibres.
(B4): since we are dealing with full branch systems that $\bar{\rho}_{\alpha}$, the invariant density of $\bar{f}_{\alpha}$, is uniformly bounded above and below, then to verify
$\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\|\bar{p}_{\alpha,i}\|_{C^{1}}<C,$
it is enough to verify
$\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\|\Theta_{\alpha,i}\|_{C^{1}}<\infty,$
which follows as in (B1). Now, to verify the rest of the assumption, we need to verify
(5.13) $\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\|\dot{\bar{p}}_{\alpha,i}\|_{C^{0}}\leq C_{p}.$
Notice that $\|\partial_{\alpha}{\bar{\rho}_{\alpha}|}_{\alpha_{0}}\|_{C_{0}}<C$ by [8, Theorem 3.2] and since the density $\bar{\rho}_{\alpha}$ is uniformly bounded above and below, verifying (5.13), reduces to showing
$\sum_{i\in\mathbb{I}}\sup_{\alpha\in J}\|\partial_{\alpha}\Theta_{\alpha,i}\|_{C^{0}}<\infty,$
which follows by [8, Lemma 5.1].
(C1): It is well known that $\bar{f}_{\alpha}$ satisfies (C1) with the first return time function $\tau_{\alpha}:\bar{\Omega}\to\mathbb{R}$ satisfying
$m\{\tau_{\alpha}>k\}\approx k^{-1/\alpha}.$
Therefore, integrable with respect to the Lebesgue measure on $\bar{\Omega}$.
5.3.2. Unfolding operator
Here we verify the three items of assumption (C2). Using (1.14) and [8, Equation (7)], we obtain the following formula for the density $\rho_{\alpha}$ of the absolutely continuous invariant probability measure $\eta_{\alpha}$ for $f_{\alpha}$:
(5.14) $\rho_{\alpha}=\frac{1}{E}\sum_{i\in\mathbb{I}}(\bar{\rho}_{\alpha}\circ\Theta_{\alpha,i})\Theta_{\alpha,i}^{\prime},$
where $E>0$ is the normalising constant in (1.14). Noting that $\bar{\rho}_{\alpha}=E\rho_{\alpha}|_{\bar{\Omega}}$, it follows that
$\sum_{i\in\mathbb{I}}\frac{\rho_{\alpha}\circ\Theta_{\alpha,i}}{\rho_{\alpha}}\Theta_{\alpha,i}^{\prime}=1.$
In this case, the unfolding operator can be written for all section $\bar{\bm{\nu}}$ on $\bar{\Omega}$ and all $\omega\in\Omega$ as
(5.15) $(\mathcal{U}_{\alpha}\bar{\bm{\nu}})_{\omega}=\sum_{i\in\mathbb{I}}\frac{\rho_{\alpha}(\Theta_{\alpha,i}(\omega))}{\rho_{\alpha}(\omega)}\Theta^{\prime}_{\alpha,i}(\omega)\,(g^{\ell(i)}_{\Theta_{\alpha,i}(\omega)})_{*}\bar{\nu}_{\Theta_{\alpha,i}(\omega)}.$
In the next result we establish (C2)(C2)(a). In fact, the proof shows that the constant can be taken to be $C=1$.
Lemma 5.6.
$\mathcal{U}_{\alpha}$ defines a uniformly bounded linear operator from $\bar{\mathbf{E}}_{3}^{0}$ to $\bm{\mathbf{L}}_{3}^{0}$.
Proof.
Given any $\bar{\bm{\nu}}\in\bar{\mathbf{E}}_{3}^{0}$, we may write for all $\omega\in\Omega$ and $\phi\in C^{3}_{0}(X)^{*}$
$\begin{split}\langle(\mathcal{U}_{\alpha}\bar{\bm{\nu}})_{\omega},\phi\rangle&=\sum_{i\in\mathbb{I}}\frac{\rho_{\alpha}(\Theta_{\alpha,i}(\omega))}{\rho_{\alpha}(\omega)}\Theta^{\prime}_{\alpha,i}(\omega)\,\langle\bar{\nu}_{\Theta_{\alpha,j}(\omega)},\phi\circ g^{\ell(i)}_{\Theta_{\alpha,i}(\omega)}\rangle\\ &\leq\sum_{j=1}^{\infty}\frac{\rho_{\alpha}(\Theta_{\alpha,i}(\omega))}{\rho_{\alpha}(\omega)}\Theta^{\prime}_{\alpha,i}(\omega)\,\|\bar{\nu}_{\Theta_{\alpha,j}(\omega)}\|_{3}\|\phi\|_{C^{3}}.\end{split}$
Taking the supremum over all test functions $\phi\in C^{3}_{0}(X)^{*}$ with $\|\phi\|_{C^{3}}\leq 1$, we obtain
$\|(\mathcal{U}_{\alpha}\bar{\bm{\nu}})_{\omega}\|_{3}\leq\|\bar{\nu}_{\Theta_{\alpha,j}(\omega)}\|_{3},$
for all $\omega\in\Omega$. Since $\bar{\bm{\nu}}\in\bar{\mathbf{E}}_{3}^{0}$, we have that $\|\bar{\nu}_{\Theta_{\alpha,j}(\omega)}\|_{3}\leq\|\bar{\bm{\nu}}\|_{3}$, and so $\|(\mathcal{U}_{\alpha}\bar{\bm{\nu}})_{\omega}\|_{3}\leq\|\bar{\bm{\nu}}\|_{3}$ for all $\omega\in\Omega$. This implies that $\mathcal{U}_{\alpha}\bar{\bm{\nu}}\in\mathbf{L}_{3}^{0}$ and $\|\mathcal{U}_{\alpha}\bar{\bm{\nu}}\|_{\mathbf{L}_{3}}\leq\|\bar{\bm{\nu}}\|_{3}$, thus showing that $\mathcal{U}_{\alpha}$ defines a bounded linear operator from $\bar{\mathbf{E}}_{3}^{0}$ to $\bm{\mathbf{L}}_{3}^{0}$ with $\|\mathcal{U}_{\alpha}\|\leq 1$. ∎
Before verifying the second and third conditions in (C2), we present some preparatory material, essentially collecting a few facts from [8]. Given $\alpha_{0}\in(0,1)$, let $r>1$ be such that $r<1/\alpha_{0}$. It follows from (1.18) that
(5.16) $\rho_{\alpha_{0}}\in L^{r}(m).$
Let $s>1$ be such that
(5.17) $\frac{1}{r}+\frac{1}{s}=1.$
Fix $\varepsilon>0$ so that $\varepsilon s<1$ and set
$J_{\alpha_{0}}=(\alpha_{0}-\varepsilon,\alpha_{0}+\varepsilon).$
Define $\mathcal{B}$ as the set of continuous functions $\phi:(0,1]\to\mathbb{R}$ such that $|\omega^{\varepsilon}\phi(\omega)|$ is bounded, and define
(5.18) $\|\phi\|_{\mathcal{B}}=\sup_{\omega\in(0,1]}|\omega^{\varepsilon}\phi(\omega)|.$
It is straightforward to check that $\|\cdot\|_{\varepsilon}$ is a norm on the vector space $\mathcal{B}$ and $\mathcal{B}$ endowed with this norm is a Banach space.
Lemma 5.7.
There exists $C>0$ such that for all $\phi\in\mathcal{B}$,
$\int|\phi|d\eta_{\alpha_{0}}\leq C\|\phi\|_{\mathcal{B}}.$
Proof.
Consider $s>1$ as in (5.17). Given $\phi\in\mathcal{B}$, we may write
(5.19) $\int_{[0,1]}|\phi(\omega)|^{s}dm(\omega)=\int_{[0,1]}\frac{|\omega^{\varepsilon}\phi(\omega)|^{s}}{\omega^{\varepsilon s}}dm(\omega)\leq\|\phi\|_{\mathcal{B}}^{s}\int_{[0,1]}\frac{1}{\omega^{\varepsilon s}}dm(\omega).$
Since we take $\varepsilon s<1$, this last integral is finite. On the other hand, using (5.16), (5.17) and Hölder inequality, we get
$\int|\phi|d\eta_{\alpha_{0}}=\int|\phi|\rho_{\alpha_{0}}dm\leq\left(\int|\phi|^{s}dm\right)^{1/s}\left(\int\rho_{\alpha_{0}}^{r}dm\right)^{1/r},$
which together with (5.19), yields the conclusion. ∎
Set
$\alpha_{*}=\alpha_{0}-\frac{1}{2}\varepsilon\quad\text{and}\quad\gamma=\alpha_{0}+\frac{1}{2}\varepsilon.$
For each $i\in\mathbb{I}$, define $h_{\alpha,i}=\rho_{\alpha}\circ\Theta_{\alpha,i}$. Notice that by [8, Theorem 3.1] and the chain rule, the derivatives $\dot{\rho}_{\alpha}$ and $\dot{h}_{\alpha,i}$ exist in the following sense: for $\alpha\in(\alpha_{0}-\varepsilon/4,\alpha_{0}+\varepsilon/4)$
(5.20) $\lim_{\alpha\to\alpha_{0}}\sup_{\omega\in(0,1]}\omega^{\alpha_{*}+\varepsilon}\Big|\frac{1}{\alpha-\alpha_{0}}(h_{\alpha,j}-h_{\alpha_{0},j})-\dot{h}_{\alpha_{0},j}\Big|=0$
and
(5.21) $\lim_{\alpha\to\alpha_{0}}\sup_{\omega\in(0,1]}\omega^{\alpha_{*}+\varepsilon}\Big|\frac{1}{\alpha-\alpha_{0}}(\rho_{\alpha}(\omega)-\rho_{\alpha_{0}}(\omega))-\dot{\rho}_{\alpha_{0}}(\omega)\Big|=0.$
Lemma 5.8.
For all $\alpha\in(\alpha_{0}-\varepsilon/4,\alpha_{0}+\varepsilon/4)$ and $i\in\mathbb{I}$, we have
$\alpha\longmapsto\frac{{\rho}_{\alpha}(\Theta_{\alpha,i}(\omega))}{\rho_{\alpha}(\omega)}\Theta^{\prime}_{\alpha,i}(\omega)$
is differentiable in $\mathcal{B}$ at $\alpha_{0}$.
Proof.
We first notice that for all $\phi\in\mathcal{B}$, we have
(5.22) $\|\phi\|_{\mathcal{B}}=\sup_{\omega\in(0,1]}|\omega^{\alpha_{*}+\varepsilon}{\omega^{-\alpha_{*}}}\phi(\omega)|.$
Thus, recalling (1.18), we have
$\Big|\frac{1}{\omega^{\alpha_{*}}}\left(\frac{1}{\rho_{\alpha}(\omega)}-\frac{1}{\rho_{\alpha_{0}}(\omega)}\right)\Big|\leq C\omega^{\alpha+\alpha_{0}-\alpha_{*}}|\rho_{\alpha}(\omega)-\rho_{\alpha_{0}}(\omega)|\leq C\omega^{\alpha_{0}+\frac{\varepsilon}{4}}|\rho_{\alpha}(\omega)-\rho_{\alpha_{0}}(\omega)|.$
Therefore, by the result of [8] the right hand side of the above inequality also converges to 0 as $\alpha\to\alpha_{0}$. Using this, together with (5.20), (5.21) and (5.22) shows that
$\lim_{\alpha\to\alpha_{0}}\Big\|\frac{1}{\alpha-\alpha_{0}}\left(\frac{h_{\alpha,i}}{\rho_{\alpha}}-\frac{h_{\alpha_{0},i}}{\rho_{\alpha_{0}}}\right)-\Bigg(\frac{\dot{h}_{\alpha_{0},i}\rho_{\alpha_{0}}-\dot{\rho}_{\alpha_{0}}h_{\alpha_{0},i}}{\rho_{\alpha_{0}}^{2}}\Bigg)\Big\|_{\mathcal{B}}=0.$
Noting that
$\|\Theta_{\alpha,i}-\Theta_{\alpha_{0},i}\|_{C^{0}}\to 0\quad\text{and}\quad\|\dot{\Theta}_{\alpha,i}-\dot{\Theta}_{\alpha_{0},i}\|_{C^{0}}\to 0,\quad\text{as $\alpha\to\alpha_{0}$}$
and
$\sup_{i\in\mathbb{I}}\|\Theta_{\alpha,i}\|_{C^{0}}<\infty.$
completes the proof of the second part of the lemma. ∎
The next result establishes (C2)(C2)(c).
Lemma 5.9.
The map $\alpha\mapsto\mathcal{U}_{\alpha}({\bar{\bm{\nu}}}_{\alpha_{0}})$ is differentiable in $\mathbf{L}_{3}^{0}$ at $\alpha_{0}$.
Proof.
We proceed in two steps. First notice that
$\alpha\longmapsto\frac{\rho_{\alpha}\circ\Theta_{\alpha,i}}{\rho_{\alpha}}\Theta^{\prime}_{\alpha,i}\cdot(g^{j}_{\Theta_{\alpha,i}})_{*}\bar{\nu}_{\Theta_{\alpha,i}}$
is differentiable in $\mathbf{L}_{3}^{0}$ at $\alpha_{0}$. This follows from Lemma 5.8, the fact that $\alpha\mapsto\bar{\bm{\nu}}_{\alpha}$ is differentiable at $\alpha_{0}$ in the the $\bar{\mathbf{E}}_{3}^{0}$ and Lemma 5.7. It remains to show that the norm of the derivatives are summable. We have
$\begin{split}\sum_{i\in\mathbb{I}}\int_{\Omega}\partial_{\alpha}&\left(\frac{\rho_{\alpha}(\Theta_{\alpha,i})}{\rho_{\alpha}(\omega)|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}\langle\bar{\nu}_{\Theta_{\alpha,i}},\varphi\circ g^{j}_{\Theta_{\alpha,i}}\rangle\right)d\eta_{\alpha_{0}}(\omega)\\ &\leq\underbrace{\sum_{i\in\mathbb{I}}\int_{\Omega}\partial_{\alpha}\left(\frac{\rho_{\alpha}(\Theta_{\alpha,i})}{\rho_{\alpha}(\omega)|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}\right)\cdot\langle\bar{\nu}_{\Theta_{\alpha,i}},\varphi\circ g^{j}_{\Theta_{\alpha,i}}\rangle d\eta_{\alpha_{0}}(\omega)}_{\text{(I)}}\\ &\quad+\underbrace{\sum_{i\in\mathbb{I}}\int_{\Omega}\frac{\rho_{\alpha}(\Theta_{\alpha,i})}{\rho_{\alpha}(\omega)|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}\partial_{\alpha}\left(\langle\bar{\nu}_{\Theta_{\alpha,i}},\varphi\circ g^{j}_{\Theta_{\alpha,i}}\rangle\right)d\eta_{\alpha_{0}}(\omega)}_{\text{(II)}}.\end{split}$
We have
[…]
For (Ia), we first consider
$\left(\frac{\dot{h}_{\alpha,i}\rho_{\alpha}-\dot{\rho}_{\alpha}h_{\alpha,i}}{\rho_{\alpha}^{2}}\right).$
Since $\Theta_{\alpha,i}\in[\frac{1}{2},1]$, $|\dot{h}_{\alpha,i}|<C$ and ${h}_{\alpha,i}<C$,
(5.23) $\left(\frac{\dot{h}_{\alpha,i}\rho_{\alpha}-\dot{\rho}_{\alpha}h_{\alpha,i}}{\rho_{\alpha}^{2}}\right)\leq C\omega^{\alpha}+C\omega^{2\alpha}\dot{\rho}_{\alpha}<C\omega^{\alpha},$
where we used in the last step the result of [8] to obtain $\omega^{\alpha}\dot{\rho}_{\alpha}<C$. Plugging the estimate (5.23) into (Ia), and using Lemma 5.7, we obtain
$\text{(Ia)}\leq C\|\bar{\bm{\nu}}_{\alpha}\|_{3}\|\varphi\|_{C^{3}}\sum_{i\in\mathbb{I}}\omega^{\alpha+\varepsilon}\frac{1}{|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}$
which is uniformly bounded by [8, Lemma 5.2.a)].
For (Ib) we use the fact that $\Theta_{\alpha,i}\in[\frac{1}{2},1]$. Therefore, $\rho_{\alpha}(\Theta_{\alpha,i})<C$. By Lemma 5.7 and [8, Lemma 5.2.c] we obtain
$\sum_{i\in\mathbb{I}}\omega^{\alpha+\varepsilon}\partial_{\alpha}\left(\frac{1}{|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}\right)<C.$
For the second summand we have
(5.24) $\begin{split}\text{(II)}&\leq C\|\dot{\bar{\bm{\nu}}}_{\alpha}\|_{3}\|\varphi\|_{C^{3}}\sum_{i\in\mathbb{I}}\int_{\Omega}\frac{\rho_{\alpha}(\Theta_{\alpha,i})}{\rho_{\alpha}(\omega)|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}d\eta_{\alpha_{0}}(\omega)\\ &\leq C\|\dot{\bar{\bm{\nu}}}_{\alpha}\|_{3}\|\varphi\|_{C^{3}}\sum_{i\in\mathbb{I}}\int_{\Omega}\frac{\rho_{\alpha}(\Theta_{\alpha,i})}{\rho_{\alpha}(\omega)|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}d\eta_{\alpha_{0}}(\omega)\\ &\leq C\|\dot{\bar{\bm{\nu}}}_{\alpha}\|_{3}\|\varphi\|_{C^{3}}\sum_{i\in\mathbb{I}}\sup_{\omega\in(0,1]}\omega^{\alpha+\varepsilon}\frac{\rho_{\alpha}(\Theta_{\alpha,i})}{|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|},\end{split}$
where in the last step we used (1.18) and Lemma 5.7. Recalling (5.14), we have $\rho_{\alpha}(\Theta_{\alpha,i})<C$ and the remaining sum is uniformly bounded by [8, Lemma 5.2.a]. ∎
The next result establishes condition (C2)(C2)(b).
Lemma 5.10.
For every $\bar{\bm{\nu}}\in\bar{\mathbf{E}}_{3}^{0}$, the map $\alpha\mapsto\mathcal{U}_{\alpha}(\bar{\bm{\nu}})$ is continuous in $\mathbf{L}_{3}^{0}$ at $\alpha_{0}$.
Proof.
First notice that
$\alpha\to\frac{\rho_{\alpha}(\Theta_{\alpha,i})}{\rho_{\alpha}(\omega)|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}(g^{j}_{\Theta_{\alpha,i}})_{*}\bar{\nu}_{\Theta_{\alpha,i}}$
is continuous in $\mathbf{L}_{3}^{0}$ at $\alpha_{0}$. By the same argument as (5.24), for all $\omega\in\Omega$, we have
$\sum_{i\in\mathbb{I}}\int_{\Omega}\frac{\rho_{\alpha}(\Theta_{\alpha,i})}{\rho_{\alpha}(\omega)|(f_{\alpha}^{j})^{\prime}(\Theta_{\alpha,i})|}\langle\bar{\nu}_{\Theta_{\alpha,i}},\varphi\circ g^{j}_{\Theta_{\alpha,i}}\rangle d\eta_{\alpha_{0}}(\omega)\leq C,$
thus concluding the proof. ∎
Appendix A Dual spaces
This appendix collects the definitions and conventions for the functional spaces and norms used in the analysis. These include Lipschitz and smooth observables, their dual spaces, and the Wasserstein–$1$ distance on probability measures, together with the associated push-forward operators. We adopt the standard duality notation $\langle\nu,\phi\rangle$ for the action of a functional $\nu$ on a test function $\phi$.
Lipschitz.  We denote by
$\operatorname{Lip}(X)=\bigl\{\,\phi:X\to\mathbb{R}\;\big|\;\operatorname{Lip}(\phi)<\infty\,\bigr\}$
the space of real-valued Lipschitz functions on $X$, where the Lipschitz seminorm is defined by
$\operatorname{Lip}(\phi)=\sup_{x\neq y}\frac{|\phi(x)-\phi(y)|}{d_{X}(x,y)}.$
We turn this seminorm into a genuine norm by fixing a base point $x_{0}\in X$ and setting
(A.1) $\|\phi\|_{\operatorname{Lip}}=|\phi(x_{0})|+\operatorname{Lip}(\phi).$
Equipped with this norm, $\operatorname{Lip}(X)$ is a Banach space, and the resulting topology is equivalent to the usual one defined using $\|\phi\|_{\infty}+\operatorname{Lip}(\phi)$. We denote by $\operatorname{Lip}(X)^{*}$ the topological dual of $\operatorname{Lip}(X)$, endowed with the dual norm
$\|\nu\|_{\operatorname{Lip}^{*}}=\sup_{\|\phi\|_{\mathrm{L}}\leq 1}\langle\nu,\phi\rangle,\qquad\nu\in\operatorname{Lip}(X)^{*}.$
We further define $\operatorname{Lip}_{0}(X)^{*}$ as the subspace of $\operatorname{Lip}(X)^{*}$ consisting of functionals that vanish on constant functions. For $\nu\in\operatorname{Lip}_{0}(X)^{*}$, adding a constant to $\phi$ does not affect the pairing, and therefore the norm reduces to
(A.2) $\|\nu\|_{\operatorname{Lip}^{*}}=\sup_{\operatorname{Lip}(\phi)\leq 1}\langle\nu,\phi\rangle.$
Wasserstein.  Consider
$\mathcal{M}_{1}(X)=\bigl\{\text{Borel probability measures on }X\bigr\}.$
We equip $\mathcal{M}_{1}(X)$ with the Wasserstein–$1$ distance, defined via the Kantorovich–Rubinstein duality as
(A.3) $W_{1}(\mu,\nu)=\sup_{\mathrm{Lip}(\phi)\leq 1}\Bigl\{\int\phi\,d\mu-\int\phi\,d\nu\Bigr\},\qquad\mu,\nu\in\mathcal{M}_{1}(X).$
Since $X$ is compact, the supremum is finite and $W_{1}$ is well defined. From the dual perspective adopted here, $W_{1}$ coincides with the norm of the signed measure $\mu-\nu$ viewed as an element of $\mathrm{Lip}_{0}(X)^{*}$:
$W_{1}(\mu,\nu)=\|\mu-\nu\|_{\mathrm{Lip}^{*}}.$
We will not appeal directly to the classical optimal transport formulation; instead, we systematically exploit this dual characterization, which is particularly convenient for studying push-forwards and parameter-dependent families of measures. Moreover, under our assumption that $X$ is compact, the metric space $(\mathcal{M}_{1}(X),W_{1})$ is complete and compact, and the Wasserstein–$1$ distance metrizes the weak* topology on probability measures; see [44].
Smooth.  For $k\geq 0$, let $C^{k}(X)$ denote the space of $k$-times continuously differentiable functions $\phi:X\to\mathbb{R}$. For $\phi\in C^{0}(X)$ we set
$\|\phi\|_{C^{0}}=\sup_{x\in X}|\phi(x)|.$
For $k\geq 1$, we introduce a norm analogous to (A.1). Define the seminorm
$|\phi|_{C^{k}}=\sum_{j=1}^{k}\|D^{j}\phi\|_{C^{0}},$
and, fixing a base point $x_{0}\in X$, define the full norm
(A.4) $\|\phi\|_{C^{k}}=|\phi(x_{0})|+|\phi|_{C^{k}}.$
With this norm, $C^{k}(X)$ is a Banach space, and (A.4) is equivalent to the standard $C^{k}$ norm. We denote by $C^{k}(X)^{*}$ the dual space of $C^{k}(X)$, endowed with the dual norm
(A.5) $\|\nu\|_{k^{*}}=\sup_{\|\phi\|_{C^{k}}\leq 1}\langle\nu,\phi\rangle,\qquad\nu\in C^{k}(X)^{*}.$
Let $C^{k}_{0}(X)^{*}\subset C^{k}(X)^{*}$ be the subspace of functionals vanishing on constant functions. For $\nu\in C^{k}_{0}(X)^{*}$, the value of $\langle\nu,\phi\rangle$ depends only on $|\phi|_{C^{k}}$, and hence
(A.6) $\|\nu\|_{k^{*}}=\sup_{|\phi|_{C^{k}}\leq 1}\langle\nu,\phi\rangle.$
Push-forward.  The notion of push-forward for measures extends naturally to the classes of linear functionals introduced above, both in the Lipschitz and in the smooth settings. Let $\nu$ be a continuous linear functional on a space of observables $\operatorname{Lip}(X)$ or $C^{k}(X)$. We define the push-forward $h_{*}\nu$ by duality,
(A.7) $\langle h_{*}\nu,\phi\rangle=\langle\nu,\phi\circ h\rangle,$
whenever the composition $\phi\circ h$ belongs to the same class as $\phi$.
In particular, if $h$ is Lipschitz and $\nu\in\operatorname{Lip}(X)^{*}$, then $h_{*}\nu\in\operatorname{Lip}(X)^{*}$; if $h$ is of class $C^{k}$ and $\nu\in C^{k}(X)^{*}$, then $h_{*}\nu\in C^{k}(X)^{*}$. In both cases, the definition is consistent with the usual push-forward of measures when measures are viewed as elements of the corresponding dual spaces. Moreover, since constant functions are preserved under composition, the push-forward leaves invariant the subspaces $\operatorname{Lip}_{0}(X)^{*}$ and $C^{k}_{0}(X)^{*}$ of functionals vanishing on constant functions.